### v4

# Extracting memorized pieces of (copyrighted) books from open-weight language models

Abstract
Plaintiffs and defendants in copyright lawsuits over generative AI often make sweeping, opposing claims about the extent to which large language models (LLMs) have memorized plaintiffs’ protected expression in their training data. Drawing on both machine learning and copyright law, we show that these polarized positions dramatically oversimplify the relationship between memorization and copyright. To do so, we extend a recent probabilistic extraction technique to measure memorization of $50$ books in $17$ open-weight LLMs. Through thousands of experiments, we show that the extent of memorization varies both by model and by book. With respect to our specific extraction methodology, we find that most LLMs do not memorize most books—either in whole or in part. However, we also find that Llama 3.1 70B entirely memorizes some books, like the first Harry Potter book and 1984. In fact, the first Harry Potter is so memorized that, using a seed prompt consisting of just the first few tokens of the first chapter, we can deterministically generate the entire book near-verbatim. We discuss why our results have significant implications for copyright cases, though not ones that unambiguously favor either side.
1 Introduction
In the dozens of copyright lawsuits over training large language models (LLMs), both sides tend to present opposing, simplified interpretations of the underlying technology. Plaintiffs—such as book authors—say that LLMs are giant copy machines that store (infringing) copies of their specific works and recombine them in outputs [kadrey]. Defendants—typically tech companies—claim that LLMs merely contain non-expressive “statistical correlations” [concordanthropic], and therefore do not copy any particular creative work. We show that the situation is more complicated than either side suggests.
During training, an LLM processes pieces of text in its training data and learns probabilistic relationships across the dataset as a whole. This informal intuition aligns with the “statistical correlations” that defendants describe as encoded in an LLM’s weights. However, sometimes, for some training data, these learned patterns are not so high-level or abstract; rather, they are highly specific [cooper2024files]. For reasons that are not fully understood, certain training data are not so transformed by training: they are “memorized” by the model—encoded in some form inside the LLM’s weights [carlini2023quantifying, lee2023talkin, cooper2024files, carlini2025blog]. Such memorized content can sometimes be “extracted” later; they can be reproduced in an LLM’s outputs at generation time [carlini2021extracting, hayes2025measuringmemorizationlanguagemodels, nasr2023scalable].
Models do not memorize all of their training data, but they do memorize some of it. An LLM may therefore memorize parts of some plaintiffs’ copyrighted works, but not others’. When such memorization occurs, it involves copying that may have significant consequences for ongoing copyright litigation. Extraction generates a “copy” of training data, but it also demonstrates the existence of a “copy” of that training data is memorized inside the model itself. The model being a “copy” (in a technical sense that copyright cares about) has important implications. Notably, a model—not just extracted training data—could be deemed an infringing “copy” of the training data it has memorized [cooper2024files, lee2023talkin]. Copyright law offers the destruction of infringing materials as a remedy. So, just as courts have ordered the destruction of bootleg DVDs, they could order the destruction of infringing models [lee2023talkin, pamscience, townsend2024deletion].
When seeking such remedies, plaintiffs often cite technical papers that demonstrate memorization by extracting training data from open-weight models and production systems. But this reasoning is flawed for a key reason: existing research does not quantify the kinds of information most relevant to a copyright infringement claim [cooper2024files]. This work typically shows that on average models memorize some portion of their training data; it does not show how much models memorize specific works.
To address this gap, we investigate the extent to which open-weight models memorize verbatim content from copyrighted books. For our experiments, we use the Books3 dataset, a (now notorious) torrented corpus of nearly 200,000200{,}000 books [gao2020pile, reisnerbooks3, lee2023talkin]. This dataset is known to be included in the training data of several LLMs [biderman2023pythia, touvron2023llamaopenefficientfoundation, llama2, llama3], and has been a focus of copyright litigation [reisnerbooks3, kadreyamendedconsolidated]. To gauge whether a particular sequence of text is memorized, we leverage a recent probabilistic extraction method [hayes2025measuringmemorizationlanguagemodels] that computes the risk of extracting that sequence at generation time (Section 2).
Extraction risk matters for copyright because the practicability of producing memorized data in a model’s outputs may bear on infringement claims (Section 3). Building on this connection, we develop and validate a procedure that uses probabilistic extraction to surface as much memorization as possible within a book; it uses a sliding window to reveal how much and where memorization occurs, and quantifies the risk that a model will reproduce verbatim sequences in its outputs (Sections 4 & 5). We then run thousands of experiments on $50$ books and $17$ LLMs (Section 6). Our results show that, while most models do not memorize most books either in whole or in part, some models memorize some books entirely, such as the first Harry Potter book [hp1] and 1984 [1984]. Following from these results, we show how to deterministically reconstruct a near-perfect copy of the first Harry Potter using Llama 3.1 70B (Section 7). Last, we discuss the significant implications of these findings for copyright cases, though not ones that unambiguously favor either side. Our results complicate defendants’ fair use story, but they also complicate plaintiffs’ efforts to bring class action lawsuits (Section 8). In summary:
  • Section 2. We provide relevant background on memorization and extraction risk [hayes2025measuringmemorizationlanguagemodels].
  • Section 3. We connect memorization and extraction risk to “copies” in U.S. copyright law.
  • Section 4. We present our sliding-window procedure, which computes extraction risk throughout a book, and enables specific claims about how much an LLM has memorized specific books.
  • Section 5.  We validate that our approach captures true instances of extraction (and thus memorization), not spurious cases where a model generates verbatim non-training data by chance.
  • Section 6. We then run experiments for $17$ LLMs on a set of $50$ books in Books3, which surface novel insights about how book-specific memorization varies across models and books.
  • Section 7. We reconstruct a near-perfect copy of the first Harry Potter using Llama 3.1 70B.
  • Section 8. We offer key takeaways for copyright law and policy.
2 Quantifying memorization through extraction
Extraction and memorization are related concepts, but differ in subtle ways. While extraction refers to reconstructing specific training data in a model’s generated outputs, memorization is broader: it involves reconstructing specific training data by examining the model through any means [cooper2023report, cooper2024files].
Successful extraction of a piece of training data in a model’s generation is evidence of memorization of training data inside the model [cooper2024files, carlini2025blog, schwarzschild2024rethinking, lee2023talkin, feldman2020mem]. This is uncontroversial in machine learning (ML). As Carlini explains, when a sufficiently large and unique piece of training data is extracted, “the only possible explanation is that the model has somewhere internally stored [that piece of training data]” [carlini2025blog]. As Cooper and Grimmelmann note, “in order to be able to extract memorized content from a model at generation time, that memorized content must be encoded in the model’s parameters. There is nowhere else it could be. A model is not a magical portal that pulls fresh information from some parallel universe into our own” [cooper2024files, p. 25].
Models do not memorize all of their training data, but they do memorize some of it. And so, an LLM may memorize parts of some plaintiffs’ copyrighted works, but not others’ (Sections 46). To show this, we first need more background on the probabilistic extraction metric we use in most of our experiments (Section 2.1) and provide guidance on its interpretation (Section 2.2).
2.1 Probabilistic discoverable extraction
Figure 1: Generating the exact completion of a quote from The Great Gatsby [gatsby] with Llama 1 30B.
Figure 2: Plotting $p_{\bm{z}}$ for the “careless people” quote from The Great Gatsby (Figure 1).
Extraction of training data in a model’s outputs—rather than direct inspection of a model’s parameters—is the most common form of evidence used to quantify memorization in LLMs [nasr2025scalable, hayes2025measuringmemorizationlanguagemodels, prashanth2024recite, carlini2023extracting, lee2022dedup, zhang2023counterfactual]. The standard metric in many frontier model release reports [reid2024gemini, gemma2, llama3] measures verbatim discoverable extraction: given a sequence of text from the training data, split it into a prefix and target suffix, prompt the model with the prefix, and deem extraction to be successful if the LLM reproduces the suffix exactly [carlini2021extracting, carlini2023quantifying, nasr2023scalable]. On its own, this type of method is well-suited to completion models that continue the prompt text (as in Figure 1), rather than instruction-tuned chatbots that reply in a conversational style. We therefore only study completion models, referring to these models by name, version, and size. For an example, the middle row of Figure 1 shows verbatim extraction of a quote from The Great Gatsby [gatsby] using Llama 1 30B (Appendix A.1).
Discoverable extraction reduces memorization to a binary (yes-or-no) outcome by issuing a single prompt to the model and checking whether the deterministic, greedy-decoded generation matches the target suffix. But LLMs are probabilistic: under realistic, non-deterministic decoding schemes (e.g., top-$k$ sampling), a model can produce different outputs in response to the same prompt, just as in Figure 1. As cooper2024files note, a copyright-relevant claim is therefore more likely to be concerned with how often a target suffix can be extracted, i.e.:
  • a model, when (a) given a particular type of input, will (b) produce a particular type of memorized output, (c) with a particular probability. That probability could be …1% … [or] … 35% …. The issue for copyright law …is what to do …with the fact that element (c)—the probabilistic element—is inescapable [cooper2024files].
We elaborate on copyright-specific aspects of this point in Sections 3 and 8. For now, we note that hayes2025measuringmemorizationlanguagemodels introduce a metric called probabilistic discoverable extraction (probabilistic extraction, for short), which captures this notion. For a training sequence $\bm{z}$} of length a+ja{+}j, one prompts with the $a$-token prefix ${\bm{z}}_{1:a}$ and computes the probability $p_{\bm{z}}\in[0,1]$ of the LLM $\theta$ generating the exact $j$-token suffix ${\bm{z}}_{a+1:a+j}$ when using decoding scheme $\phi$. That is,
$p_{\bm{z}}\coloneqq p_{{\bm{z}},\theta,\phi}=\Pr_{\theta,\phi}\!\big({\bm{z}}_{a+1:a+j}\,\big|\,{\bm{z}}_{1:a}\big)=\prod_{t=a+1}^{a+j}\Pr_{\theta,\phi}\!\big({\bm{z}}_{t}\,\big|\,{\bm{z}}_{1:t-1}\big).$ (1)
This equation simply captures the probability of the LLM generating the exact suffix ${\bm{z}}_{a+1:a+j}$, obtained by multiplying the conditional probabilities of each token $z_{t}$ given the prefix ${\bm{z}}_{1:a}$ and any earlier suffix tokens. [^1] For numerical stability, we compute this as $\exp$ of the sum of the token conditional $\log$ probabilities. The decoding scheme determines which tokens are considered at each generation step $t$, and how their logits are normalized into conditional probabilities (Section 4.1 & Appendix A.2).
In Figure 2, we plot $p_{\bm{z}}$ for the “careless people” quote using top-$k$ (T=1T{=}1, k=40k{=}40) decoding with Llama 1 models, which were trained on The Great Gatsby [touvron2023llamaopenefficientfoundation]. The probabilities for this sequence are enormous. In general, $p_{{\bm{z}}}$ is a product of many numbers between 0 and $1$, so even relatively short suffixes should have small values—tending toward 0 as sequence length increases. For scale, a $50$-token suffix with $p_{\bm{z}}\!=\!0.5\%$ means that each token has a very high conditional probability of approximately $90\%$ ($0.5\%\!=\!0.005\!\approx\!0.9^{50}$, i.e., the geometric mean of the token conditional probabilities is $0.005^{\frac{1}{50}}\!\approx\!0.9\!=\!90\%$). [^2] A $100$-token suffix where each token has conditional probability $90\%$ has $p_{\bm{z}}=0.9^{100}\!\approx\!0.0027\%$—about $250\times$ less likely than a $50$-token sequence with the same per-token conditional probabilities. The exception is when the model assigns unusually high probabilities to each token in the sequence. For sequences in the model’s training data, this almost always happens only when the model has memorized those sequences.
Determining extraction success.  In traditional discoverable extraction, success means the greedy, deterministic suffix exactly matches the target suffix. Because Equation 1 defines $p_{\bm{z}}$ as the probability that the model generates the target suffix exactly given the prefix, for probabilistic extraction we must decide how high of a probability is “enough” for success. For a training sequence $\bm{z}$} and chosen threshold $\tau_{\text{min}}$ (Appendix C), we consider extraction successful if the target suffix’s probability exceeds $\tau_{\text{min}}$:
$s({\bm{z}};\tau_{\text{min}})\;=\;\mathbf{1}\!\big[\,p_{{\bm{z}}}\geq\tau_{\text{min}}\,\big].$ (2)
In our probabilistic extraction experiments, we always use a suffix length of $50$ tokens, top-$k$ (T=1T{=}1, k=40k{=}40) decoding, and a conservative global threshold of $\tau_{\text{min}}\!=\!0.1\%$, meaning that the (geometric) average token conditional probability is $0.1\%^{\frac{1}{50}}\!\approx\!87.1\%$. We discuss our choice of decoding scheme in detail in Appendix D.2, and the validity of our choice of $\tau_{\text{min}}$ in Section 5.
2.2 Interpreting target suffix probabilities as extraction risk
For a fixed prefix–suffix pair, $p_{{\bm{z}}}$ is the model’s conditional probability (under a given decoding scheme $\phi$) of generating the exact target suffix given the prefix (Equation 1). Equivalently, over independent prompts using the same prefix and $\phi$, $p_{{\bm{z}}}$ is the expected fraction of prompts that result in the LLM reproducing the verbatim target suffix. The expected number of prompts needed for the first reproduction of the target suffix is $\nicefrac{{1}}{{p_{{\bm{z}}}}}$. For example, $p_{{\bm{z}}}\!=\!50\%$ means the model will output roughly $1$ verbatim suffix every $2$ prompts with the prefix; $p_{{\bm{z}}}\!=\!0.1\%$—our choice for $\tau_{\text{min}}$—is about $1$ in 1,0001{,}000 prompts. [^3] In practice, we compute $p_{{\bm{z}}}$ from a single forward pass through the LLM $\theta$ on $\bm{z}$}. We obtain the logits, apply $\phi$ (e.g., greedy, top-$k$) at each suffix token to get the conditional probabilities of the target tokens, and multiply them as in Equation 1. No repeated sampling is required (Appendix A.3). ($p_{{\bm{z}}}\!=\!0\%$ means the suffix will never be generated given the prefix.)
We can therefore interpret $p_{{\bm{z}}}$ as a measure of extraction risk [hayes2025measuringmemorizationlanguagemodels]. For a fixed decoding scheme $\phi$ and prefix from the training data, this enables cross-model comparisons: if $p_{{\bm{z}}}^{(A)}>p_{{\bm{z}}}^{(B)}$, then model $A$ will, in expectation, reproduce the target suffix more often than model $B$ (when prompted with that same prefix). For instance, all of the probabilities in Figure 2 exceed $\tau_{\text{min}}\!=\!0.1\%$: they indicate that every Llama 1 model memorized the “careless people” quote in Figure 1. But the quote is not as easily extractable from each model. Llama 3.1 30B exhibits the highest extraction risk for the quote: in expectation, it would take fewer than $3$ prompts (i.e., $\nicefrac{{1}}{{35.2\%}}\approx 2.84$) with the prefix to extract the suffix.
3 Memorization, extraction risk, and U.S. copyright
The dozens of pending copyright lawsuits concerning LLMs [chatgptiseating] generally present three interrelated issues: (1) whether training an LLM on copyrighted material is permitted fair use (i.e., limited use of the copyrighted material can, under certain circumstances, be used without permission from the copyright owner); (2) whether the model itself is a copy or derivative work (technical terms in copyright law) of the works on which it is trained (Section 8); and (3) whether the model reproduces copyrighted material in its outputs. Some suits present only one issue, while others present all three questions. Some suits are based on content owned by a single company, while others are class action lawsuits purporting to represent all book authors. [^4] There are also other copyright issues [lee2023talkin, pamscience, sag2023safety, goodyear2025infringement, sobelstyle, cooper2024unlearning], for example, whether novel outputs of generative-AI systems are themselves copyrightable [lee2023talkin, uscopyrightofficereport2]. These lawsuits have primarily focused on (1) training and on (3) model outputs, and the two decisions so far have held that training an LLM can be fair use, albeit with significant limitations [kadreyjudgment, bartzjudgment]. They also raise issues about the acquisition, storage, or use of shadow libraries of copyrighted material as training data, such as Books3 [reisnerbooks3, lee2023talkin].
Our work does not speak to the first issue; training and fair use have been discussed extensively elsewhere [lemley2021fairlearning]. Rather, our findings affect the other two issues: we show that copies are sometimes made in the model itself (i.e., memorized) and those copies can sometimes be generated as outputs (i.e., extracted). Because many LLMs are known to have been trained on Books3, we use books from this dataset to study memorization and extraction of copyrighted material (Sections 46).
Memorization and copies inside the model.  If an LLM memorizes all or a substantial portion of a copyrighted work (near-)verbatim, the LLM itself may be or contain an infringing copy or derivative work [lee2023talkin, cooper2024files]. This is because memorized training data are encoded inside the model (Sections 1 & 2). Others have argued that encoding the work in the form of model weights satisfies the technical definition of “copy” in the U.S. Copyright Act [cooper2024files]. In a technical sense, that copy may be of interest to copyright even if there is no practical way to access or extract it. Like the copies of copyrighted works used as training data during training, that copy in model might be fair use, but the analysis of fair use would look somewhat different than internal use of copyrighted works in training processes. This is particularly true for open-weight models like DeepSeek and Llama models, which are not merely used internally by their developers, but are themselves shared externally with others (Section 8).
Extracting copies in model outputs.  Because some memorized works in the training data are extractable, LLMs can sometimes output copies of those works to users [cooper2023report, lee2023talkin, henderson2023fair, lee2024talkinshort, cyphert2024memlaw, sag2024fairness, charlesworth2025]. Works with higher extraction probabilities exhibit greater extraction risk; it takes fewer attempts to produce an output that is a verbatim copy of such a work (Section 2.2). That output at generation time will be judged separately from the model itself, and is less likely to be a fair use [henderson2023fair, lee2023talkin].
Extraction is inherently probabilistic under realistic decoding schemes (Section 2), so the relevant question is not only whether a copy could ever appear for some prompt, but also how often that copy appears in order to be “meaningful” [cooper2024files]. When we measure suffix probabilities $p_{\bm{z}}$ that are very close to $0\%$, we do not count those sequences as extracted; one could perhaps understand such generations as instances of “a monkey at the typewriter” [borelmonkey, simpsons] (Section 5). Very large $p_{\bm{z}}$ (e.g., $90\%$) indicates near-certain reproduction; this is clearly extraction (and therefore evidence of memorization). However, most cases are not this clear, and it is not immediately obvious where to draw a line in the enormous range between these extremes. To avoid false positives, we use a conservative threshold $\tau_{\min}\!=\!0.1\%$ for scientific evidence of extraction (Sections 2.2 & 5). But for copyright law and policy, the practicability of extraction may require a different bar: if reproducing a work just once would require thousands of prompts, the real-world risk may be relatively limited, even if nonzero (Section 8).
Altogether, setting sensible copyright policy requires a specific understanding of both of these sets of issues. However, ML research does not tend to investigate these types of work-specific questions. In the next section, we begin to address this gap by quantifying memorization of individual books.
4 Measuring memorization of individual books
Memorization is the main focus in the next three sections. While the methods we use rely on extraction of $50$-token suffixes (Section 2), we center our claims on what these measurements reveal about memorization inside model weights. We briefly address LLM outputs later (Section 7). To estimate how much a given model memorizes a given book, we introduce a sliding-window procedure that computes probabilistic extraction along the length of a book (Section 4.1). This procedure surfaces book-specific, contextual information about memorization—information that is not readily apparent from average extraction rates over randomly sampled sequences, which prior work typically relies on to quantify memorization (Section 4.2). In the subsequent sections, we assess the validity of our approach (Section 5) and demonstrate how it enables comparisons across models and books (Section 6).
4.1 Sliding-window probabilistic extraction procedure
(a) Each sequence’s extraction probability $p_{\bm{z}}$ by its start position within the overall book
(b) The maximum extraction probability $p_{\bm{z}}$ over all suffixes that span each character position
Figure 3: Visualizing results of the sliding-window probabilistic extraction procedure. For each $100$-token window ($50$-token prefix ++ $50$-token suffix) across 1984 [1984], we compute $p_{\bm{z}}$, the probability that Llama 3.1 70B generates the exact $50$-token suffix given the $50$-token prefix, with respect to top-$k$ decoding (T=1T{=}1, k=40k{=}40). (a) Scatterplot of all extracted sequences $\bm{z}$} shown by their start position in the book. The $100$-token sequences overlap significantly. (b) Condensed heatmap view. At each character position, we plot the maximum $p_{\bm{z}}$ across all overlapping sequences whose suffix covers that position. In both plots, colors encode $p_{\bm{z}}$ on a $\log$ scale. We deem extraction successful if $p_{\bm{z}}\!\geq\!\tau_{\text{min}}\!=\!0.1\%$, considering values below this to not reflect extraction (i.e., rounding $p_{\bm{z}}$ down to $0\%$). See Appendix D.
To identify regions of memorization, we develop the following “panning for gold” approach. For a given book, we start at the beginning of the text file in Books3. We take a chunk of text that is sufficiently long to contain $100$ tokens, slide $10$ characters forward in the book text and repeat this process. We do this for the entire length of the book, which results in approximately $1$ sequence every $10$ characters, e.g., The Great Gatsby has 270,870270{,}870 characters and thus roughly 27,00027{,}000 sequences. For most of our experiments, we follow the standard procedure of evaluating $100$-token sequences [carlini2023quantifying, hayes2025measuringmemorizationlanguagemodels], which we divide into $50$-token prefix prompts and $50$-token suffixes that we attempt to extract (Equation 1).
As a result, every character in the book (except for at the beginning and end) is present in multiple sequences—in both prefixes and suffixes. This significant degree of overlap is deliberate. Since we typically do not know how open-weight models were trained, it is not exactly clear how we should break up books into token sequences when attempting extraction. By testing various sequence starts and ends in the same area, we aim to surface as much total memorization as possible. We are attempting to surface extraction “hot-spots” within a book (Appendix D.2).
In Figure 3, we provide visualizations of the results of this procedure for one setting: Llama 3.1 70B (trained on Books3) on 1984 [1984] (a book contained in Books3) using top-$k$ decoding (T=1T{=}1, k=40k{=}40). Figure 3(a) shows a scatterplot of every $100$-token sequence whose suffix we successfully extracted (i.e., $p_{\bm{z}}\!\geq\!\tau_{\text{min}}\!=\!0.1\%$). We plot each sequence’s extraction probability $p_{\bm{z}}$ according to its location (start character position) within the book. White vertical gaps at a given point on the $x$-axis indicate that a sequence with that start position was not extracted (i.e., has $p_{\bm{z}}\!=\!0\%$ ). A continuous horizontal band of color (associated with a given $p_{\bm{z}}$ on the $y$-axis) would indicate that we can extract the entire book—in pieces, with respect to $100$-token sequences—with that associated probability.
As alternate view, Figure 3(b) condenses the same results into a heatmap. Unlike the scatterplot, this heatmap does not convey how many sequences are extractable at different probabilities in each area of the book. Instead, at each character position, we plot the maximum extraction probability $p_{\bm{z}}$ across all overlapping sliding-window sequences whose suffixes cover that character position. As we show in Section 6, such heatmaps facilitate comparisons of memorization “hot-spots” across models at the same book location. This heatmap shows high extraction probabilities exist throughout the entirety of 1984. On its own, this does not mean that the whole book can be extracted in one continuous segment at generation time; these results show how much of the book we estimate to be memorized in the model’s parameters, with respect to probabilistic extraction (Equation 1) measured with $50$-token prefixes and suffixes. We return to reconstructing an entire book at generation time in Section 7.
4.2 Average extraction rates obscure book-specific memorization
Model Book % of book extracted with $p_{\bm{z}}\!\geq$
$\tau\!=\!1\%$ $\tau\!=\!10\%$ $\tau\!=\!50\%$ $\tau\!=\!75\%$
Llama 3.1 70B Harry Potter 90.89% 75.44% 43.26% 16.75%
Sandman Slim 0.38% 0.28% 0.12% 0.00%
Llama 1 65B Harry Potter 25.48% 15.00% 4.40% 1.69%
Sandman Slim 0.27% 0.13% 0.10% 0.09%
Pythia 12B Harry Potter 0.40% 0.10% 0.08% 0.00%
Sandman Slim 0.34% 0.32% 0.26% 0.22%
Figure 4: Average extraction rates are low, but book-specific extraction varies widely. (left) Comparing extraction rates (Equation 3) of random $100$-token sequences ($50$-token prefix ++ $50$-token suffix) from Books3 for different models trained on Books3. We show the greedy (blue) and probabilistic (orange, $p_{\bm{z}}\!\geq\!\tau_{\text{min}}\!=\!0.1\%$, top-$k$ decoding with T=1T{=}1 and k=40k{=}40) discoverable extraction rates (Appendix C). For both, average extraction is low. (right) Extraction coverage (Equation 4) for specific books reveals a more nuanced picture. For Harry Potter and the Sorcerer’s Stone [hp1] and Sandman Slim [sandmanslim], we show coverage for $\tau\in\{1\%,10\%,50\%,75\%\}$ for $3$ models (Appendix D.4). Nearly half of Harry Potter and the Sorcerer’s Stone can be extracted with respect $p_{\bm{z}}\!\geq\!\tau\!=\!50\%$.
Our sliding-window procedure provides a detailed, work-specific view of memorization that is distinct from prior work in ML. That is, while prior work shows that models memorize certain amounts of the data they were trained on [e.g., carlini2021extracting, nasr2023scalable, hayes2025measuringmemorizationlanguagemodels, lee2022dedup, biderman2023pythia, reid2024gemini, gemma2, biderman2023emergent], it typically quantifies memorization through overall extraction rates—much like those in Figure 4 (left). For some (part of a) training dataset, researchers draw (typically at random) a set $\mathbb{Z}$} of sequences $\bm{z}$} of a specified length (e.g., $100$ tokens). For each sequence, they prompt with its prefix and count extraction as successful if the resulting generation matches the target suffix (Section 2). The extraction rate is computed as the number of attempted extractions that succeeded, relative to the total number of attempts:
$\mathsf{extraction\_rate}({\mathbb{Z}};s)\;=\;\frac{1}{|{\mathbb{Z}}|}\sum_{{\bm{z}}\in{\mathbb{Z}}}\mathbf{1}[\,s({\bm{z}})\,],$ (3)
where $s({\bm{z}})$ is an extraction success condition. As noted in Section 2.1, for traditional discoverable extraction, the indicator in Equation 3 evaluates to $1$ when the greedy completion exactly matches the target suffix. In probabilistic extraction, the indicator evaluates to $1$ if the target suffix’s probability $p_{\bm{z}}$ exceeds a chosen threshold $\tau_{\text{min}}$ (Equation 2, Appendix C).
These reported averages are generally small, just as we observe in Figure 4 (left) for $9$ models trained on Books3. The highest rate we observe on random $100$-token sequences ($50$-token prefix ++ $50$-token suffix) from Books3 is just over $0.6\%$ (for Llama 3.1 70B). Small average extraction rates often form the basis of defendants in copyright infringement suits calling memorization a rare “bug” [cooper2024files, openairesponse]. However, while average extraction rates are useful for estimating overall memorization of Books3 text, they clearly tell an incomplete story. Low extraction rates signal that models likely do not memorize most text in Books3, but specific works can be highly memorized. On average, Llama 3.1 70B may have only memorized about $0.6\%$ of sequences in Books3, but concealed within that small number is that it memorized effectively all of 1984 (Figure 3).
We can see this intuitively in Figure 3; target suffixes whose extraction probability exceeds $\tau_{\text{min}}{=}0.1\%$ can be found nearly everywhere in the book. [^5] Exceptions are this particular edition’s forward at the beginning of the book and notes near the end of the book—located before some selections of other writing by George Orwell, which are also highly memorized. However, we can also make this more concrete: using the results of our sliding-window procedure, we can quantify how much of a whole book is extractable. We compute extraction coverage as the fraction of characters in a book that lie within at least one suffix whose extraction probability $p_{\bm{z}}$ exceeds a threshold $\tau$ (Appendix D.4)—where $\tau\in[\tau_{\text{min}},1]$ because $\tau_{\text{min}}$ is the minimum probability we consider to reflect extraction success (Sections 2.1 & 5). Let $\mathsf{span}({\bm{z}})$ be the set of character indices spanned by a given $50$-token suffix of sequence $\bm{z}$}. Then
[…]
Figure 5: Extraction coverage (Equation 4) differs across models for the $50$ books we evaluate, illustrated for Llama 1 65B and Llama 3.1 70B. Results use $100$-token sequences ($50$-token prefix ++ $50$-token suffix), top-$k$ decoding (T=1T{=}1, k=40k{=}40), and coverage threshold $p_{\bm{z}}\!\geq\!\tau\!=\!10\%$.
For Llama 3.1 70B and 1984, Figure 3(a) highlights in red the sequences used to compute Equation 4 at $p_{\bm{z}}\!\geq\!\tau\!=\!25\%$. These sequences cover $54.39\%$ of 1984: for each, reproducing the verbatim $50$-token suffix would require (in expectation) only $4$ prompts with the corresponding $50$-token prefix.
The majority of the $50$ books and $17$ models we test do not have this degree of extraction coverage, as is illustrated in Figure 5 for Llama 1 65B and Llama 3.1 70B at $\tau\!=\!10\%$. For both models, most books at this more permissive threshold exhibit $2\%$ coverage. There are nevertheless important exceptions, like 1984. Several books have $10\%$ coverage, and Llama 3.1 70B exhibits higher coverage on more books—more than any other model we test (Appendix D.4). [^6] To have an accurate, general sense of what extraction coverage distributions look like for different models, we would need to run a much larger sample than $50$ of the ${\sim}200{,}000$ books. We defer this to future work.
For a more detailed comparison of book-specific coverage, in Figure 4 (right) we show extraction coverage for Harry Potter and the Sorcerer’s Stone [hp1] and Sandman Slim [sandmanslim]—one of the books by plaintiff Richard Kadrey in Kadrey et al. v. Meta, Inc. [kadreyamendedconsolidated]). Sandman Slim is hardly memorized at all. Extraction coverage for this book is less than $0.4\%$ for all models and thresholds $\tau$. [^7] Prior studies (and Figure 4) show that larger LLMs exhibit higher extraction rates [carlini2023quantifying]. This pattern does not cleanly generalize to specific works. At $\tau{=}75\%$, Llama 3.1 70B exhibits $0.00\%$ extraction coverage of Sandman Slim, while the smaller Llama 1 65B and Pythia 12B exhibit $0.09\%$ and $0.22\%$, respectively. In contrast, the first Harry Potter [hp1] is highly memorized in Llama models (Section 6). For Llama 3.1 70B, $43.26\%$ of the book is covered with $\tau\!=\!50\%$. [^8] The version we evaluate from Book3 includes the first chapter of the second Harry Potter book at the end. For lower $\tau$, there is higher coverage; $90.89\%$ of the book is covered with $\tau\!=\!1\%$. While lower thresholds correspond to extracting text less reliably—i.e., requiring more prompts to the model—for an LLM, they reflect enormous probabilities (Section 2.1).
5 Validating our extraction procedure
In prior sections (Section 24), we emphasized that the minimum probability we count as successful extraction, $\tau_{\text{min}}$ (Equation 2), is deliberately conservative. But how can we be confident this is the case: that when we measure a successful extraction, we are capturing a true instance of memorization, rather than happenstance generation of a verbatim suffix (a “monkey at the typewriter” event, Section 3)? We address this question through a set of experiments designed to validate our approach. As before, we apply our sliding-window procedure with top-$k$ decoding (T=1T{=}1, k=40k{=}40) and $50$-token suffixes (Section 4.1), but here we also vary prefix length (up to $800$ tokens) to observe how conditioning on additional context affects our measurements.
Figure 6: Evaluating Llama 3.1 70B, Llama 3.1 8B, and Phi 4 on We Were Eight Years in Power [wewere] (a book in Books3), using prefix lengths in $\{25,50,100,200,400,800\}$ and $50$-token suffixes with top-$k$ decoding (T=1T{=}1, k=40k{=}40). The book is in Llama’s training data, but not Phi 4’s. Longer prefixes reveal more memorized sequences, though not all sequences in the book are extractable. For Phi 4, the only extractable suffixes are widely duplicated quotes that likely appeared in its training data from other sources. These results support that when our procedure identifies extraction, it is detecting true instances of memorization rather than chance generations.
As a baseline, we run experiments on a book from Books3 and Llama 3.1 models, which were trained on Books3 (Section 5.1). As expected [carlini2023quantifying], longer prefixes surface more memorized training data. Then, we run a set of negative controls: we use the same procedure for models on non-training data, which tests whether our method spuriously registers false positives on data where memorization is impossible (Section 5.2). Unlike the experiments on training data, regardless of prefix length, suffix probabilities remain below $\tau_{\text{min}}$ (with interesting exceptions [^9] These sequences are boilerplate text (copyright notices, publisher addresses, etc.) and popular quotes—both of which are duplicated across many sources aside from Books3. (See Appendices D.5, E.6, & G.) ), demonstrating our approach does not mistakenly register non-training data as extractable. Together, these experiments show that, when our measurements indicate extraction, we can be confident they are a valid reflection of memorization. We briefly discuss the implications of these results (Section 5.3), which inform the experiments we run in the following section. Additional results and discussion can be found in Appendix E.
5.1 Longer prefixes surface more memorized training data
When a sequence is known to be in the training data, memorization is the most plausible explanation for high-probability generation of its (sufficiently long) verbatim suffix given the corresponding prefix (Section 2). So, as a baseline, we apply the sliding-window procedure to this setting (Figure 6): we attempt to extract sequences in We Were Eight Years in Power [wewere] (a book in Books3) from Llama 3.1 70B and Llama 3.1 8B (models trained on Books3). Longer prefixes reveal additional memorized sequences. This is what carlini2023quantifying call the discoverability phenomenon: providing additional context (a longer ground-truth prefix) can raise the conditional probability of the target suffix $p_{\bm{z}}$ enough that it becomes extractable ($p_{\bm{z}}\!\geq\!\tau_{\text{min}}$). However, many sequences remain unextractable: no matter how long the prefix, their suffix probabilities $p_{\bm{z}}$ never surpass $\tau_{\text{min}}$. [^10] Our results suggest that longer prefixes beyond $800$ tokens would enable extraction of additional sequences from this book, but the gains diminish, indicating that not every sequence is extractable (Appendix E.3).
For example, Llama 3.1 70B memorizes most of the book, yet consistent sections remain unextractable across all prefix lengths (e.g., white gaps in Figure 6, such as from ${\sim}75$K–120120K characters). The same is even clearer for Llama 3.1 8B: being smaller, it memorizes significantly less [carlini2023quantifying, hayes2025measuringmemorizationlanguagemodels, carlini2021extracting]. Longer prefixes do surface more memorized training data, but most of the book remains unextractable. These results are consistent with prior work: high-quality models do not—indeed cannot [lee2022dedup, carlini2021extracting, mahdavi2024memorizationcapacitymultiheadattention, collins2017capacitytrainabilityrecurrentneural]—memorize all of the training data in their massive training datasets. Even as prefix length increases, our measurements do not suddenly capture extraction success for everything; they capture success for sequences that have been memorized. [^11] We observe the same for Llama 3.1 70B on other books in its training data. Some appear effectively not memorized at all, with only limited increases in extraction success at longer prefix lengths (Appendix E.3). We further support this point in the next section, where we compare these results to experiments on non-training data.
5.2 Measurements on non-training data do not mistakenly register extraction
Using the above results as a basis for comparison, we next test our method in settings where memorization is impossible. This is the role of negative controls: experiments in conditions where the outcome of interest should not occur. If (false) positives appear in such a setting, they indicate that the measurement procedure is capturing spurious signal. In our context, running the sliding-window procedure on non-training data provides this sanity check. In general (with important exceptions noted below), observing suffix probabilities above $\tau_{\min}$ would indicate false positives, i.e., would count successful extractions where none should exist. The absence of such cases is evidence that the extractions we do count are valid reflections of memorization.
We run two sets of negative controls. First, we apply our procedure to Books3 books with Phi 4 [phi4]—a model that was not trained on whole copyrighted books. Second, we run Llama 3.1 70B (trained on Books3), on books from 2025, i.e., books that post-date its 2024 training cutoff. In our plots, we distinguish these conditions by color: blue for models trained on Books3 on Books3 text, orange for Phi 4 on Books3 text, and gray for results on post-cutoff books.
Negative control 1: A model not trained on Books3 $\times$ Books3 text.  We run the same experiment as in Section 5.1 on Phi 4, which was not trained on in-copyright books like We Were Eight Years in Power [wewere]. To support the validity of our measurement approach, unlike with Llama 3.1 models, we should not observe extraction success for unique text from this book. Our results for Phi 4 in Figure 6 confirm this expectation: increasing the context length does not push the conditional probability of non-training-data suffixes high enough to cross the $\tau_{\text{min}}$ threshold. We do observe increased extraction signal, but it only concentrates in four small areas that do not appear to be false positives. Instead, they plausibly reflect memorized text: the suffixes our procedure counts as extracted are famous quotes that are widely duplicated on the Internet, and thus very likely to have been included in Phi 4’s training data via sources other than this copyrighted book [phi4]. For example, at ${\sim}280$K characters, we extract sequences containing passages from the King James Bible [deuteronomy]; at ${\sim}546$K characters, we extract a quote from Barack Obama’s 2004 Democratic National Convention speech [obama2004]. Notably, Llama 3.1 models also exhibit high extraction probabilities in these same regions. While they were trained on this book, it is plausible that duplicates from other sources contributed to the memorization and extractability of these sequences [lee2022dedup].
Figure 7: Validating probabilistic extraction by computing $p_{\bm{z}}$ for non-training data: Llama 3.1 70B and top-$k$ (T=1T{=}1, k=40k{=}40) decoding on Great Big Beautiful Life [greatbig] (published spring 2025), with respect to $50$-token suffixes and prefix lengths in $\{200,400,800\}$ tokens. Except for the copyright notice at the beginning, we do not register any $p_{\bm{z}}\!\geq\!\tau_{\text{min}}\!=\!0.1\%$. We register no other suffix probabilities $p_{\bm{z}}\!\geq\!10^{-13}\%$, and only 11 (out of ${\sim}62{,}000$ sequences) for which $p_{\bm{z}}{\geq}10^{-14}\%$.
Negative control 2: Models trained on Books3 $\times$ books published after their training.  We apply the same procedure to Llama models on books published in 2025—after the July 2024 cutoff for Llama 3.1 training. In Figure 7, we show one example for Great Big Beautiful Life [greatbig]. Apart from the (highly duplicated, boilerplate) copyright notice, for all prefix lengths there are no suffix probabilities that, in analogous experiments on training data, we would count as successful extraction. Not only does $p_{\bm{z}}$ never surpass $\tau_{\text{min}}\!=\!0.1\%\!=\!10^{-1}\%$, there are no sequences with $p_{\bm{z}}\geq 10^{-13}\%$. Exactly $1$ sequence (out of ${\sim}62{,}000$) has $p_{\bm{z}}\geq 10^{-14}\%$ (the faint gray band at ${\sim}20$K characters). This reflects chance generation of non-training data—effectively a “monkey at the typewriter” event. But importantly, this is not a sequence that our method counts as extractable: the suffix probability is $13$ orders of magnitude smaller than our setting for $\tau_{\text{min}}$. With a single attempt, the odds of generating the verbatim suffix are $\nicefrac{{1}}{{p_{\bm{z}}}}\!\approx\!\nicefrac{{1}}{{10^{-14}\%}}$, or $1$ in $10$ quadrillion (Appendix E.4.2). To give a sense of just how small this number is, the chance of winning the Powerball jackpot in a single draw is about $1$ in $292$ million—roughly $34$ million times more likely.
5.3 Implications for measuring probabilistic extraction on books
Three important observations follow from these results (and related experiments in Appendix E.4); we highlight them here and defer additional discussion to Appendix E. First, the negative controls provide strong evidence that our sliding-window probabilistic extraction procedure is valid. Second, they indicate that our choice of $\tau_{\text{min}}$ for counting extraction success is indeed very conservative. Given the very low probabilities we observe for generating non-training data (Figure 7, Appendix E.4.2), we could reasonably set $\tau_{\text{min}}$ below $0.1\%$. As such, our results almost certainly under-count memorization, and should be read as a lower bound. Nevertheless, we prefer to err on the side of caution. To lower $\tau_{\text{min}}$ and remain confident about validity, we would need to run a much larger set of negative controls, which exceeds the compute budget for this work. Part of the reason is that, as others have noted [cooper2024files], LLMs are not actually like monkeys randomly outputting tokens. They have learned to generate structured, grammatical sentences, so the distribution of plausible outputs is far smaller than random chance would suggest (Appendix E.5). Exploring what this implies for how high $p_{\bm{z}}$ must be to count as evidence of memorization is an important question for future work. For now, we prefer to under-count memorization rather than risk over-counting it.
Third, having validated our procedure, we can now apply it more broadly. In particular, we are justified in extending our experiments beyond models like Llama, where we know that Books3 was part of the training data. If suffix probabilities for Books3 sequences surpass $\tau_{\text{min}}$, it is reasonable to interpret these as true extractions and thus evidence of memorization. This remains the case regardless of whether the model was literally trained on Books3, or whether the memorized training data originate from other overlapping sources—for example, highly duplicated famous quotations, as we saw with Phi 4 in Figure 6 (Appendix E.6). On this basis, in the following section, we expand our analysis to include $3$ additional models for which it is unknown if Books3 was included in the training data.
6 Comparing memorization across books and models
We now present the main results of applying the sliding-window method. We describe the experimental setup (Section 6.1) and highlight the types of comparisons our approach facilitates across both books and models (Section 6.2). We defer discussion of the copyright implications of these results to Section 8.
6.1 Experimental setup
We use the same setup as in Section 4.1: the sliding-window procedure with top-$k$ decoding (T=1T{=}1, k=40k{=}40) for $100$-token sequences ($50$-token prefixes ++ $50$-token suffixes). We run this procedure on $17$ models, which we organize into three groups: (a) models trained on Books3 (Llama [touvron2023llamaopenefficientfoundation, llama2, llama3] and Pythia [biderman2023pythia], blue); (b) models that were very likely trained on (at least parts of) books also contained in Books3, based on measurements of $p_{\bm{z}}\!\geq\!\tau_{\text{min}}\!=\!0.1\%$ for unique content from those books (DeepSeek v1, Qwen 2.5, and Gemma 2, red); and (c) a model not trained on copyrighted books (Phi 4, orange). Of the ${\sim}200{,}000$ books in Books3, we include $50$ in our analysis. These are the same books for which we analyze extraction coverage in Figure 5. They include works associated with plaintiffs in the amended class action complaint of Kadrey et al. v. Meta, Inc. [kadreyamendedconsolidated, pp. 4-5], [^12] That suit was recently decided in favor of the defendants (Meta) [kadreyjudgment]. as well as generally popular books [e.g., hp1, hobbit, sisyphus, catch22]), and some academic texts [e.g., zittrain, dante]). We suspected that some popular books would exhibit higher degrees of memorization (e.g., due to duplicated text from other sources). This is why we did not select books at random; we deliberately chose titles intended to capture variation across Books3. Further details on these decisions are in Appendix D.1. We ran thousands of experiments on these $50$ books and $17$ models (Appendix D.3). (In total, these experiments cost $6$ GPU months.) Nevertheless, our results should not be read as a complete account of memorization across the entire Books3 dataset.
6.2 Key patterns in our results
We focus here on only a few representative books and on the first two categories of models—(a) models known to be trained on Books3 and (b) models where it is not disclosed whether Books3 was included in the training data. Recall that extraction probability $p_{\bm{z}}$ reflects the expected frequency that a model would reproduce the verbatim $50$-token suffix given its prefix (Section 2). Because of this, $p_{\bm{z}}$ values can be compared across models, allowing us to evaluate which models are more or less at risk of outputting verbatim $50$-token suffixes. [^13] Different LLMs have different token vocabularies and tokenizers, which segment text differently. So, we mean this in the high-level sense of comparing heatmap regions at the same within-book character locations. With this metric, we draw conclusions about different degrees of memorization: how memorization varies for a specific book across models and across books for a given model. We omit heatmaps for Llama 3 70B, as they resemble those for Llama 3.1 70B.
(a) Models trained on Books3
(b) Models that we conclude were trained on (at least parts of) some books that are also contained in Books3
Figure 8: Heatmaps clarify how memorization varies widely across books and models. We apply the sliding-window procedure with top-$k$ decoding (T=1T{=1}, k=40k{=}40) and $100$-token sequences ($50$-token prefixes ++ $50$-token suffixes). We show (a) models known to be trained on Books3 (blue) and (b) models for which it is unknown if Books3 was included in the training data (red). With respect to these settings, Llama 3.1 70B memorizes effectively all of Harry Potter and the Sorcerer’s Stone.
Comparing a single book across models.  Each book in Figure 8 illustrates a different high-level pattern of memorization across models. There are books like M. Butterfly [mbutterfly], for which our sliding-window procedure reveals almost no memorization: aside from the copyright notice, we extract little—if any—text from these books across models of different sizes trained by different organizations. Most books we test with $50$-token prefixes resemble this pattern. However, for a given book, we are often able to extract some substantive sequences from Llama 3 and 3.1 70B.
Books like Beloved [beloved] represent a more intermediate case: for these books, memorization is generally low for most models with respect to $50$-token prefixes (Appendix E.3), but with important exceptions. Llama 3 and 3.1 70B memorize a considerable number of fragments in the book. Other Llama models memorize significantly less, and category (b) models (DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B) memorize even less. For Beloved, the $7$ largest models we test—category (b) and all large Llama models—memorize the same passages at ${\sim}170$K [belovedquote1, belovedquote3] and ${\sim}480$K characters [belovedquote2]. Both are famous quotes from the book, which we confirmed are highly duplicated on the Internet. As with Phi 4 (Section 5.2), models that otherwise exhibit low memorization (except for such passages) may not have been trained on the whole book; rather, they may have memorized these passages due to their inclusion and duplication in other sources (Appendix E.6).
Last, select books exhibit enormous degrees of memorization for the $7$ largest models, with lesser amounts of memorization in smaller models like Llama 1 13B. These are widely popular books like Harry Potter and the Sorcerer’s Stone [hp1] (Figure 8) and 1984 [1984] (Figure 3). While memorization of such books tends to be significant for larger models, it is especially high for Llama 3 and 3.1 70B, which appear to memorize several books in their entirety [e.g., hp4, gatsby, ulysses, hobbit].
Comparing different books within a model.  Small models generally memorize less than larger ones, but there is variation across books. For example, Llama 1 13B memorizes more of Harry Potter and the Sorcerer’s Stone than M. Butterfly and Beloved; Llama 3.1 8B memorizes portions of We Were Eight Years in Power [wewere] (Figure 6), but less from other books (Appendix E.3). Variation is more pronounced in larger models. As noted above, for all models, most results resemble those shown in Figure 8 for M. Butterfly. This is true for most of the plaintiff-authored books in Kadrey et al. v. Meta, Inc. [kadreyamendedconsolidated]—with exceptions (e.g., Llama 3.1 70B on wewere; see Figure 6). Extraction coverage distributions reinforce this conclusion: they skew toward low coverage (Section 4.2, Appendix D.4). However, there are clearly several exceptions. Llama 3 and Llama 3.1 70B memorize significantly more from the $50$ books than any other model we tested. Many of these books yield results similar to those for Beloved [e.g., handmaid, davinci, sisyphus, curious]. In extreme cases, both models appear to memorize entire books, such as Harry Potter and the Sorcerer’s Stone. We address this further in Section 7. Llama 1 65B, Llama 2 70B, and category (b) models also exhibit high memorization for some, but not all, of these books.
Altogether, these patterns underscore the importance of training choices in shaping memorization. Models do not memorize all of their training data [lee2022dedup, carlini2021extracting, mahdavi2024memorizationcapacitymultiheadattention, collins2017capacitytrainabilityrecurrentneural]. The mere existence of particular sequences within a training dataset does not imply a particular conclusion about memorization [lee2022dedup, lee2023explainers, lee2023talkin, lee2024talkinshort]. For example, similarly sized models trained by different organizations display different degrees of memorization of the same book—even when the evidence strongly suggests the entire book (e.g., Harry Potter and the Sorcerer’s Stone) was included in their respective training datasets (Appendix E.6). This is also evident with respect to just Llama models, which were all trained by the same organization (Meta) on Books3. All Llama models (by size) memorize more than the other models we evaluate. But notably, within size classes, newer versions typically memorize more from specific books than earlier ones: Llama 3 models appear to have memorized more from books than Llama 2, which in turn memorize more than Llama 1.
7 Reconstructing Harry Potter and the Sorcerer’s Stone with one seed prompt
Our sliding-window probabilistic extraction experiments estimate how much of a given book is memorized within an LLM’s weights, with respect to $100$-token sequences ($50$-token prefixes ++ $50$-token suffixes). For a $50$-token suffix where $p_{\bm{z}}=1\%$, it would take (in expectation) $100$ prompts with the prefix to extract the verbatim suffix (Section 2.2). Therefore, these results do not imply that one would definitively be able to extract large, continuous portions of a book all at once during generation.
However, for Llama 3 70B and 3.1 70B on some books (Figures 3 & 8), our results show that there exist extraordinarily high extraction probabilities throughout. In fact, these probabilities approach $100\%$ when we increase the prefix length. For example, for Harry Potter and the Sorcerer’s Stone, increasing the prefix to $200$ tokens pushes $p_{\bm{z}}$ (for $50$-token suffixes) to over $90\%$ almost everywhere in the book (Appendix E.3). This means we can extract almost any $50$-token segment with approximately $1$ prompt, using the corresponding $200$-token prefix. With this degree of memorization across the entire book, one could possibly reconstruct the entire book near-verbatim, starting from a single short prompt of ground-truth text to seed the generation process. Using beam search to decode generations—which deterministically approximates the highest probability sequence under the model for a given configuration—the model could reasonably continue the seed prompt to complete the book.
7.1 A simple reconstruction procedure
As a simple test of this idea, we implement this reconstruction procedure with about $100$ lines of boilerplate generation code using HuggingFace APIs for Llama 3.1 70B. We begin with an $s$-token seed prompt (e.g., s=60s{=}60, the first line of Harry Potter and the Sorcerer’s Stone), and set a maximum sliding context length of $n$ tokens. The model generates $m$ tokens at a time using beam search, until reaching the context window length $n$. At this point, we remove the first $m$ tokens from the context and append the generated $m$ tokens, creating a new $n$-token prompt. This updated prompt becomes the input for the next generation step. We repeat this process, sliding along the length of the book by prompting with $n$ tokens to produce the next $m$ tokens. After the first $n$ tokens, the clipped, $n$-length context window progresses beyond the initial seed prompt. All subsequent prompts do not contain any ground-truth text drawn from the book; they consist entirely of generated text from prior iterations. In essence, this reconstruction procedure is just standard autoregressive generation. The only difference is that we clip the context to a fixed window length nn—rather than letting it grow by $m$ tokens every iteration—by discarding the oldest tokens as new ones are generated.
With appropriate choices of $s$, $n$, $m$, and configuration of beam search, [^14] Beam search is a deterministic decoding method that explores multiple candidate sequences in parallel to approximate the highest-probability continuation under the model. At each generation step, it keeps the top-$k$ highest-probability sequences so far. At the final step, it outputs the sequence with the highest probability. we attempt to generate the entire book with high fidelity. On our hardware, after testing just a handful of configurations for beam search and a few different short seed prompts, we were able to deterministically generate a near-exact copy of the entire 304-page book. We show results in this section using a seed prompt of just "Mr. and Mrs. D" (s=6s{=}6 tokens, drawn from the first line of the first chapter, prepended with the beginning-of-sequence token) and a beam width of k=8k{=}8; we use a sliding context length of n=3000n{=}3000 tokens and generate m=50m{=}50 tokens at each step. In practice, the only minor complication was that Llama 3.1 70B tended to predict end-of-sequence (EOS) tokens at the ends of chapters. To address this, we removed the EOS token from the generated $m$ tokens before appending them to the context, and replaced it with the tokens for "CHAPTER {c+1}" (with {c+1} spelled out in words, e.g., when c=2c{=}2, we insert "CHAPTER THREE"). With this simple adjustment, the model successfully continued to generate the book.
In summary, it was straightforward to produce a copy of Harry Potter and the Sorcerer’s Stone that is nearly identical to the original. Further, because beam search is a deterministic decoding algorithm, our results are reproducible. We provide the complete (very short) code, and discuss the stability of this result and future work beyond this “existence proof” in Appendix F.
Figure 9: Samples of the diff between the ground-truth text of Harry Potter and the Sorcerer’s Stone [hp1] from Books3 and the entire book text we generated using Llama 3.1 70B, starting with a short seed prompt from the first chapter. See Appendix F.1.
7.2 Assessing the fidelity of the reconstructed book
Qualitatively, the two documents are close to identical. As illustrated in Figure 9, differences predominantly involve small inconsistencies in formatting: white space, capitalization, use of underscores (_) to indicate italics, etc. The version of Harry Potter and the Sorcerer’s Stone in Books3 uses British spelling (e.g., "Mum" instead of "Mom"), so there are also some minor spelling differences. Very occasionally, when there is only a single line in a paragraph, the model skips that line during generation.
We also quantitatively assess the similarity of the two texts, using metrics for which two identical documents have a score of $1$. The cosine similarity of the two documents’ TF-IDF vectors is a near-perfect $0.9999$. TF-IDF is a limited metric, as it treats documents as bags of words and thus fails to capture word order. So, as two additional points of comparison, we compute similarity using greedy longest common subsequence matching at both the word and sentence levels (via difflib’s SequenceMatcher). We obtain a word-level similarity of $0.992$ and a sentence-level similarity (which is more sensitive to formatting differences) of 0.9340.934.151515We apply minimal text normalization before computing these values: removing _ (used to signify italics in the Books3 version) and aligning ellipses to "..." (which appear as " . . ." in Books3).
8 Takeaways for copyright law and policy
Our experiments show that the extent of memorization in models varies by model, model family, model size, the book tested, and even within different parts of an individual book (Section 46). These results complicate the narratives both plaintiffs and defendants typically present in copyright cases when describing how LLMs work. They support the positions of plaintiffs in some respects and of defendants in others. We see three primary implications for copyright disputes: LLMs could be derivative works of the books they memorize (Section 8.1); in some cases, it may be practical to generate infringing copies (Section 8.2); and, class actions are ill-suited for copyright claims about memorization (Section 8.3).
8.1 LLMs could be derivative works of the books they memorize
In the case of some models, there is quite a lot of memorization of some books, though most books are not memorized at all—with respect to our specific extraction methodology—either in whole or in part. Such evidence of memorization in the model matters for the dispute over whether models themselves are derivative works (Section 3), an argument that courts have thus far not been receptive to [bartzjudgment, kadreyjudgment].
A work is not a derivative work unless it is “substantially similar” in significant part to the original work [e.g., litchfield]. While prior debates over whether models can qualify as derivative works have been largely theoretical [cooper2024files, lee2023talkin], the results of our probabilistic extraction procedure offer empirical evidence that is directly relevant to this issue. They indicate that models are not, as plaintiffs sometimes contend, mere copies of all of the works on which they were trained. However, they also suggest that at least some LLMs (e.g., Llama 3.1 70B) may be derivative works of at least some books (e.g., Harry Potter and the Sorcerer’s Stone)—because those LLMs have memorized a significant amount of protectable expression from those books. The law does not require that the entire work be included in the derivative; it is enough that the derivative incorporates a substantial amount of protectable expression. We find that this turns out to be true for some (but not all) books for some (but not all) models.
Fair use.  The inquiry does not end here; the model, like the training dataset, may be protected under copyright’s fair use doctrine [lemley2021fairlearning]. But the fair use analysis of the model itself may be different than the fair use analysis of the training dataset [lee2023talkin]. Copyright law treats intermediate copies, such as those used internally within a system, differently from external-facing copies. A sequence of copyrighted data used for training is an intermediate copy used internally by AI companies for model training, rather than sold as a product. Unless that sequence is extracted, it is only used internally in the course of the model producing outputs. Beyond extraction, individual sequences may to lesser or greater extents influence model outputs, but many of these outputs will not be substantially similar to any training data [henderson2023fair, lee2023talkin].
A model, too, is an intermediate work in many cases. If the model memorizes content but only generates noninfringing outputs, that is a strong argument that how the model makes use of that content is transformative and ought to be permissible fair use. But for companies that sell or release their models to others under open source licenses, the model itself is the product, and sometimes one that is being sold for commercial gain. That does not defeat the fair use argument, but it makes the analogy to training data less precise. In these instances, it may be harder to rely on the cases that justify training on copyrighted data as fair use [e.g., authorsguild, sega, sonyconnectix].
Potential consequences of a model being a copy of works it has memorized.  A finding that the model itself is a copy of some copyrighted works, and that a distribution of the model was thus a distribution of a copy of those works, could be dramatic for the AI industry. Llama 3.1 70B was downloaded 105,029105{,}029 times from HuggingFace in May 2025 (Appendix G)—far less from the height of its release in 2024. If we say conservatively that the model has been downloaded 1 million times since its release, [^16] Meta announced in March 2025 that total downloads for all of its Llama models surpassed $1$ billion [metadownloads]. In 2024, the Times reported that Llama 2 models had been downloaded $180$ million times [nytllama2]. then those 1 million downloads could be seen as 1 million potentially infringing distributions of reproductions of Harry Potter and the Sorcerer’s Stone—as well as any other books in copyright for which a more than de minimis amount of copyrighted expression has been memorized.
In general, damages for copyright infringement can run very high: statutory damages can be as high as \$150,000 per infringed work, and actual damages can be much higher. If the model itself is a copy and is not protected by fair use, courts might also order that it not be distributed or even that copies of the model itself be destroyed [17usc503]. We have only studied a relatively small number of books ($50$) and models ($17$), so cannot draw strong conclusions about the extent of memorization of books in general. In principle, if the overall quantity of memorized books in open-weight models is particularly high, it is not an exaggeration to say that the risk of damages across all of those books and of orders of destruction are an existential threat to the companies training these models.
Policy considerations.  The effect of these copyright rules may be to treat openly distributed models less favorably than models gated behind closed system, because the distributed models contain a latent copy or derivative of some copyrighted works. However, there is no good policy reason to treat open-weight models more harshly in copyright law than proprietary ones. Policymakers may want to act to protect models from being declared unlawful copies, in order to prevent crushing financial liability on developers for providing a model that generates transformative, noninfringing outputs. However, a blanket protection for all open-weight models may also be unwarranted. The extent to which models memorize their training data is significantly influenced by choices during training, such as de-duplicating the training dataset [cooper2024files, carlini2023quantifying, lee2023talkin, lee2022dedup, hayes2025measuringmemorizationlanguagemodels] Policymakers may be less inclined to protect models if the developers made training choices known to substantially increase the risk of memorization—for example, deliberately duplicating particular books.
8.2 It may be practical to generate infringing copies in some cases
In our experiments with the sliding-window probabilistic extraction procedure (Sections 4 & 6), we use relatively short $50$-token prompts drawn from books to see if it is possible for an LLM to generate verbatim the following $50$ tokens with high probability. As is standard in ML, we use extraction of text snippets for evidence of memorization inside the model. But no one would realistically use this technique to extract an entire book. Doing so would require access to at least half of the book to use as $50$-token prefixes in order to even attempt to produce the other half in $50$-token suffixes. And, for sequences that are memorized, it might take in expectation up to $\nicefrac{{1}}{{\tau_{\text{min}}}}\!=\!1{,}000$ tries to produce each verbatim suffix, with no indication of which of the many responses represents the actual text of the book.
Traditional extraction measurements vs. reconstruction of (whole) books.  Reliable extraction of large segments of text would require other approaches, like the reconstruction procedure that we apply to Harry Potter and the Sorcerer’s Stone with Llama 3.1 70B (Section 7). This result is an extreme case and is not representative of most of the other books we have tested (Appendix F.3). Nevertheless, it is significant: in contrast to claims by some defendants [concordanthropic], it illustrates that, for LLMs that exhibit high degrees of memorization of specific works, it may be possible to generate copies or derivative works of those works. General assessments of the extent of this type of behavior, and how easy it is to elicit it, require more detailed exploration.
Practicality of extraction and responsibility for infringement.  Even though our results suggest it may be possible to reconstruct large segments of copyrighted text from an LLM, they do not imply that it will always be easy to do so. This may affect who courts view as directly liable for any output infringement [lee2023talkin]. As a policy matter, the case for liability against the model developer is stronger if simple prompts generate infringing outputs. The case is weaker if a user (e.g., a plaintiffs’ lawyer) employs more intensive strategies to try to generate infringing material. In that case, the adversarial user—who perhaps runs hundreds or thousands of prompts to try to get one output that is infringing—seems more like a direct volitional actor misusing the model [costar]; they should likely bear some responsibility for the infringement. [^17] Models are used at scale—in some cases, by millions of users. Hypothetically, a million users interacting with a particular model might independently prompt the model with the prefix of a memorized sequence. Even for a verbatim suffix that (in expectation) appears only once for every 1,0001{,}000 prompts (i.e., $p_{\bm{z}}{=}\tau_{\text{min}}{=}0.1\%$), in expectation 1,0001{,}000 of these users could expect to see the verbatim memorized suffix in the output. By contrast, in several cases (e.g., the first Harry Potter book, 1984), it does not take a thousand or even hundreds of generations to get large amounts of text; it takes only a handful. And in one case that we have identified so far (Harry Potter and the Sorcerer’s Stone), it is possible to generate the book directly from a short starting prompt and boilerplate generation code (Section 7). We will address this further in future work.
Considerations for LLM training and infringement.  The fact that extraction is probabilistic may affect resolution of the “is training infringement” issue. However, it is not determinative. Using a work as training data does not mean that it will necessarily be memorized. And, even if it is memorized, it may still be difficult to extract that work. Large parts of a work would need to be extractable in order to create superseding copies (i.e., a copy that can stand in for/ replace the work). Based on our current results, this seems unlikely (though possible, see Section 7) in practice for most works.
Nonetheless, the fact that memorization occurs at all creates a point of distinction from Authors Guild v. Google [authorsguild]. In this case, the court held that Google’s scanning of books to create a searchable database, and Google’s public display of limited text snippets on Google Books, was fair use. The court also considered and rejected an argument that hackers could gain unauthorized access to Google’s internal books database that supported the service, and thereby obtain plaintiffs’ entire works [804fd, at 227-28]. In contrast, our results show that it is possible to recreate some of the content of some books by “hacking” the model itself. We think it unlikely as a practical matter that people will use the model in this way; there are easier and more effective ways to pirate a book. But at a minimum, it complicates the copyright fair use analysis.
8.3 Class actions are ill-suited for copyright claims about memorization
To be certified as a class action, plaintiffs generally have to show that all putative class members share common legal and factual issues, as well as suffer common injuries. Because many of the pending cases are proceeding as class actions, plaintiffs will therefore need to demonstrate that all book owners meet this standard so that a court can certify the class and treat their claims together [class].
As is clear from our results, this may be impossible. Basic questions about whether a particular model actually incorporated any significant expression from any individual plaintiff’s book cannot be generalized. Rather, it varies case by case: the extent of memorization differs widely from model to model, and even from work to work within a model (Section 6). This makes it difficult to assess, on a class-wide basis, whether a particular model copied a particular work and whether, for that model, infringing outputs based on memorization are even possible. Indeed, we show that memorization of works varies for many of the actual named authors in lawsuits, and even across books for the same plaintiffs (e.g., Ta-Nehisi Coates, see Figure 6; Appendices D.3.6D.3.8, & E.3). Answering questions about the degree to which a specific LLM has memorized a specific work requires experimental evidence—for example, running every book through tests similar to the ones we performed here. Some plaintiffs may be able to show copying, but others will not. Courts generally deny class certification in such circumstances [e.g., walmart].
9 Conclusion
Building on recent work on probabilistic discoverable extraction [hayes2025measuringmemorizationlanguagemodels] (Section 2), we develop (Section 4) and validate (Section 5) a sliding-window procedure to estimate the extent to which open-weight models memorize different books contained in the Books3 training dataset. Our results show that the extent of verbatim memorization of books in open-weight LLMs is more significant than previously described. They also show that memorization varies widely from model to model and from book to book within each model, as well as varying in different parts of individual books (Section 6). We find that, with respect to our sliding-window probabilistic extraction procedure, most models do not memorize most books. But this is not always the case. In extreme instances, a model may memorize a book so significantly that it is possible to generate the entire book near-verbatim, as we have done with Llama 3.1 70B for Harry Potter and the Sorcerer’s Stone (Section 7).
Altogether, our results complicate current disputes over copyright infringement, both by rejecting easy claims made by both sides about how models work and by demonstrating that there is no single answer to the question of how much a model memorizes (Sections 3 & 8). There is much more work to do. We only ran book-specific, verbatim-extraction experiments for $50$ books in Books3. In future work, we intend to test significantly more books and models, as well as possible explanations for memorization variability and for the high level of memorization of some specific works by particular models.
Acknowledgments and Disclosure of Funding
The computing for this project was performed on the Sherlock cluster. We would like to thank Stanford University and the Stanford Research Computing Center for providing computational resources and support that contributed to these research results. We thank Jamie Hayes for chatting about additional interesting aspects concerning the $(n,p)$-discoverable extraction method [hayes2025measuringmemorizationlanguagemodels] that we did not cover in that original paper. We thank Timothy B. Lee for his extremely clear explanation of our work [lee2025arstechnica], which helped us simplify the presentation of the math in this paper. We would additionally like to thank Alexandra Chouldechova, danah boyd, Danny Wilf-Townsend, Florian Tramèr, Franziska Boenisch, James Grimmelmann, Katherine Lee, Kevin Klyman, Luis Villa, Matthew Jagielski, Matthew Sag, Milad Nasr, Nicholas Carlini, Pamela Samuelson, Paul Cantor, Stella Biderman, and others for helpful questions, comments, and feedback about earlier versions of this project.
A. Feder Cooper is a postdoctoral affiliate at Stanford University, in Percy Liang’s group in the Department of Computer Science and Daniel E. Ho’s group at Stanford Law School. Cooper is a full-time employee of Microsoft, working as a postdoctoral researcher in the FATE group within Microsoft Research, before being appointed as an Assistant Professor of Computer Science at Yale University in 2026. The opinions expressed about the results in this paper should not be attributed to Microsoft.
Mark A. Lemley is a full-time professor at Stanford Law School. He also has a consulting practice as a partner at Lex Lumina LLP. As part of that practice he has represented AI companies in copyright litigation. Nothing in this paper should be attributed to those companies.
Appendix Table of Contents
  • A
    Additional notes on memorization and probabilistic extraction ........................................................................................................................................................................A
    • A.1
      Additional notes on the “careless people” example ........................................................................................................................................................................A.1
    • A.2
      Metrics ........................................................................................................................................................................A.2
    • A.3
      Computing sequence probabilities in one forward pass ........................................................................................................................................................................A.3
  • B
    Testing our measurement pipeline ........................................................................................................................................................................B
  • C
    Experiments on extraction rates ........................................................................................................................................................................C
    • C.1
      Setup ........................................................................................................................................................................C.1
    • C.2
      Experimental procedure ........................................................................................................................................................................C.2
    • C.3
      High-level takeaways ........................................................................................................................................................................C.3
  • D
    Sliding-window experiments with our probabilistic extraction procedure ........................................................................................................................................................................D
    • D.1
      Setup ........................................................................................................................................................................D.1
    • D.2
      Sliding-window probabilistic extraction procedure ........................................................................................................................................................................D.2
    • D.3
      Book-specific results ........................................................................................................................................................................D.3
    • D.4
      Estimating how much of a book is memorized by a model ........................................................................................................................................................................D.4
      • D.4.1
        Quantifying extraction coverage ........................................................................................................................................................................D.4.1
      • D.4.1
        Extended extraction coverage results........................................................................................................................................................................D.4.2
    • D.5
      Selection of extracted sequences ........................................................................................................................................................................D.5
  • E
    Measurement validity ........................................................................................................................................................................E
    • E.1
      Motivating our validity experiments ........................................................................................................................................................................E.1
      • E.1.1
        Setting $\tau_{\text{min}}$ and extending to models with unknown training data ........................................................................................................................................................................E.1.1
      • E.1.2
        Relating greedy-decoded and probabilistic discoverable extraction ........................................................................................................................................................................E.1.2
      • E.1.3
        The role of prefix length in our experiments ........................................................................................................................................................................E.1.3
    • E.2
      Setup ........................................................................................................................................................................E.2
    • E.3
      Results of varying prefix length on known training data ........................................................................................................................................................................E.3
    • E.4
      Results of negative control experiments ........................................................................................................................................................................E.4
      • E.4.1
        Results on Phi 4 ........................................................................................................................................................................E.4.1
      • E.4.2
        Results on books from after training cutoffs ........................................................................................................................................................................E.4.2
    • E.5
      Additional notes on “the monkey at the typewriter” metaphor ........................................................................................................................................................................E.5
    • E.6
      Additional notes on extraction and membership inference ........................................................................................................................................................................E.6
      • E.6.1
        The relationship between membership inference and extraction ........................................................................................................................................................................E.6.1
      • E.6.2
        We know ground-truth membership for most of our experiments ........................................................................................................................................................................E.6.2
      • E.6.3
        We are inferring membership for models with undisclosed training data ........................................................................................................................................................................E.6.3
      • E.6.4
        Our inferences about membership are about sequences, not whole books ........................................................................................................................................................................E.6.4
      • E.6.5
        Absence of extraction does not imply anything about membership ........................................................................................................................................................................E.6.5
      • E.6.6
        Extractability and copyright ........................................................................................................................................................................E.6.6
  • F
    Reconstructing books with one seed prompt ........................................................................................................................................................................F
    • F.1
      Reconstruction method ........................................................................................................................................................................F.1
    • F.2
      Running partial reconstruction for books with partial high memorization signal ........................................................................................................................................................................F.2
    • F.3
      Running reconstruction for books with low memorization signal ........................................................................................................................................................................F.3
  • G
    Extended discussion of results ........................................................................................................................................................................G
Appendix A Additional notes on memorization and probabilistic extraction
We provide additional notes on the background material in this paper, including the “careless people” quote (Appendix A.1), more details on the key probabilistic extraction metric in this paper (called $(n,p)$-discoverable extraction, Appendix A.2), as well as our implementation (roughly speaking) for this metric (Appendix A.3). Please also refer to Section 2 of the main paper text.
A.1 Additional notes on the “careless people” example
The headline figure we include in Figure 1 (Section 2) is meant to be illustrative—an instance of, not a general statement about, memorization and extraction. This example also plays a role in some of our validity experiments (Section 5 and Appendix E). The text is drawn from The Great Gatsby [gatsby]; it is one of the most famous quotes from the book. It is also a short, $57$-token sequence: the prefix is $25$ tokens and the suffix is $32$ (with the Llama 1 tokenizer). We do not focus on such short suffixes in our analysis, for reasons that are discussed both in the main text and in our experimental details below. While extraction of verbatim $32$-token suffixes is often accepted as valid evidence for memorization [biderman2023pythia], it is now more common practice to use $50$-token suffixes. (Regarding validity, also refer to Section 5 and Appendix E.) We compute the extraction probability with respect to top-$k$ sampling ($T=1$, $k=40$). We do this by transforming the model logits according to the sampling parameters of this scheme. (See more about this below in Appendix A.3.)
It is also the case that very famous quotes like this are duplicated in many locations in training-data sources (e.g., blogs, essays, personal web pages, etc.) We are not claiming (here, or in general) that this sequence can be extracted because a given book (here, The Great Gatsby) was included in Books3; it could be extracted (or more extractable) because it is duplicated in many places in the training data. Of course, one of those copies is in Books3, if the entirety of Books3 is included in the model’s training data. We discuss this further in Appendix E.
A.2 Metrics
The metric that we use for memorization comes from hayes2025measuringmemorizationlanguagemodels. The authors provide a different framing for the metric. Instead of directly comparing $p_{\bm{z}}$ values—for the same sequence across LLMs, across multiple sequences for the same LLM, etc.—hayes2025measuringmemorizationlanguagemodels specify a threat model with an adversary that can query the model an arbitrary number of times with the same prompt—an adversary that has $n$ independent attempts to extract a sequence $\bm{z}$} with at least probability $p$. This is where their metric derives its name, $\bm{(n,p)}$-discoverable extraction:
Definition 1 ($(n,p)$-discoverable extraction, from hayes2025measuringmemorizationlanguagemodels).
Given a training sequence $\bm{z}$} that is split into an $a$-length prefix ${\bm{z}}_{1:a}$ and a $j$-length suffix ${\bm{z}}_{a+1:a+j}$, $\bm{z}$} is $(n,p)$-discoverably extractable if
$\Pr\Big(\cup_{w\in[n]}(g_{\phi}\circ f_{\theta})_{w}^{j}({\bm{z}}_{1:a})={\bm{z}}_{1:a+j}\Big)\geq p,$
where $(g_{\phi}\circ f_{\theta})_{w}^{j}({\bm{z}}_{1:a})$ represents the $w$-th (of $n$) independent execution of the autoregressive process of generating a distribution over the token vocabulary, sampling a token from this distribution, and adding the token to the sequence $j>0$ times, starting from the same initial sequence ${\bm{z}}_{1:a}$.
On its own, this type of method is well-suited to completion models that continue the prompt text, rather than instruction-tuned chatbots that reply in a conversational style. We therefore only study completion models and, for brevity, refer to these models by name, version, and size. Also note that this definition can be (at least in principle) easily adapted to non-verbatim extraction by testing if a generation is within $\epsilon$ distance from the target suffix for a given distance metric.
Despite these differences in framing, in practice for verbatim extraction, computing this metric just involves calculating sequence probabilities $p_{{\bm{z}},\theta,\phi}$. For sequence $\bm{z}$}, this is the suffix probability with respect to an LLM’s base distribution $\theta$ and decoding scheme $\phi$ (e.g., greedy, top-$k$) that determines which tokens are considered at each step and how their logits are renormalized into conditional probabilities (Appendix A.3). That is,
$p_{\bm{z}}\coloneqq p_{{\bm{z}},\theta,\phi}$ $=\Pr_{\theta,\phi}\!\big({\bm{z}}_{a+1:a+j}\,\big|\,{\bm{z}}_{1:a}\big)$
$=\prod_{t=a+1}^{a+j}\Pr_{\theta,\phi}({\bm{z}}_{t}\mid{\bm{z}}_{1:t-1})$
$=\exp\!\Bigg(\sum_{t=a+1}^{a+j}\log\Pr_{\theta,\phi}\!\big({\bm{z}}_{t}\,\big|\,{\bm{z}}_{1:t-1}\big)\Bigg).$ (5)
We will always note the particular LLM $\theta$ and decoding scheme $\phi$ being evaluated, and so, for brevity, will refer to this quantity as simply $p_{\bm{z}}$. In practice, we compute $p_{\bm{z}}$ as the $\exp$ of the sum of conditional $\log$ probabilities of the tokens in the sequence, as this is more numerically stable than multiplying together the conditional probabilities of the tokens in the sequence. (See Section 2.)
In Figure 10 (left), we show examples of $p_{\bm{z}}$ for the “careless people” quote (split into prefix and suffix as in Figure 1) for different models. (This is Figure 2 in Section 2, with Pythia 12B added.) Once one has computed $p_{\bm{z}}$, they can use $p_{\bm{z}}$ to compute $n$ for a given $p$ (for $p_{\bm{z}}\in(0,1)$) with
$1-(1-p_{\bm{z}})^{n}\geq p\;\Longrightarrow\;n\geq\frac{\log(1-p)}{\log(1-p_{\bm{z}})}.$ (6)
Figure 10: Plotting extraction probability $p_{\bm{z}}$ for the “careless people” quote from The Great Gatsby (see Section 2). (left) $p_{\bm{z}}$ (1) for different models. (right) Translating $p_{\bm{z}}$ into how many prompts $n$ it would take to extract the sequence $z$ with at least probability $p$ (A.2).
Since $p_{\bm{z}}$ is the probability of generating the suffix with $1$ prompt to the model, the probability of not generating it in $1$ prompt is $(1-p_{\bm{z}})$. The probability of not generating it in $n$ independent prompts to the model (with the same prompt) is $(1-p_{\bm{z}})^{n}$, and so the probability of generating it in $n$ prompts is $1$ minus this probability, i.e., $1-(1-p_{\bm{z}})^{n}$.
As an intuition for this presentation of the metric, think about a fair coin flip coming up heads. Here, $p_{\bm{z}}=0.5$ (the probability of heads in one flip). With two actual flips ($n=2$), the probability of flipping heads at least once is $1-(1-0.5)^{2}=0.75=p$; when $n=10$, $p=0.9990$. We visualize this intuition in Figure 11 (left) with coin flips, and show how it translates to suffix generation probabilities for Llama 1 30B and the “careless people” quote, for which $p_{\bm{z}}=35.2\%$ (right). As is clear Equation (6), for a given $p_{\bm{z}}$, one can pick a probability threshold $p$ and get the corresponding number of prompts $n$ (or vice versa). Figure 10 (right) visualizes how $n$ changes for different settings of $p$ for the “careless people” quote.
In this paper, we re-frame $(n,p)$-discoverable extraction to emphasize the quantity $p_{\bm{z}}$ (A.2) that we actually compute: the probability of extracting a sequence $\bm{z}$} (for the given LLM, hyperparameter-configured decoding scheme, and suffix start index location within sequence $\bm{z}$}). This is also useful for plotting distributions over $p_{\bm{z}}$ for a given book—to see how probabilities vary for sequences across a book, or for a given sequence across different models. We provide such comparisons in Section 6.
In practice, for LLMs, we can compute $p_{\bm{z}}$ for verbatim extraction without prompting $n$ times. We compute sequence probabilities directly from the logits we obtain from running sequences through the LLM, rather than generating multiple suffixes for each prefix. (See Appendix A.3.) These two approaches are functionally equivalent for verbatim extraction: the computed sequence probabilities correspond to the statistically expected frequencies of verbatim extracted outputs we would observe if we were to generate a large number of sequences for the same prompt. hayes2025measuringmemorizationlanguagemodels, the authors of this extraction approach, confirm this with ample experimental evidence. We do, as well, though omit these results for brevity because they follow directly from what $p_{\bm{z}}$ is measuring.
We limit ourselves to verbatim extraction in this paper; it is significantly computationally cheaper to do so because we can operate directly on model logits with one forward pass. (See Section 2 and Appendix A.3 for more details.) As a result, our work here (in several respects) only scratches the surface: there are various different ways to instantiate this extraction methodology in practice, we only explore a limited set of open-weight models, and we only test $50$ of the nearly 200,000200{,}000 books in Books3.
We further discuss how we determine extraction success for probabilistic extraction in Appendix C, where we describe how extraction success for different sequences can be aggregated into an overall extraction rate. In the main text, we discuss determining extraction success in Section 2.1.
Figure 11: Showing how the probability $p$ of generating a targeted outcome at least once changes as a function of $n$ independent trials. Following the intuition of flipping a fair coin (where heads has $p_{\bm{z}}=0.5=50\%$), we show how the probability of flipping heads at least once changes with more flips (left). We show how the probability $p$ of generating the verbatim suffix of the “careless people” quote for Llama 1 30B ($p_{\bm{z}}\approx 35.2\%$) changes as a function of the number of independent prompts to the model with the prefix (right). These curves both use $1-(1-p_{\bm{z}})^{n}$ (6), plugging in the respective $p_{\bm{z}}$ (for the coin flip and for generating the “careless people” suffix given the prefix) and different values of $n$ to compute $p$.
A.3 Computing sequence probabilities in one forward pass
Given that the core algorithm for $(n,p)$-discoverable extraction is so elegantly concise, we include a (simplified) version of the code in this appendix. Our implementation differs from hayes2025measuringmemorizationlanguagemodels, which we confirm in discussion with the authors of that work. In that work, the authors generated the suffix a token at a time and summed up the per-generated-token conditional $\log$ probabilities. Here, we observe that this is not necessary. We can get the logits for each token in the sequence with just one forward pass through the model, and effectively only one line of code. (See Listing 1, line 28.) For a $50$-token suffix, this results in up to a $50\times$ decrease in compute—an enormous savings. (It is of course less with KV caching enabled during generation, but still nonzero). This also makes computing this metric cheaper than traditional greedy-decoded discoverable extraction [carlini2021extracting], which generates the greedy-decoded sequence (i.e., for a $50$-token suffix, there are $50$ forward passes through the model, even if those passes involve caching).
Note that this is the entire interaction with the model in our experiments using our probabilistic extraction procedure: for a given sequence, we perform inference on the GPU, get the logits tensor, and move the logits to the CPU for processing. (There are perhaps additional opportunities for improving the efficiency of our implementation that involve KV-caching and streaming logits, rather than performing a giant matmul. We will explore this in the future.) We are not making any changes to the model in our extraction measurements; we are just getting the raw logits from running a given training-data sequence through the model. (As shown in Listing 1, we do this for a batch of sequences at a time, yielding a batch of logits.)
The logits tensors we get are of shape [batch_size, sequence_length, vocabulary_size]. This is a three-dimensional tensor:
  • Dimension 1: We have a slot for each sequence in the batch (with batch_size sequences overall).
  • Dimension 2: We have a position for each index in the token sequence (with sequence_length=100{=}100 being the number of total positions in most of our experiments).
  • Dimension 3: We have a position for each token in the whole token vocabulary for this model (with vocabulary_size being the overall number of positions, e.g., this is 32,00032{,}000 for the Llama 1 family).
So, at position [i, j, k] in this tensor, we have the logit (unnormalized $\log$ probability) value for the i-th sequence in the batch, at the j-th position in the token sequence, for the k-th token in the whole token vocabulary. That is, for a given token sequence in the batch (at [i,j]), we have the logit values for all of the next-possible tokens in the whole token vocabulary (a vocabulary_size-length list of numbers, one for each token in the whole vocabulary); these values can be normalized to compute the probability distribution over the next token in the sequence.
We then implement our sequence probability computations (i.e., compute $p_{\bm{z}}$) as a post-processing operation on the logits we get from the model. (See Listing 2.) That is, for any given decoding scheme, we can transform the logits—for example, change the shape according to temperature, truncate them to the top-kk—before normalizing them into the $\log$ probability distribution.
1import torch
2"""
3Runs inference to get logits for sequences z, which we will use
4to compute (n,p)-discoverable extraction (i.e., to compute p_z).
5 - dataloader: a torch.utils.DataLoader wrapping a torch.utils.data.Dataset containing the tokenized sequence z
6
7 - model: loaded transformers.AutoModelForCausalLM (in this snippet, assumed to have layers distributed across GPUs with device_map="auto", though for smaller models we use data parallelism)
8
9 - returns all_logits (logits indexed by batch number). In practice, we do not do this. We asynchronously do computations on logits on CPU to save log probs.
10"""
11def compute_logits_for_batch(dataloader, model):
12
13 all_logits = {}
14 with torch.inference_mode():
15 for batch_idx, batch in enumerate(dataloader):
16 input_ids = batch["input_ids"]
17 attention_mask = batch["attention_mask"]
18 seq_metadata_batch = batch["metadata"]
19
20 device = model.device
21
22 inputs = {
23 "input_ids": input_ids.to(device),
24 "attention_mask": attention_mask.to(device)
25 }
26
27 # Run inference; shape: [batch_size, seq_len, vocab_size]
28 logits_batch = model(**inputs).logits.detach().cpu()
29
30 input_ids = input_ids.detach().cpu()
31 attention_mask = attention_mask.detach().cpu()
32
33 all_logits[batch_idx] = logits_batch
34 return all_logits
Listing 1: Computing logits with one forward pass through the model. This is the key speedup in our implementation and this is the only code that interacts with a given model that we study in our main extraction experiments.
This computation is written out on the right side of Equation A.2. The key point here is that, for a given sequence that we are evaluating, all we do (starting at the first index in the suffix) is sum up the conditional $\log$ probabilities of the actual next tokens in the sequence, one at a time (Listing 2, lines 32-40). This is often called teacher forcing.
We work with the conditional $\log$ probabilities (adding them together), rather than the conditional probabilities (and multiplying them together), as this is more numerically stable to compute. However, note that these computations are equivalent. For example, consider the sequence "the orange cat" (where, for simplicity, we will assume each word is a single token). When processing "the", the logits reflect the probabilities of the next token. We find the logit value in the distribution that corresponds to "orange" and add the associated conditional $\log$ probability, and then continue on (doing the same for "orange", where we then get the conditional $\log$ probability for the next token being "cat"). We get the conditional $\log$ probability of the whole suffix by adding together the conditional, per-token $\log$ probabilities of the actual suffix. From the suffix $\log$ probability, we can easily obtain the probability $p_{\bm{z}}$: the probability of the model (under the given decoding scheme we are using) generating that suffix conditioned on the prior context (the prefix).
Any given sequence should be low probability. That is, for a suffix length of $50$ tokens, this computation is effectively multiplying together $50$ per-token conditional probabilities (by adding together conditional $\log$ probabilities). For a sequence where each token has really high probability, this can still mean the overall sequence has low probability.
Consider that each token in a suffix has conditional probability $0.9$. This is really high; it means that a single token in the logit vector (at each position in the suffix) has $90\%$ of the probability mass; for a token vocabulary of size $32,000$, that means the remaining $31,999$ tokens all share the remaining $10\%$. For this suffix, the probability of generating it is $0.9^{50}\approx 0.005$—i.e., a $0.5\%$ chance of generation. This is a relatively small number in the scheme of things; but it is also a really large number in general, if we consider what this means for the underlying token probabilities. In some cases in this paper, we observe sequences that have over $90\%$ probability—an enormous number, especially considering what the underlying computation is.
Such unusually high probabilities associated with verbatim suffixes in the training data are indicative of memorization (where it is reasonable to consider $0.5\%$, or even smaller, to be unusually high). As cooper2024files say, “the pattern [the model has learned from the training data] is the memorized training data.”
Also note that top-$k$ sampling truncates the logit vector to only the top-$k$ tokens (e.g., for a vocabulary of 32,00032{,}000 tokens, there are 32,00032{,}000 logits, but top-$k$ with k=40k{=}40 decoding will only consider the $40$ highest-valued logits). This means that, if a given sequence has a token at a given position that is not in the top-$k$ of the logit vector, then we will not be able to complete the probability computation. This is what the has_impossible_token check is referring to in Listing 2, lines 36–38. In this case, the given sequence has 0 probability; it is not extractable with any probability.
In practice, our code deviates from Listing 2 for efficiency reasons. (The core logic is nevertheless the same as what we show here.) As we process conditional $\log$ probabilities, we save a lot of metadata in order to produce the plots that we include in this paper. For additional speedups (which, ultimately, total around $200\times$ prior work on $(n,p)$-discoverable extraction), as we compute logits on the GPU (Listing 1), we delegate our post-processing code to happen in parallel on CPU.
1import torch, torch.nn.functional as F, math
2"""
3For a batch of logits, computes the log probability (and probability)
4of the suffix given the prefix for a batch of sequences.
5 -inputs: a batch of sequences
6 -logits: a batch of logits (computed doing inference, above)
7 -suffix_start_idx: index in sequences where suffix begins
8 -transforms base distribution for temperature (temp) and top_k (k).
9 -returns list of logprobs (and probs)
10"""
11def compute_logprob_batch(inputs, logits, suffix_start_idx, temp, k):
12 input_ids = inputs_batch["input_ids"]
13 attention_mask = inputs_batch["attention_mask"]
14 if temp != 1.0 or k is not None:
15 logits = apply_temperature_and_topk(logits, temp, k)
16
17 # normalize the logits to a probability distribution;
18 # surfaces each next token’s log prob
19 # shape: [batch_size, seq_len, vocab_size]
20 log_probs = F.log_softmax(logits, dim=-1)
21 # log_probs[b, i, j]: for sequence at batch index b, log probs of
22 # predicting token j at position i in the sequence
23 log_probs_list = [], prob_list = [], batch_size = input_ids.shape[0]
24 for b in range(batch_size):
25 input_ids_b = input_ids[batch_b]
26 # shape: [seq_len, vocab_size]
27 log_probs_b = log_probs[b]
28 real_seq_len = attention_mask[b].sum().item()
29 total_log_prob = 0.0, has_impossible_token = False
30 # Compute log prob for each token in suffix conditioned on prefix
31 for i in range(suffix_start_idx - 1, real_seq_len - 1):
32 # get actual next token in suffix
33 actual_next_token_id = input_ids_b[i + 1]
34 # log prob for actual_next_token i+1 based on context up to i
35 log_prob = log_probs_b[i, actual_next_token_id].item()
36 # possible -inf from top-k filtering
37 if not math.isfinite(log_prob):
38 has_impossible_token = True
39 break
40
41 total_log_prob += log_prob
42
43 if has_impossible_token:
44 total_log_prob = float("-inf")
45 prob = 0.0
46 else:
47 prob = math.exp(total_log_prob)
48 log_probs_list.append(total_log_prob)
49 prob_list.append(prob)
50
51 return log_probs_list, prob_list
52
53def apply_temperature_and_topk(logits, temp, k):
54 if temperature != 1.0:
55 logits = logits / temperature
56 if top_k is not None:
57 # top-k values from logits; shape: [batch, seq_len, top_k]
58 top_k_vals = torch.topk(logits_batch, top_k, dim=-1)
59 # k-th element in top-k values for each item in batch
60 kth_vals = top_k_vals.values[:, :, -1].unsqueeze(-1)
61 mask = logits < kth_vals
62 logits = logits.masked_fill(mask, float("-inf"))
63 return logits
Listing 2: Computing logprobs
Appendix B Testing our measurement pipeline
To confirm the accuracy of our implementation of probabilistic discoverable extraction, as well as to benchmark its efficiency in comparison to the implementation in hayes2025measuringmemorizationlanguagemodels, we re-run the experiments from hayes2025measuringmemorizationlanguagemodels for small Pythia models on the Enron dataset. To ensure consistency, we discussed the original run times and sample of Enron with the authors of hayes2025measuringmemorizationlanguagemodels. We provide some brief details about these experiments to show due diligence for our implementation in our particular setting.
We ran these experiments with both float16 and float32, and report results for float32 to align with hayes2025measuringmemorizationlanguagemodels. (However, we note no significant differences between the two.) In Table 1, we show the runtime for our experiments on the 10,00010{,}000, $100$-token sequences hayes2025measuringmemorizationlanguagemodels drew from Enron. In Table 2, we report the same metrics as hayes2025measuringmemorizationlanguagemodels: the greedy-decoded discoverable extraction rate and (as a point of comparison) the maximum probabilistic rate. We run these experiments on the same $4$ A100s as all of our other probabilistic extraction experiments.
Time (mm:ss) Parallelism Batch size
Pythia 1B 02:30.42 Data (four GPUs) 250
Pythia 2.8B 03:36.35 Data (four GPUs) 250
Table 1: Runtime for different Pythia models on a 10,00010{,}000-sequence dataset drawn from Enron.
Pythia 1B on Enron subset hayes2025measuringmemorizationlanguagemodels (No BOS token) With BOS token
Greedy extraction rate 0.76% 0.74%
Max. probabilistic rate 5.27% 5.52%
Pythia 2.8B on Enron subset hayes2025measuringmemorizationlanguagemodels (No BOS token) With BOS token
Greedy extraction rate 1.3  % 1.82%
Max. probabilistic rate 9.04% 9.47%
Table 2: Greedy discoverable extraction and maximum probabilistic rate for Pythia models on a 10,00010{,}000-sequence subset from Enron. We successfully replicate the exact results in hayes2025measuringmemorizationlanguagemodels, which did not include the beginning of sequence (BOS) token in the sequences they use for their experiments using Pythia models. We compare to the extraction rates where the BOS token is included at the start of the sequence. Note that the probabilistic rate increases slightly with the presence of the BOS token.
We successfully replicate (exactly) the results of these experiments in hayes2025measuringmemorizationlanguagemodels. However, in the process of doing so, we identified a small bug in their original implementation. The Pythia tokenizer is a GPT-2-style tokenizer, and does not add a beginning of sequence (BOS) token by default to the sequences it tokenizes. However, for our setting, it is best practice to include this token at the start of a sequence, as we submit these sequences as inputs for inference to LLMs (that are trained to expect this token at the beginning of the sequence). When we manually prepend this token and re-run the experiments from hayes2025measuringmemorizationlanguagemodels, we observe elevated extraction rates compared to hayes2025measuringmemorizationlanguagemodels. See Table 2. We similarly prepend the BOS token in our Books3 experiments using Pythia models.
Appendix C Experiments on extraction rates
Similar to other extraction papers, we also compute average extraction rates. These serve as a reference point for our discussion concerning how average rates can conceal interesting patterns of work-specific memorization. In the main text, we refer briefly to these experiments in Section 4.2.
Computing average extraction rates.
Let $\mathbb{Z}$} be a set of sequences $\bm{z}$} drawn from some (part of a) training dataset. Each $\bm{z}$} is split into a prefix ${\bm{z}}_{1:a}$ and target suffix ${\bm{z}}_{a+1:a+j}$ (e.g., a=j=50a{=}j{=}50 for $100$-token sequences).
All extraction rates take the form
$\mathsf{extraction\_rate}({\mathbb{Z}};s)\;=\;\frac{1}{|{\mathbb{Z}}|}\sum_{{\bm{z}}\in{\mathbb{Z}}}\mathbf{1}[\,s({\bm{z}})\,],$ (7)
where $s({\bm{z}})$ is a success predicate defined according to the chosen criterion.
Fix a model $\theta$ and a decoding scheme $\phi$ (e.g., base distribution T=1T{=}1, or top-$k$ with k=40k{=}40), which induces a conditional next-token distribution that can be used to compute suffix probabilities $p_{\bm{z}}\in[0,1]$. For a sequence $\bm{z}$} with prefix ${\bm{z}}_{1:a}$ and suffix ${\bm{z}}_{a+1:a+j}$, define the suffix probability under $\theta$ and $\phi$ as
$p_{\bm{z}}\coloneqq p_{{\bm{z}},\theta,\phi}\;=\;\prod_{t=a+1}^{a+j}\Pr_{\theta,\phi}({\bm{z}}_{t}\mid{\bm{z}}_{1:t-1}).$
This is the probability that the model generates the exact target suffix token by token under decoding scheme $\phi$; it is the same definition as Equation A.2 that we use in Section 2 and Appendix A.
For traditional, greedy-decoded discoverable extraction [carlini2023quantifying, lee2022dedup] (Section 2), $\phi{=}\mathsf{greedy}$ (i.e., T=1,k=1T{=}1,k{=}1) and the success predicate can be written simply as explicit string equality:
$s_{\theta,\mathsf{greedy}}({\bm{z}})\;=\;\mathbf{1}\!\big[\,\mathsf{greedy}(\theta;{\bm{z}}_{1:a})={\bm{z}}_{a+1:a+j}\,\big].$ (8)
More generally, for any LLM $\theta$ and decoding scheme $\phi$, the success predicate just thresholds the suffix probability. Given a minimum threshold $\tau_{\text{min}}\in(0,1]$,
$s_{\theta,\phi}({\bm{z}};\tau_{\text{min}})\;=\;\mathbf{1}\!\big[\,p_{{\bm{z}},\theta,\phi}\geq\tau_{\text{min}}\,\big].$ (9)
Again, note that greedy decoding is a special case: each top-$1$ token has probability $1$ under $\phi=\mathsf{greedy}$, so $p_{{\bm{z}},\theta,\phi}=1$ and success reduces to equality with the suffix. In the main paper, we omit $\theta$ and $\phi$ in our notation. We always mention which specific model we are evaluating, and we are always considering $\phi$ to be top-$k$ decoding with T=1T{=}1 and k=40k{=}40. (See Appendix D.2 for more details on this choice.) When using non-deterministic decoding schemes and suffix probabilities to quantify extraction, we very conservatively set $\tau_{\text{min}}{=}0.1\%$ (Section 5 & Appendix E).
C.1 Setup
Data.
We perform our average extraction rate experiments with the Books3 dataset. We obtained this dataset from a previous (2022) download of The Pile [gao2020pile], which is stored on a university cluster and which we use for research purposes on language modeling.
The status of Books3 as a research artifact (as well as a training corpus for LLMs) remains unresolved. (See Section 3.) Nevertheless, this dataset remains widely available in the research community; for example, it can be found in the HuggingFace-hosted version of The Pile [hfpilededup], as well as in various other data repositories [e.g., books3a, books3b, books3c]. It also features as part of a benchmark task in the popular HELM evaluation suite.
For this set of experiments, we draw random samples from Books3, performing $3$ separate runs each on 40,00040{,}000 sequences. We discuss how we sample these sets of sequences in Appendix C.2.
Models.
From $13$ different models, we try to extract 40,00040{,}000 sequences from Books3 ($3$ times each). We test Llama 1 7B, Llama 1 13B, Llama 1 65B [touvron2023llamaopenefficientfoundation], Llama 2 7B, Llama 2 13B, Llama 2 70B [llama2], Llama 3.1 8B, Llama 3.1 70B [llama3], Pythia 6.9B, Pythia 12B [biderman2023pythia], DeepSeek v1 7B [deepseekv1], Gemma 2 9B [gemma2], and Phi 4 [phi4] (a 14B model). We group these models into three size groups: small, medium and large.
We use float16 for all models except for Llama 3 and Llama 3.1 models (and Qwen 2.5 models in sliding-window experiments). These models were explicitly trained to work with bfloat16.
We know for certain that Pythia and Llama models were trained on Books3, and can conclude from our validity experiments (Section 5 and Appendix E) that DeepSeek v1, Qwen 2.5 and Gemma 2 models were trained on (at least parts of) some books that are also contained in Books3. Phi 4 was deliberately not trained on copyrighted (whole) books, but we find that we can still extract some snippets from copyrighted books.
Metrics.
We compute extraction rates using Equation 7 for both traditional, greedy-decoded discoverable extraction (using the success condition in Equation 8) and probabilistic extraction with top-$k$ (T=1T{=}1, k=40k{=}40) decoding (using the success condition in Equation 9) and $\tau_{\text{min}}{=}0.1\%$.
Compute resources.
We run all of our average extraction rate experiments in a slurm cluster environment, using the same node with 4 A100 GPUs.
C.2 Experimental procedure
Sampling books from Books3.
To collect 40,00040{,}000 sequences for which we compute an extraction curve for a given model, we use the following procedure. We sample 4,0004{,}000 books without replacement from Books3. Then, from each of these books, we sample $10$ non-overlapping $100$-token sequences ($50$-token prefixes ++ $50$-token suffixes). We make sure that the start of the prefix text is a space (to prevent unusual tokenization). We run these 40,00040{,}000 sequences through each of the $13$ models to compute extraction metrics. We follow this procedure $3$ separate times (i.e., for a total of 120,000120{,}000 sequences across 12,00012{,}000 different books) to get a sense of the variance across extraction metrics for different (empirical) population estimates. Each run of 40,00040{,}000 sequences took between approximately $8$ minutes and $45$ minutes on $4$ A100s. We omit detailed timing results for brevity.
Probabilistic extraction configuration.
We report greedy-decoded discoverable extraction and probabilistic discoverable extraction with top-$k$ (T=1T{=}1, k=40k{=}40) and $\tau_{\text{min}}{=}0.1\%$. The maximum probabilistic rate (also reported in Appendix B) should be understood as a very loose upper bound on extraction of memorized training data; it counts all sequences with non-zero $p_{\bm{z}}$ (i.e., it also includes sequences for which $p_{\bm{z}}\!<\!\tau_{\text{min}}\!=\!0.1\%$). For sequences with these low $p_{\bm{z}}\!<\!\tau_{\text{min}}$, it may also capture generalization.
C.3 High-level takeaways
We defer discussion of validity to Appendix E, where we talk about how to draw conclusions for different models—some for which we know with certainty that Book3 is in their training data, and others for which we can conclude that at least some content that overlaps with Books3 was in their training data.
We attempt to use Phi 4 as a reference. Unlike the other models we test, Phi 4 was trained predominantly on synthetic data. Based on the Phi 4 technical report [phi4], Phi 4 was deliberately not trained on copyrighted (whole) books. As a result, Books3 (the specific corpus) was almost certaintly not included in its training data. Nevertheless, we are sometimes able to extract data that is contained in Books3. These sequences likely found their way into the training data from web-scraped sources that quote these books. We discuss this further in Section 5, Appendix E, and Appendix G.
Figure 12 shows average extraction rates grouped roughly by model size. As we expected (and discuss in Section 3), average extraction rates are relatively low. Using the very conservative threshold of $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$, the lowest rate we observe is for Phi 4 ( $0.2\%$), and the highest is for Llama 3.1 70B ($\sim\!0.6\%$). Consistent with prior work, our results show that larger models tend to memorize more than smaller ones [hayes2025measuringmemorizationlanguagemodels, carlini2023quantifying, nasr2023scalable, nasr2025scalable]. This holds within the same model family. Phi 4 is an exception; despite being a 14B model, its extraction rates on Books3 sequences are comparable to the smaller (7–9B) models we evaluate. Small Llama 3 models memorize on-par with medium Llama 1 and 2 models. In general, later generations of Llama model families tend to memorize more Books3-contained text. For just one example, Llama 3.1 70B memorizes more on average than Llama 2 70B, which memorizes more on average than Llama 1 65B. This pattern does not always hold for specific books. We defer additional, related observations about Llama 3.1 70B to Appendix G.
Altogether, it is important to keep in mind that any given sequence that we deemed extracted is, by definition, a member of the training data; however, it is not necessarily a member of the training data because it came from the Books3 corpus. The sequence may be part of another source (or multiple other sources) in the training data. For Pythia and Llama models, the Books3 copy of a sequence may just be one of several duplicate copies in the whole training dataset. The same can be said for models like Gemma 2, where we are not sure if the Books3 corpus was included in the training dataset.) Nevertheless, it is clear from our results that sequences that are contained in Books—but perhaps originate from other sources—are included in the training data (Section 5 & Appendix E).
Figure 12: Comparing greedy discoverable and probabilistic extraction rates (for a our conservative setting, $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$). As a reference point, we also include the maximum possible rate of generating $50$-token sequences that match our suffixes drawn from Books3 for small (7–9B), medium (12–14B), and large models (65–70B). (This is akin to looking at all $p_{\bm{z}}>0$.) Probabilistic rates are computed with top-$k$ sampling using T=1T{=}1 and k=40k{=}40. The maximum probabilistic rate should be understood as a very loose upper bound on extraction of memorized training data; it also captures generalization. Average extraction rates (i.e., $\tau_{\text{min}}{=}0.1\%$) on randomly sampled Books3 sequences are relatively low for all models. Aligning with prior work, we observe that larger models memorize more. For our setting, this is true for both within and across model families. Phi 4, a 14B model, is an exception: its extraction rates are on par with smaller models. Error bars here indicate variance across the 33 samples of 40,00040{,}000 sequences that we run.
Appendix D Sliding-window experiments with our probabilistic extraction procedure
We discuss details for experiments that use a sliding window across a book to surface memorization “hot-spots” at particular book locations. We discuss the overall setup (Appendix D.1), our sliding-window procedure (Appendix D.2), detailed results for $50$ books and $10$ models (Appendix D.3), our extraction coverage metric (Appendix D.4), and instances of extracted sequences (Appendix D.5). We describe other experiments in different appendices. We defer discussion of sliding-window validity experiments on non-training data and different prompt prefix lengths to Appendix E. Additional experiments concerning extraction rates and our reconstruction procedure are discussed in Appendix C and Appendix F, respectively.
D.1 Setup
Data and book selection.
We draw our prompts (prefixes) from a selection of books in Books3; we verify extraction against the corresponding suffixes drawn from Books3. Books3 is a common pretraining dataset that is known to have been used to train several open-weight LLMs (Appendix C.1, [reisnerbooks3, lee2023talkin]. (See models, below). We list the $50$ selected books in Table LABEL:app:tab:sliding-books.
We began our project with The Great Gatsby [gatsby]. This choice stems from a conversation in 2023 while writing a different paper [lee2023talkin], when Anthropic announced its 100,000100{,}000-token context for Claude with the example of prompting with the entirety of The Great Gatsby [claudegatsby]. We then continued our selection process with the books listed with the associated plaintiffs in the (amended) class action complaint of Kadrey et al. v. Meta, Inc. [kadreyamendedconsolidated, pp. 4-5]. This is how we sourced the first $13$ in-copyright books from Books3. We expanded to additional books by these plaintiffs, and also added in some generally popular books that we chose among the project team (e.g., hp1, hobbit, sisyphus, catch22) and some (less publicly popular) academic books based on our personal preferences (e.g., zittrain, dante). (This team, after all, is composed of academics.) We deliberately include selections from the public domain (alice, gatsby, ulysses), and an example of a book that was published under a permissive CC license (downandout).We added in alice because of concurrent work that was published on arXiv just as we posted our own [ma2025memorizationcloselookbooks]. To round out the list, we also selected books at random from the manifest file that accompanies the Books3 archive (e.g., onions, emirate, dungeons). Once we observed that certain very popular books were highly memorized by Llama 3.1 70B, we added additional bestsellers to our set (e.g., thrones, davinci, twilight, 1984, leanin).
Models.
Altogether, we run this set of experiments on $17$ completion-style (i.e., non-chatbot) models. We includes results for $10$ of these models: Pythia 12B [biderman2023pythia], Llama 1 13B, Llama 1 65B [touvron2023llamaopenefficientfoundation], Llama 2 70B [llama2], Llama 3 70B, Llama 3.1 70B [llama3], DeepSeek v1 67B [deepseekv1], Qwen 2.5 72B [qwen2.5], Gemma 2 27B [gemma2], and Phi 4 (a 14B model) [phi4]. For brevity, the $4$ (smaller) models we do not include in our results are Pythia 6.7B, Llama 1 7B, Llama 2 7B, Llama 2 13B, Llama 3 8B, Llama 3.1 8B, and DeepSeek v1 7B. We note that Llama 3 8B and Llama 3.1 8B exhibit an enormous degree of memorization on some books, for models in this size class (e.g., 1984).
We load all models with float16—except for Llama 3, Llama 3.1, and Qwen 2.5 models. These models were explicitly trained to work with bfloat16. In an earlier version of this work, we reported results for Llama 3.1 that used float16. Slight differences in results in this version are attributed to the change to bfloat16. (We have otherwise confirmed our results are deterministic, with respect to the hardware that we use throughout.) Regardless, the overall results are similar, and the conclusions remain the same.
We picked these models because they each fit (with respect to $16$-bit format) on $4$ A100s, and because they fall into three different categories that we wanted to investigate: (1) models that we know with certainty were trained on Books3, (2) models where we have reason to believe were trained on copyrighted books (whether drawn directly from Books3 or from other sources), and (3) a model that we trust was not trained on whole copyrighted books.
For the first category, when we generate verbatim completions for these models on Books3 data, we can very reasonably conclude that we have extracted data that these models have memorized during training. We know that these data were members of the training data; extraction success (and thus memorization) is a standard conclusion in this setting [e.g., biderman2023pythia, hayes2025measuringmemorizationlanguagemodels, prashanth2024recite]. For the second category, we would like to be able to do valid membership inference of copyrighted data. The third category is also to help confirm validity; as a baseline, we intentionally include a model for which we do not expect to be able to extract much (if any) copyrighted books data.
  1. 1.
    Category 1: Pythia 12B, Llama 1 13B, Llama 1 65B, Llama 2 70B, Llama 3 70B, and Llama 3.1 70B. We know that Pythia and Llama models were training on Books3. Pythia was trained on the Pile [gao2020pile] (which contains Books3), based on EleutherAI’s extensive documentation. We include only the largest ($12$-billion parameter) Pythia model.
    Llama 1 and Llama 2 are known with absolute certainty to have been trained on Books3. The Llama 1 report explicitly says so. It came out in discovery during Kadrey et al. v. Meta, Inc. [kadrey] that Llama 2 models were all also definitively trained on Books3.
    Both in documents in discovery and Meta’s otherwise public documentation, it is repeatedly noted that the development of Llama 3 continued to use the same data practices as prior versions. For minor-versioned Llama 3 models (like Llama 3.1), it is similarly documented by Meta in release reports that the same internal workflows were used, including for dataset curation [llama3]. While this is not as definitive evidence that these models were trained on Books3, based on what is public information about the training process and prior Llama models, it is widely believed in the ML community that Llama 3 (and associated minor versions) models were trained on Books3 (as well as additional books data from LibGen). We consider this sufficient to conclude with confidence that Llama 3 and Llama 3.1 were trained on Books3: Books3 text sequences were members of the training dataset. (And regardless of this reasoning, our validity experiments enable us to independently conclude this. See Section 5 and Appendix E.)
  2. 2.
    Category 2: DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B. Although we do not know with certainty whether these models were trained on Books3, we aim to use our extraction methodology to assess whether they have memorized content from books. In particular, we want to design a measurement procedure that lets us conclude—with high confidence—that a specific sequence was in the training data; we want to be able to conclude that reproducing such a sequence in an LLM’s output is valid extraction (i.e., evidence of memorization). Based on our validity experiments (Section 5 and Appendix E), we can be very confident that when we successfully generate Books3 data, that those data were memorized from the training data. We discuss those experiments elsewhere, and note here that when we observe high probabilities $p_{\bm{z}}$ for these models, we can safely conclude that those sequences $\bm{z}$} have been memorized.
  3. 3.
    Category 3: Phi 4. For validity reasons, we also want to compare on a model that definitively should not have been trained on Books3. (See also Appendix E.) We pick Phi 4 for these reasons, as it is a very high quality for its size class and phi4 note in the Phi 4 report that it was trained predominantly on synthetic data:
    • Synthetic data constitutes the bulk of the training data for phi-4 and is generated using a diverse array of techniques, including multi-agent prompting, self-revision workflows, and instruction reversal. …[Phi 4 was also trained on] web content, licensed books, and code repositories to extract seeds for the synthetic data pipeline. phi4 (emphasis added)
    Phi 4 is similar in size to Pythia 12B and Llama 1 13B, which facilitates cross-model comparisons of extraction risk. (This is because the amount of memorization, reasonably, also varies according to model size; see also carlini2023quantifying.)
Compute resources.
We run all of our sliding-window experiments in a slurm cluster environment, using the same node with $4$ A100 GPUs. All told, this set of experiments took $3$ GPU months. Altogether, this was the most costly set of experiment we ran. In total, all of our experiments took approximately $6$ GPU months.
Table 3: The $50$ books drawn from Books3 that we evaluated with our sliding-window probabilistic extraction procedure. See Sections 4 and 6, as well as Appendices D.1 & D.2.
Author Title Year Status Books3 path
1 Margaret Atwood The Handmaid’s Tale [handmaid] 1985 © the-eye.eu/public/Books/Bibliotik/M/Margaret Atwood (1985) The Handmaid_s Tale [retail].epub.txt
2 Teodolinda Barolini Dante and the Origins of Italian Literary Culture [dante] 2006 © the-eye.eu/public/Books/Bibliotik/D/Dante and the Origins of Italia - Barolini, Teodolinda;.epub.txt
3 Dan Brown The Da Vinci Code [davinci] 2003 © the-eye.eu/public/Books/Bibliotik/D/Dan Brown - The Da Vinci Code.epub.txt
4 Albert Camus (Justin O’Brien, translator) The Myth of Sisyphus [sisyphus] 1955 © the-eye.eu/public/Books/Bibliotik/2/2013(orig1942) Albert Camus - The Myth of Sisyphus[Transl Justin O’Brien]_Ral.epub.txt
5 Lewis Carroll Alice’s Adventures in Wonderland [alice] 1865 PD the-eye.eu/public/Books/Bibliotik/A/Alice’s Adventures in Wonderlan - Lewis Carroll.epub.txt
6 Ta-Nehisi Coates The Beautiful Struggle [struggle] 2009 © the-eye.eu/public/Books/Bibliotik/T/The Beautiful Struggle - Ta-Nehisi Coates.epub.txt
7 Ta-Nehisi Coates We Were Eight Years in Power: An American Tragedy [wewere] 2017 © the-eye.eu/public/Books/Bibliotik/W/We Were Eight Years in Power - Ta-Nehisi Coates.epub.txt
8 Ta-Nehisi Coates The Water Dancer: A Novel [waterdancer] 2019 © the-eye.eu/public/Books/Bibliotik/T/The Water Dancer - Ta-Nehisi Coates;.epub.txt
9 Jon Cogburn Dungeons and Dragons and Philosophy [dungeons] 2012 © the-eye.eu/public/Books/Bibliotik/2/2012 Jon Cogburn - Dungeons and Dragons and Philosophy - Raiding the Temple of Wisdom_Rsnl.epub.txt
10 Junot Díaz Drown [drown] 1996 © the-eye.eu/public/Books/Bibliotik/J/Junot Diaz - Drown.epub.txt
11 Junot Díaz The Brief Wondrous Life of Oscar Wao [oscarwao] 2007 © the-eye.eu/public/Books/Bibliotik/T/The Brief Wondrous Life of Oscar Wao.epub.txt
12 Junot Díaz This Is How You Lose Her [loseher] 2012 © the-eye.eu/public/Books/Bibliotik/T/This Is How You Lose Her-Diaz.epub.txt
13 Cory Doctorow Down and Out in the Magic Kingdom [downandout] 2004 $C$-BY-NC-SA the-eye.eu/public/Books/Bibliotik/D/Down and Out in the Magic Kingd - Cory Doctorow.epub.txt
14 Carol Ann Duffy The World’s Wife [worldswife] 2001 © the-eye.eu/public/Books/Bibliotik/T/The World_s Wife - Carol Ann Duffy.epub.txt
15 Jennifer Egan A Visit from the Goon Squad [goonsquad] 2010 © the-eye.eu/public/Books/Bibliotik/A/A\ Visit\ from\ the\ Goon\ Squad\ -\ Jennifer\ Egan.epub.txt
16 Christopher Farnsworth The President’s Vampire [vampire] 2011 © the-eye.eu/public/Books/Bibliotik/T/The President’s Vampire - Christopher Farnsworth.epub.txt
17 F. Scott Fitzgerald The Great Gatsby [gatsby] 1925 PD the-eye.eu/public/Books/Bibliotik/T/The Great Gatsby - F. Scott Fitzgerald.epub.txt
18 Malcom Gladwell Blink: The Power of Thinking Without Thinking [blink] 2005 © the-eye.eu/public/Books/Bibliotik/B/Blink - Malcolm Gladwell.epub.txt
19 Christopher Golden Dead Ringer [deadringer] 2016 © the-eye.eu/public/Books/Bibliotik/C/Christopher Golden - Dead Ringers [retail].epub.txt
20 Christopher Golden Ararat [ararat] 2017 © the-eye.eu/public/Books/Bibliotik/A/Ararat - Christopher Golden.epub.txt
21 Andrew Sean Greer The Confessions of Max Tivoli [maxtivoli] 2005 © the-eye.eu/public/Books/Bibliotik/T/The Confessions of Max Tivoli - Andrew Sean Greer.epub.txt
22 John Grisham Theodore Boone: The Fugitive [fugitive] 2015 © the-eye.eu/public/Books/Bibliotik/T/The Fugitive - John Grisham.epub.txt
23 Mark Haddon The Curious Incident of the Dog in the Night-Time [curious] 2003 © the-eye.eu/public/Books/Bibliotik/T/The Curious Incident of the Dog in the Night-Time - Mark Haddon.epub(1).txt
24 Joseph Heller Catch-22 [catch22] 2018 © the-eye.eu/public/Books/Bibliotik/C/Catch-22 - Joseph Heller.epub.txt
25 David Henry Hwang M. Butterfly [mbutterfly] 1988 © the-eye.eu/public/Books/Bibliotik/M/M. Butterfly - David Henry Hwang.epub.txt
26 Betty E.M. Jacobs All the Onions [onions] 1999 © the-eye.eu/public/Books/Bibliotik/A/All the Onions - Betty E. M. Jacobs.epub.txt
27 James Joyce Ulysses [ulysses] 1922 PD the-eye.eu/public/Books/Bibliotik/U/Ulysses - James Joyce - Penguin Group - 2000.epub.txt
28 Richard Kadrey Sandman Slim [sandmanslim] 2009 © the-eye.eu/public/Books/Bibliotik/S/Sandman Slim - Richard Kadrey.epub.txt
29 Matthew Klam Who Is Rich? [rich] 2017 © the-eye.eu/public/Books/Bibliotik/W/Who Is Rich_ - Matthew Klam.epub.txt
30 Laura Lippman After I’m Gone [afterimgone] 2014 © the-eye.eu/public/Books/Bibliotik/A/After I’m Gone - Laura Lippman.epub.txt
31 Laura Lippman Sunburn [sunburn] 2018 © the-eye.eu/public/Books/Bibliotik/S/Sunburn - Laura Lippman.epub.txt
32 George R.R. Martin A Game of Thrones [thrones] 1996 © the-eye.eu/public/Books/Bibliotik/A/A Game of Thrones - George R. R. Martin.epub.txt
33 Stephenie Meyer Twilight [twilight] 2005 © the-eye.eu/public/Books/Bibliotik/T/Twilight - Stephenie Meyer.epub.txt
34 Toni Morrison Beloved [beloved] 1987 © the-eye.eu/public/Books/Bibliotik/B/Beloved - Toni Morrison.epub.txt
35 Yōko Ogawa The Memory Police [memory] 2019 © the-eye.eu/public/Books/Bibliotik/T/The Memory Police - Yoko Ogawa.epub.txt
36 George Orwell Nineteen-Eighty Four [1984] 1949 © the-eye.eu/public/Books/Bibliotik/N/Nineteen Eighty-Four (The Annotated Edition) - George Orwell.epub.txt
37 Philip Pullman The Subtle Knife [subtleknife] 1997 © the-eye.eu/public/Books/Bibliotik/P/Pullman - The Subtle Knife.epub.txt
38 Ramzi Rouighi The Making of a Mediterranean Emirate [emirate] 2011 © the-eye.eu/public/Books/Bibliotik/T/The Making of a Mediterranean E - Rouighi, Ramzi;.epub.txt
39 J.K. Rowling Harry Potter and the Sorcerer’s Stone [hp1] 1998 © the-eye.eu/public/Books/Bibliotik/H/Harry_Potter_ and_the_Sorcerers_Stone-Rowling.epub.txt
40 J.K. Rowling Harry Potter and the Goblet of Fire [hp4] 2000 © the-eye.eu/public/Books/Bibliotik/H/Harry_Potter_ and_the_Goblet_of_Fire-Rowling.epub.txt
41 J.D. Salinger The Catcher in the Rye [catcher] 1951 © the-eye.eu/public/Books/Bibliotik/T/The Catcher in the Rye - J. D. Salinger.epub.txt
42 Sheryl Sandberg Lean In [leanin] 2013 © the-eye.eu/public/Books/Bibliotik/L/Lean In - Sheryl Sandberg.epub.txt
43 Sarah Silverman The Bedwetter [bedwetter] 2010 © the-eye.eu/public/Books/Bibliotik/T/The Bedwetter - Sarah Silverman.epub.txt
44 Rachel Louise Snyder No Visible Bruises [novisible] 2019 © the-eye.eu/public/Books/Bibliotik/N/No Visible Bruises - Rachel Louise Snyder.epub.txt
45 Lysa TerKeurst Unglued [unglued] 2012 © the-eye.eu/public/Books/Bibliotik/L/Lysa TerKeurst, Unglued.epub.txt
46 Lysa TerKeurst Embraced [embraced] 2018 © the-eye.eu/public/Books/Bibliotik/E/Embraced - Lysa Terkeurst.epub.txt
47 J.R.R. Tolkien The Hobbit [hobbit] 1937 © the-eye.eu/public/Books/Bibliotik/T/The Hobbit (Houghton Mifflin Harcourt) (75th Anniversary Edition) [Epub] - J.R.R. Tolkien.epub.txt
48 Jacqueline Woodson Brown Girl Dreaming [browngirldreaming] 2014 © the-eye.eu/public/Books/Bibliotik/B/Brown Girl Dreaming - Jacqueline Woodson.epub.txt
49 Jacqueline Woodson Another Brooklyn [anotherbrooklyn] 2016 © the-eye.eu/public/Books/Bibliotik/A/Another Brooklyn - Jacqueline Woodson.epub.txt
50 Jonathan Zittrain The Future of the Internet and How to Stop It [zittrain] 2008 © the-eye.eu/public/Books/Bibliotik/J/Jonathan Zittrain - The Future of the Internet.epub.txt
D.2 Sliding-window probabilistic extraction procedure
In this appendix, we provide more details on our main “panning for gold” extraction experiments (Section 4). These experiments take the approach of constructing $100$-token sequences $\bm{z}$} out of Books3 books, which we attempt to extract. These sequences are constructed using a sliding window across the length of an entire book, and we run these experiments for each of the $50$ books in Table LABEL:app:tab:sliding-books. We use a very short, $10$-character shift for our sliding window. This is deliberate; it is meant to help us, effectively, “pan for gold”—to identify regions within specific books where there are high-probability stretches of memorized content. The approach that we take reveals the position (in character space) in each book where these regions occur (and if they occur at all). We explain the intention behind these experiments, and how the associated results should be interpreted.
Sequence length.
We use $100$-token sequences $\bm{z}$}, with a prefix (prompt) length of $50$ tokens and a (target) suffix length of $50$ tokens. We pick $100$ tokens and a 5050/$50$ split because this is accepted in the literature as a reasonable minimum for being confident that extracted suffixes are reflective of memorization [carlini2023quantifying, carlini2021extracting, lee2022dedup, hayes2025measuringmemorizationlanguagemodels]. $50$ tokens for the prefix is less for validity reasons. Of course, this prefix length uses an amount of context that is equivalent to the output, and so is compelling in this respect. However, when we are confident that our measurements are capturing valid instances of extraction, using longer prefixes can help discover more underlying memorization. (Indeed, this is part of where discoverable extraction gets its name; carlini2023quantifying demonstrate the “discoverability phenomenon” with respect to showing how extraction success can increase with longer-context prompts.) We discuss this further (and include experiments with varied prefix lengths) in Section 5 and Appendix E. Other work uses shorter sequences (e.g., $64$ tokens and a 3232/$32$ token split [biderman2023pythia]).
Sliding-window sequence chunking.
For a given book, we start at the beginning of the text file in Books3. We sample a chunk of text that is sufficiently long to contain $100$ tokens of corresponding tokenized text: to do this, we take a chunk of $800$ characters, tokenize, take the first $100$ tokens of the resulting tokenized sequence as the sequence, and discard the rest. (Each $100$ token sequence is typically, but not always, reflective of 300300-$400$ characters of text. Rare words and formatting make this vary considerably. This is why we use a character-chunk size of 800800—to make sure we have sufficient head-room to always end up with a $100$ token sequence.) We then shift $10$ characters forward in the book text and repeat this process. We do this for the entire length of the book, which results in approximately one sequence every $10$ characters, i.e., $\texttt{len(book\_characters)}/10$ total sequences $\bm{z}$}, for which we compute $p_{\bm{z}}$. For example, The Great Gatsby has 270,870270{,}870 characters, which results in roughly 27,00027{,}000 sequences, for which we compute $p_{\bm{z}}$ (Equation A.2). (See Section 2 and Appendix A.2.)
This means that the $100$-token sequences overlap significantly. This is deliberate. We typically do not know how different open-weight models were trained; a priori, it is not clear exactly where we should begin sequences (or where they should end). With this approach, for extractable content, we expect to surface high-probability regions of memorized content, which we can then explore in more detail and more precisely in follow-up experiments. We picked a sliding window of $10$ characters by running initial experiments with $1$-, $10$-, $50$-, and $100$-character windows. $10$ characters exhibited an insignificant loss in extraction signal (at $10\times$ cheaper cost than $1$ character). We intend to investigate reducing this cost further in the future.
Initially, we did not expect this approach to result in high per-book extraction rates. This is because, even though this approach should surface memorization “hot-spots” by getting good coverage of $100$-token sequences across the book, we are not being particularly clever about the start location for each sequence. Since we indiscriminately slide the window $10$ characters each time, we expect that many prompts will begin in “odd” locations—locations in the middle of tokens, which lead to “broken” or otherwise “strange” tokenizations. We expected to yield near-zero probability of extraction.
However, with the final presentation of our results—in which we use a very high minimum probability $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$ in this work for deeming extraction successful (Sections 2.1 & 5; Appendices C & E)—this did not really have an impact on extraction rates.
So, in the subsections that follow, we show per-book extraction rates across the different models we test with this procedure. These are useful for very high-level comparisons of how many sequences we extracted with probabilistic extraction, compared to greedy-decoded discoverable extraction as a baseline (which, similar to hayes2025measuringmemorizationlanguagemodels, we observe consistently under-counts extraction).
Decoding configuration
Similar to hayes2025measuringmemorizationlanguagemodels, after some initial tests with different decoding schemes, we use temperature T=1T{=}1 for top-$k$ decoding with the k=40k{=}40. We believe that setting T=1T{=}1 makes the most sense for studying memorization, as this reflects the LLM’s base probability distribution. Since extraction is only successful when the underlying training sequence’s logits reflect very high probabilities (Section 2), we also think it makes sense to clip the full-vocabulary logit distribution to reflect some number of the top-highest probability tokens per step. This is the same intuition behind using greedy decoding for discoverable extraction—but with a twist. Greedy decoding deterministically picks the single-most-likely token per step. (This is equivalent to top-$k$ decoding with T=1T{=}1 and k=1k{=}1.) As the results in hayes2025measuringmemorizationlanguagemodels show, this can be a suboptimal decoding strategy for measuring extraction, for several reasons:
  • Greedy decoding is rarely used with LLMs in practice. Non-deterministic sampling schemes are more reflective of realistic conditions.
  • Because greedy decoding is deterministic, the same prompt always results in the same output. As a result, greedy-decoded discoverable extraction provides (in this sense) $1$-bit of information with respect to extraction: was it possible to extract the sequence? Using non-deterministic decoding schemes, we can get more detailed information about extraction risk; not all sequences are equally extractable, and this is useful information to analyze. (See Section 2.2 & Appendix A.)
  • While greedy decoding picks the locally highest probability token per step, it is not guaranteed to result in the globally highest probability sequence (of the given length) under the model. Being able to explore more of the model’s distribution with non-deterministic decoding schemes can surface sequences (of the same given length) that are higher probability under the model. Such higher probability sequences can match the target suffix, where the greedy-decoded suffix does not (i.e., fails to extract that suffix).
  • More generally, this allows for a bit more “wiggle room.” By being able to pick beyond just the top-$1$ token per step, the decoding process take locally suboptimal steps, which can still result in high-probability sequences that (even if slightly lower probability than the greedy sequence, with respect to the normalized top-$k$ distribution) reflect extraction. While we do not investigate this in detail in this work, we observe that (when this happens) such sequences are very similar to the greedy-decoded suffix; the greedy-decoded suffix is a near-miss for the verbatim suffix (it is near-verbatim). It may be reasonable under more relaxed (near-verbatim) metrics for extraction to have counted this greedy suffix as extracted, but that is not the approach we (or hayes2025measuringmemorizationlanguagemodels) take, and so these suffixes would be missed. We intend to investigate this further in future work.
The above reasons also reflect the general motivations for using probabilistic extraction (rather than traditional discoverable extraction) as our metric (Sections 2 & 3; Appendix A). Similar motivations for useful extraction measurements are also noted in older extraction papers (e.g., carlini2021extracting, carlini2023quantifying); however, this work typically opts to use greedy decoding based on the perception that it is cheaper to do so at scale. This work does not claim that greedy decoding is the best or only way to measure extraction, and in fact notes [carlini2021extracting]—and in select cases shows [carlini2023quantifying]—that other (more expensive) decoding schemes are able to extract more training data. (Again, this seems likely because these schemes, like beam search [carlini2023quantifying], are able to produce higher-likelihood sequences under the model.)
Further, as we show through optimizing the work of hayes2025measuringmemorizationlanguagemodels, it turns out that it is actually much cheaper to compute probabilistic extraction (with any top-$k$ decoding configuration, including greedy/ k=1k{=}1), as we compute extraction probabilities as a post-processing operation directly off of the logits obtained from inference, rather than (as has been done in prior work [carlini2021extracting, carlini2023quantifying]) actually generating (greedy) completions. We can compute verbatim extraction probabilities with a single forward pass through the model, which is an enormous cost savings compared to generating $50$ tokens per prompt. We discuss these details in Appendix A.3.
Note that because our implementation of probabilistic extraction makes computing $\log$ probabilities a post-processing operation on the output model logits on CPU (Appendix A), we also collect results for greedy-decoded discoverable extraction [lee2022dedup, carlini2021extracting, nasr2023scalable] by setting T=k=1T{=}k{=}1, incurring negligible cost to do so. For brevity, we do not include detailed results for this setting. To give a brief intuition for these results, we do include greedy extraction rates (with respect to the set of sliding-window sequences we construct) for each book’s results in Appendix D.3; these rates consistently under-count probabilistic extraction with top-$k$ decoding (T=1T{=}1, k=40k{=}40).
D.3 Book-specific results
For each of the $50$ books we test, we include a separate subsection with the following information for $10$ models (documented in Appendix D.1):
  • Bar plot comparing extraction and generation rates. For each model, we plot three different rates: the greedy-decoded discoverable extraction rate; the probabilistic extraction rate for top-$k$ decoding (T=1T{=}1, k=40k{=}40), counting a sequence as extracted if $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$; and the maximum possible rate of generating suffixes given their corresponding prefixes. Each rate is computed over all (overlapping) sequences from the sliding-window procedure (Appendix D.2). Because we use a consistent sequence set, this comparison is valid even with overlaps.
    The discoverable extraction rate consistently underestimates the probabilistic extraction rate, even with this conservative $\tau_{\text{min}}$. The maximum rate of suffix generation is not an extraction rate: here we do not apply $\tau_{\text{min}}$ (effectively setting it to $0\%$), so it includes all sequences with nonzero $p_{\bm{z}}$. This captures not only valid extractions but also low-probability instances of spontaneous suffix generation given a prefix. We include this number for two reasons:
    1. 1.
      It is never $100\%$, showing that not every suffix can be generated given its prefix, even in the limit of repeated prompts. This is important to emphasize for legal audiences, where there is a common misconception that, with enough queries, a probabilistic model can generate anything [toothpick]. Our results show that this is not the case.
    2. 2.
      We are conservative in globally setting $\tau_{\text{min}}=0.1\%$ to define extraction success. In future work, we plan to investigate more rigorous statistical methods for setting this threshold. As a result, our paper significantly under-counts extraction: the true amount lies above what we claim but below the maximum generation rate. Therefore, the maximum possible rate should be understood as a loose upper bound on extraction.
    We return to these points when discussing the validity of our measurement procedure (Section 5, Appendix E).
  • Summary table. This table reports exact values for the plotted rates, along with the runtime for the sliding-window probabilistic extraction procedure for each model.
  • Histograms of extraction probabilities. We plot the distribution over suffix extraction probabilities across all overlapping sequences from the sliding-window procedure. These histograms illustrate how many sequences are extractable for a given model and how their extraction probabilities vary.
  • Heatmaps of extraction probabilities by location. We also include heatmaps (as in Figure 3(b) and Section 6). At each character, the heatmap plots the maximum extraction probability among overlapping $50$-token suffixes spanning that character. These visualizations highlight “hot-spots” of high-probability extraction and show how such regions emerge and taper off within a book.
If no sequences are extracted for a given model–book pair, we omit histograms (they would be empty). In one case—The Beautiful Struggle [struggle]—we extract nothing for any model, so both histograms and heatmaps are omitted. For histograms and heatmaps, results are presented in three groups corresponding to the categories of models described in Appendix D.1: models known with certainty to have been trained on Books3 (blue), models that were likely trained on books also present in Books3 (red), and one model (Phi 4) that, based on documentation, was not trained on whole copyrighted books (orange). See Section 5 and Appendix E for further discussion about how we can be confident that our results for all three categories reflect valid extraction.
Tracking down longer extracted sequences.
Visualizing memorization “hot-spots” also allows us to identify runs of memorized text from which we can attempt to extract more than $50$ tokens. That is, in regions with very high extraction probabilities, these probabilities may be high enough to yield nontrivial extraction probabilities for longer runs.
We take the full list of $p_{\bm{z}}$ for all sequences in a given book–model experiment, along with their start and end positions, sorted by start position. We filter to only those with especially high probability (e.g., $p_{\bm{z}}\geq 50\%$), then merge overlapping sequences by taking the minimum start and maximum end. From this we construct longer candidate sequences to attempt extraction.
To illustrate why this works, consider a single $50$-token suffix with $p_{\bm{z}}=50\%$. Its per-token geometric mean probability is $0.5^{1/50}\approx 98.62\%$. For this individual token probability, the longest sequence length $x$ that maintains $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$ is
$(0.5^{1/50})^{x}=0.001\;\;\Longrightarrow\;\;x=\frac{\log 0.001}{\log(0.5^{1/50})}\approx 498\text{ tokens}.$
(However, for a suffix of this length, it would perhaps be reasonable to consider a smaller $\tau_{\text{min}}$ for determining extraction success.) If we see multiple sequences close by that exhibit $p_{\bm{z}}=50\%$, we could perhaps find a much longer sequence that fits the conditions above (or something similar).
We applied this methodology to Harry Potter and the Sorcerer’s Stone [hp1] and 1984 [1984] with Llama 3.1 70B, given the extent of memorization we observe in those cases. Figure 13 shows one successful case: a $300$-token extraction consisting of a $25$-token prefix and a $275$-token suffix with $p_{\bm{z}}=36.1\%$.
Figure 13: An example of a long sequence from Harry Potter and the Sorcerer’s Stone [hp1] that we extract from Llama 3.1 70B, with top-$k$ (T=1T{=}1 k=40k{=}40) decoding. The prompt is a $25$-token prefix; the target suffix is $275$ tokens. This suffix has $p_{\bm{z}}=36.1\%$, which corresponds to a geometric-mean individual token conditional probability of over $99.6\%$, i.e., $0.9963^{275}\approx 36.1\%$). In expectation, with this prompt, this suffix is generated verbatim more than $\nicefrac{{1}}{{3}}$ of the time.
D.3.1 The Handmaid’s Tale, handmaid
The Handmaid’s Tale
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.02% 0.04% 0.15% 00:15:12.23
Llama 1 13B 0.02% 0.07% 0.35% 00:16:32.97
Llama 1 65B 0.24% 0.54% 2.03% 00:56:07.73
Llama 2 70B 0.30% 0.74% 2.45% 00:56:03.78
Llama 3 70B 0.33% 1.14% 6.37% 01:08:44.77
Llama 3.1 70B 0.66% 1.70% 8.08% 01:08:14.35
DeepSeek v1 67B 0.01% 0.06% 0.30% 01:14:59.54
Qwen 2.5 72B 0.02% 0.09% 0.86% 01:14:45.41
Gemma 2 27B 0.01% 0.06% 0.36% 01:14:34.15
Phi 4 0.01% 0.06% 0.23% 00:26:39.96
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 14: The Handmaid’s Tale, handmaid: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 15: Heatmaps according to location in The Handmaid’s Tale. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.2 Dante and the Origins of Italian Culture, dante
Dante and the Origins of Italian Culture
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.00% 0.01% 0.04% 01:09:53.43
Llama 1 13B 0.00% 0.00% 0.09% 00:50:57.34
Llama 1 65B 0.00% 0.00% 0.20% 02:20:26.40
Llama 2 70B 0.00% 0.00% 0.27% 02:06:17.05
Llama 3 70B 0.01% 0.04% 1.06% 02:32:48.51
Llama 3.1 70B 0.02% 0.07% 1.26% 02:32:13.24
DeepSeek v1 67B 0.00% 0.01% 0.22% 02:33:57.80
Qwen 2.5 72B 0.00% 0.01% 0.19% 02:52:29.34
Gemma 2 27B 0.00% 0.01% 0.10% 02:38:54.09
Phi 4 0.00% 0.01% 0.18% 02:18:03.72
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 16: Dante and the Origins of Italian Culture, dante: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 17: Heatmaps according to location in Dante and the Origins of Italian Culture. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.3 The Da Vinci Code, davinci
The Da Vinci Code
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.15% 0.17% 0.26% 00:25:41.35
Llama 1 13B 0.16% 0.20% 0.47% 00:26:15.45
Llama 1 65B 0.21% 0.35% 1.56% 01:28:07.24
Llama 2 70B 0.24% 0.51% 2.06% 01:29:53.68
Llama 3 70B 0.39% 1.01% 17.61% 01:49:20.95
Llama 3.1 70B 0.54% 1.61% 21.80% 01:50:22.05
DeepSeek v1 67B 0.17% 0.23% 0.87% 01:55:44.93
Qwen 2.5 72B 0.17% 0.24% 0.97% 02:14:58.24
Gemma 2 27B 0.16% 0.18% 0.39% 01:54:16.65
Phi 4 0.16% 0.17% 0.22% 00:45:56.32
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 18: The Da Vinci Code, davinci: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 19: Heatmaps according to location in The Da Vinci Code. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.4 The Myth of Sisyphus, sisyphus
The Myth of Sisyphus
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.45% 0.65% 0.81% 00:10:47.48
Llama 1 13B 0.48% 0.77% 2.03% 00:11:17.22
Llama 1 65B 0.68% 1.86% 6.30% 00:38:00.27
Llama 2 70B 0.86% 2.35% 8.62% 00:38:46.04
Llama 3 70B 2.15% 5.49% 21.17% 00:46:57.15
Llama 3.1 70B 2.50% 6.80% 22.78% 00:47:50.27
DeepSeek v1 67B 0.45% 1.00% 2.32% 00:56:29.95
Qwen 2.5 72B 0.79% 1.99% 6.38% 01:03:05.52
Gemma 2 27B 0.49% 0.99% 2.29% 00:56:54.18
Phi 4 0.47% 0.79% 1.76% 00:22:30.60
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 20: The Myth of Sisyphus, sisyphus: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 21: Heatmaps according to location in The Myth of Sisyphus. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.5 Alice’s Adventures in Wonderland, alice
Alice’s Adventures in Wonderland
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.16% 0.40% 10.40% 00:42:22.36
Llama 1 13B 1.49% 4.81% 26.56% 00:32:19.81
Llama 1 65B 4.96% 16.92% 46.52% 01:26:03.29
Llama 2 70B 5.19% 17.09% 46.08% 01:17:43.11
Llama 3 70B 6.76% 21.80% 49.74% 01:34:00.43
Llama 3.1 70B 7.48% 23.30% 50.32% 01:35:36.94
DeepSeek v1 67B 4.52% 15.55% 46.19% 01:39:10.60
Qwen 2.5 72B 4.02% 14.18% 45.95% 01:52:45.29
Gemma 2 27B 2.57% 8.18% 30.32% 01:36:46.76
Phi 4 0.65% 1.82% 14.90% 01:24:06.58
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 22: Alice’s Adventures in Wonderland, alice: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 23: Heatmaps according to location in Alice’s Adventures in Wonderland. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.6 The Beautiful Struggle, struggle
The Beautiful Struggle
Model Greedy $\tau_{\text{min}}{=}0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.00% 0.00% 0.03% 00:19:22.45
Llama 1 13B 0.00% 0.00% 0.08% 00:13:44.55
Llama 1 65B 0.00% 0.00% 0.11% 00:37:14.86
Llama 2 70B 0.00% 0.00% 0.10% 00:31:25.61
Llama 3 70B 0.00% 0.00% 0.14% 00:38:11.48
Llama 3.1 70B 0.00% 0.00% 0.17% 00:39:49.66
DeepSeek v1 67B 0.00% 0.00% 0.10% 00:46:24.68
Qwen 2.5 72B 0.00% 0.00% 0.07% 00:53:11.56
Gemma 2 27B 0.00% 0.00% 0.10% 00:42:27.90
Phi 4 0.00% 0.00% 0.03% 00:39:00.05
Figure 24: Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
We omit heatmaps since they are completely empty for this book. The are no probabilities for any model that satisfy $\tau_{\text{min}}\geq 0.1\%$.
D.3.7 We Were Eight Years in Power, wewere
We Were Eight Years in Power
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.42% 0.48% 0.73% 00:21:20.90
Llama 1 13B 0.43% 0.60% 1.38% 00:21:48.74
Llama 1 65B 0.80% 1.94% 8.64% 01:13:02.59
Llama 2 70B 0.80% 1.81% 8.38% 01:13:58.30
Llama 3 70B 1.24% 2.60% 10.20% 01:29:38.62
Llama 3.1 70B 1.51% 3.21% 10.94% 01:29:39.20
DeepSeek v1 67B 0.45% 0.59% 1.39% 01:36:56.81
Qwen 2.5 72B 0.65% 1.25% 4.45% 01:50:31.25
Gemma 2 27B 0.41% 0.57% 1.14% 01:36:05.07
Phi 4 0.41% 0.44% 0.62% 00:39:18.74
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 25: We Were Eight Years in Power, wewere: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 26: Heatmaps according to location in We Were Eight Years in Power. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.8 The Water Dancer, waterdancer
The Water Dancer
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.46% 0.52% 0.71% 00:21:44.11
Llama 1 13B 0.49% 0.53% 0.67% 00:22:15.13
Llama 1 65B 0.49% 0.54% 0.76% 01:16:48.71
Llama 2 70B 0.49% 0.54% 0.83% 01:16:49.47
Llama 3 70B 0.47% 0.53% 0.86% 01:31:46.86
Llama 3.1 70B 0.49% 0.56% 0.89% 01:32:59.63
DeepSeek v1 67B 0.49% 0.57% 0.88% 01:42:56.38
Qwen 2.5 72B 0.50% 0.54% 0.81% 01:56:53.50
Gemma 2 27B 0.49% 0.53% 0.66% 01:35:49.67
Phi 4 0.46% 0.51% 0.61% 00:40:49.06
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 27: The Water Dancer, waterdancer: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 28: Heatmaps according to location in The Water Dancer. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.9 Dungeons and Dragons and Philosophy, dungeons
Dungeons and Dragons and Philosophy
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.09% 0.19% 0.94% 00:22:04.42
Llama 1 13B 0.04% 0.12% 1.21% 00:20:43.58
Llama 1 65B 0.09% 0.46% 2.02% 01:12:37.41
Llama 2 70B 0.17% 0.61% 2.21% 01:14:42.20
Llama 3 70B 0.21% 0.65% 2.42% 01:29:40.16
Llama 3.1 70B 0.22% 0.69% 2.45% 01:26:12.17
DeepSeek v1 67B 0.15% 0.62% 1.96% 01:38:55.59
Qwen 2.5 72B 0.08% 0.22% 1.72% 01:33:21.08
Gemma 2 27B 0.04% 0.11% 0.93% 01:28:45.86
Phi 4 0.02% 0.10% 0.64% 00:35:24.35
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 29: Dungeons and Dragons and Philosophy, dungeons: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 30: Heatmaps according to location in Dungeons and Dragons and Philosophy. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.10 Drown, drown
Drown
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.00% 0.00% 0.02% 00:15:39.60
Llama 1 13B 0.00% 0.00% 0.05% 00:11:21.27
Llama 1 65B 0.00% 0.00% 0.15% 00:30:33.16
Llama 2 70B 0.03% 0.06% 0.33% 00:26:45.87
Llama 3 70B 0.08% 0.09% 0.74% 00:33:39.80
Llama 3.1 70B 0.03% 0.09% 0.69% 00:32:19.88
DeepSeek v1 67B 0.00% 0.00% 0.20% 00:42:57.33
Qwen 2.5 72B 0.00% 0.00% 0.13% 00:42:54.30
Gemma 2 27B 0.00% 0.00% 0.07% 00:34:08.02
Phi 4 0.00% 0.00% 0.01% 00:31:49.62
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 31: Drown, drown: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 32: Heatmaps according to location in Drown. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.11 The Brief Wondrous Life of Oscar Wao, oscarwao
The Brief Wondrous Life of Oscar Wao
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.09% 0.16% 0.21% 00:17:12.42
Llama 1 13B 0.04% 0.15% 0.25% 00:16:17.08
Llama 1 65B 0.10% 0.23% 0.34% 00:55:33.06
Llama 2 70B 0.08% 0.29% 0.58% 00:55:40.12
Llama 3 70B 0.14% 0.38% 1.01% 01:08:54.55
Llama 3.1 70B 0.16% 0.45% 1.02% 01:07:45.63
DeepSeek v1 67B 0.06% 0.19% 0.31% 01:12:30.63
Qwen 2.5 72B 0.06% 0.19% 0.34% 01:20:54.30
Gemma 2 27B 0.03% 0.12% 0.30% 01:13:44.72
Phi 4 0.01% 0.00% 0.25% 00:27:47.94
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 33: The Brief Wondrous Life of Oscar Wao, oscarwao: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 34: Heatmaps according to location in The Brief Wondrous Life of Oscar Wao. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.12 This Is How You Lose Her, loseher
This Is How You Lose Her
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.22% 0.31% 0.35% 00:07:29.38
Llama 1 13B 0.07% 0.28% 0.33% 00:07:57.88
Llama 1 65B 0.16% 0.35% 0.57% 00:27:38.38
Llama 2 70B 0.11% 0.38% 0.75% 00:27:25.49
Llama 3 70B 0.12% 0.34% 0.71% 00:33:17.95
Llama 3.1 70B 0.08% 0.39% 0.71% 00:33:24.74
DeepSeek v1 67B 0.16% 0.32% 0.38% 00:41:50.38
Qwen 2.5 72B 0.15% 0.31% 0.34% 00:37:56.86
Gemma 2 27B 0.06% 0.27% 0.34% 00:38:33.11
Phi 4 0.00% 0.00% 0.30% 00:12:51.69
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 35: This Is How You Lose Her, loseher: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 36: Heatmaps according to location in This Is How You Lose Her. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.13 Down and Out in the Magic Kingdom, downandout
Down and Out in the Magic Kingdom
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.06% 0.18% 0.32% 00:16:40.70
Llama 1 13B 0.01% 0.06% 0.18% 00:12:37.05
Llama 1 65B 0.03% 0.09% 0.49% 00:33:25.06
Llama 2 70B 0.04% 0.10% 0.42% 00:29:13.97
Llama 3 70B 0.06% 0.15% 1.41% 00:36:03.83
Llama 3.1 70B 0.10% 0.18% 1.60% 00:35:05.98
DeepSeek v1 67B 0.12% 0.21% 0.51% 00:46:46.42
Qwen 2.5 72B 0.02% 0.18% 0.39% 00:51:43.93
Gemma 2 27B 0.01% 0.07% 0.17% 00:42:41.89
Phi 4 0.00% 0.00% 0.12% 00:35:27.50
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 37: Down and Out in the Magic Kingdom, downandout: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 38: Heatmaps according to location in Down and Out in the Magic Kingdom. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.14 The World’s Wife, worldswife
The World’s Wife
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.46% 0.86% 1.14% 00:02:16.06
Llama 1 13B 0.17% 0.29% 0.86% 00:02:01.06
Llama 1 65B 0.05% 0.52% 2.43% 00:07:50.11
Llama 2 70B 0.14% 0.52% 2.82% 00:07:47.68
Llama 3 70B 0.95% 1.92% 6.81% 00:08:56.11
Llama 3.1 70B 0.68% 1.87% 5.90% 00:09:29.81
DeepSeek v1 67B 0.32% 0.83% 1.65% 00:20:36.99
Qwen 2.5 72B 0.22% 0.56% 1.70% 00:10:05.30
Gemma 2 27B 0.00% 0.21% 1.60% 00:17:30.78
Phi 4 0.10% 0.11% 0.90% 00:04:44.13
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 39: The World’s Wife, worldswife: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 40: Heatmaps according to location in The World’s Wife. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.15 A Visit from the Goon Squad, goonsquad
A Visit from the Goon Squad
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.00% 0.04% 0.12% 00:13:21.36
Llama 1 13B 0.00% 0.00% 0.11% 00:14:14.38
Llama 1 65B 0.02% 0.05% 0.32% 00:49:25.39
Llama 2 70B 0.01% 0.05% 0.42% 00:49:29.54
Llama 3 70B 0.08% 0.15% 0.73% 00:59:39.01
Llama 3.1 70B 0.10% 0.20% 0.72% 00:59:32.69
DeepSeek v1 67B 0.00% 0.04% 0.21% 01:06:18.04
Qwen 2.5 72B 0.00% 0.05% 0.24% 01:03:48.02
Gemma 2 27B 0.00% 0.03% 0.17% 01:05:09.63
Phi 4 0.00% 0.02% 0.13% 00:25:50.75
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 41: A Visit from the Goon Squad, goonsquad: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 42: Heatmaps according to location in A Visit from the Goon Squad. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.16 The President’s Vampire, vampire
The President’s Vampire
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.13% 0.29% 0.38% 00:32:58.48
Llama 1 13B 0.10% 0.16% 0.32% 00:23:26.33
Llama 1 65B 0.10% 0.23% 0.39% 01:01:31.94
Llama 2 70B 0.11% 0.24% 0.39% 00:52:13.17
Llama 3 70B 0.09% 0.21% 0.38% 01:02:45.72
Llama 3.1 70B 0.10% 0.24% 0.37% 01:03:01.58
DeepSeek v1 67B 0.14% 0.30% 0.45% 01:11:22.92
Qwen 2.5 72B 0.10% 0.25% 0.37% 01:17:12.82
Gemma 2 27B 0.08% 0.12% 0.31% 01:09:01.82
Phi 4 0.07% 0.07% 0.26% 01:02:03.59
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 43: The President’s Vampire, vampire: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 44: Heatmaps according to location in The President’s Vampire. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.17 The Great Gatsby, gatsby
The Great Gatsby
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.16% 0.36% 0.91% 00:20:05.24
Llama 1 13B 1.33% 3.31% 21.53% 00:13:07.86
Llama 1 65B 8.22% 21.74% 48.99% 00:34:03.40
Llama 2 70B 7.26% 20.66% 51.06% 00:29:51.29
Llama 3 70B 12.33% 30.43% 64.90% 00:35:47.91
Llama 3.1 70B 16.11% 35.33% 67.86% 00:36:24.40
DeepSeek v1 67B 2.18% 5.56% 26.94% 00:46:31.71
Qwen 2.5 72B 3.24% 7.90% 29.73% 00:52:35.89
Gemma 2 27B 2.34% 7.21% 31.88% 00:44:44.88
Phi 4 1.00% 2.44% 6.68% 00:36:12.82
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 45: The Great Gatsby, gatsby: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 46: Heatmaps according to location in The Great Gatsby. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.18 Blink, blink
Blink
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.05% 0.19% 0.33% 00:14:29.08
Llama 1 13B 0.03% 0.12% 0.77% 00:13:33.43
Llama 1 65B 0.12% 0.44% 2.32% 00:48:36.54
Llama 2 70B 0.17% 0.59% 2.83% 00:48:40.22
Llama 3 70B 0.25% 1.19% 8.20% 00:59:12.29
Llama 3.1 70B 0.43% 1.65% 9.26% 00:59:11.72
DeepSeek v1 67B 0.11% 0.39% 1.10% 01:08:40.10
Qwen 2.5 72B 0.13% 0.47% 4.19% 01:02:51.21
Gemma 2 27B 0.04% 0.15% 0.89% 01:03:43.75
Phi 4 0.01% 0.07% 0.41% 00:24:16.36
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 47: Blink, blink: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 48: Heatmaps according to location in Blink. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.19 Dead Ringer, deadringer
Dead Ringer
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.11% 0.24% 0.43% 00:16:42.57
Llama 1 13B 0.11% 0.20% 0.45% 00:17:39.28
Llama 1 65B 0.10% 0.25% 0.48% 00:58:41.88
Llama 2 70B 0.12% 0.24% 0.50% 00:59:13.07
Llama 3 70B 0.09% 0.19% 0.42% 01:12:25.79
Llama 3.1 70B 0.09% 0.21% 0.43% 01:11:09.37
DeepSeek v1 67B 0.17% 0.32% 0.50% 01:20:36.57
Qwen 2.5 72B 0.11% 0.29% 0.48% 01:30:30.56
Gemma 2 27B 0.08% 0.20% 0.39% 01:20:40.55
Phi 4 0.04% 0.12% 0.26% 00:30:56.08
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 49: Dead Ringer, deadringer: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 50: Heatmaps according to location in Dead Ringer. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.20 Ararat, ararat
Ararat
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.14% 0.20% 0.33% 00:30:36.62
Llama 1 13B 0.12% 0.16% 0.34% 00:22:18.90
Llama 1 65B 0.13% 0.17% 0.42% 01:02:33.58
Llama 2 70B 0.13% 0.16% 0.41% 00:55:30.87
Llama 3 70B 0.09% 0.15% 0.44% 01:08:08.25
Llama 3.1 70B 0.09% 0.15% 0.45% 01:07:24.99
DeepSeek v1 67B 0.19% 0.26% 0.48% 01:07:58.66
Qwen 2.5 72B 0.16% 0.23% 0.39% 01:18:27.52
Gemma 2 27B 0.12% 0.16% 0.32% 01:11:25.61
Phi 4 0.08% 0.09% 0.14% 01:02:21.74
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 51: Ararat, ararat: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 52: Heatmaps according to location in Ararat. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.21 The Confessions of Max Tivoli, maxtivoli
The Confessions of Max Tivoli
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.00% 0.03% 0.08% 00:31:23.50
Llama 1 13B 0.00% 0.00% 0.04% 00:22:51.88
Llama 1 65B 0.00% 0.00% 0.07% 00:59:47.86
Llama 2 70B 0.00% 0.00% 0.08% 00:51:14.57
Llama 3 70B 0.00% 0.01% 0.10% 01:01:40.94
Llama 3.1 70B 0.00% 0.01% 0.10% 01:03:07.68
DeepSeek v1 67B 0.01% 0.04% 0.12% 01:09:46.77
Qwen 2.5 72B 0.00% 0.00% 0.06% 01:21:13.67
Gemma 2 27B 0.00% 0.00% 0.05% 01:09:55.07
Phi 4 0.00% 0.00% 0.00% 01:03:00.43
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 53: The Confessions of Max Tivoli, maxtivoli: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 54: Heatmaps according to location in The Confessions of Max Tivoli. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.22 The Fugitive, fugitive
The Fugitive
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.06% 0.12% 0.25% 00:08:14.74
Llama 1 13B 0.07% 0.12% 0.24% 00:08:12.61
Llama 1 65B 0.11% 0.14% 0.31% 00:30:03.36
Llama 2 70B 0.11% 0.14% 0.36% 00:30:37.72
Llama 3 70B 0.05% 0.12% 0.20% 00:37:01.28
Llama 3.1 70B 0.05% 0.12% 0.20% 00:37:03.15
DeepSeek v1 67B 0.10% 0.16% 0.42% 00:46:17.99
Qwen 2.5 72B 0.07% 0.16% 0.28% 00:40:28.49
Gemma 2 27B 0.07% 0.14% 0.25% 00:41:54.08
Phi 4 0.01% 0.08% 0.12% 00:15:42.06
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 55: The Fugitive, fugitive: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 56: Heatmaps according to location in The Fugitive. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.23 The Curious Incident of the Dog in the Night-Time, curious
The Curious Incident of the Dog in the Night-Time
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.03% 0.10% 0.32% 00:10:16.88
Llama 1 13B 0.06% 0.09% 0.59% 00:10:00.99
Llama 1 65B 0.10% 0.36% 2.39% 00:35:52.83
Llama 2 70B 0.13% 0.41% 3.08% 00:36:36.43
Llama 3 70B 0.63% 1.39% 10.39% 00:43:59.67
Llama 3.1 70B 0.60% 1.78% 11.98% 00:44:27.83
DeepSeek v1 67B 0.05% 0.15% 1.20% 00:53:18.28
Qwen 2.5 72B 0.13% 0.23% 1.80% 00:48:15.93
Gemma 2 27B 0.06% 0.08% 0.79% 00:53:41.68
Phi 4 0.03% 0.06% 0.49% 00:18:42.60
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 57: The Curious Incident of the Dog in the Night-Time, curious: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 58: Heatmaps according to location in The Curious Incident of the Dog in the Night-Time. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.24 Catch-22, catch22
Catch-22
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.03% 0.08% 0.23% 00:34:24.57
Llama 1 13B 0.07% 0.15% 0.61% 00:34:35.46
Llama 1 65B 0.38% 0.69% 2.50% 02:03:04.55
Llama 2 70B 0.36% 0.82% 2.87% 02:02:26.37
Llama 3 70B 1.03% 2.43% 17.62% 02:37:38.47
Llama 3.1 70B 1.56% 3.61% 21.10% 02:28:55.27
DeepSeek v1 67B 0.06% 0.16% 0.74% 02:30:53.45
Qwen 2.5 72B 0.13% 0.28% 0.87% 02:37:55.18
Gemma 2 27B 0.06% 0.15% 0.53% 02:29:00.40
Phi 4 0.03% 0.08% 0.26% 00:59:47.42
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 59: Catch-22, catch22: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 60: Heatmaps according to location in Catch-22. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.25 M. Butterfly, mbutterfly
M. Butterfly
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.23% 0.51% 1.34% 00:08:34.80
Llama 1 13B 0.07% 0.18% 1.14% 00:06:16.90
Llama 1 65B 0.15% 0.34% 1.34% 00:16:33.55
Llama 2 70B 0.11% 0.54% 1.91% 00:14:04.98
Llama 3 70B 0.11% 0.62% 2.19% 00:17:10.41
Llama 3.1 70B 0.19% 0.66% 2.34% 00:16:38.24
DeepSeek v1 67B 0.22% 0.74% 1.58% 00:27:01.18
Qwen 2.5 72B 0.12% 0.43% 1.17% 00:31:30.45
Gemma 2 27B 0.05% 0.22% 1.04% 00:25:35.94
Phi 4 0.02% 0.06% 0.71% 00:18:57.56
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 61: M. Butterfly, mbutterfly: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 62: Heatmaps according to location in M. Butterfly. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.26 All the Onions, onions
All the Onions
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.55% 0.92% 1.92% 00:02:02.96
Llama 1 13B 0.00% 0.02% 0.35% 00:02:08.65
Llama 1 65B 0.02% 0.15% 1.21% 00:07:37.66
Llama 2 70B 0.08% 0.23% 0.95% 00:07:28.70
Llama 3 70B 0.18% 0.68% 1.45% 00:09:17.09
Llama 3.1 70B 0.13% 0.71% 1.47% 00:08:58.09
DeepSeek v1 67B 0.21% 0.81% 2.47% 00:23:04.91
Qwen 2.5 72B 0.42% 0.76% 2.16% 00:09:31.26
Gemma 2 27B 0.06% 0.42% 1.47% 00:18:18.31
Phi 4 0.00% 0.00% 0.15% 00:07:00.32
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 63: All the Onions, onions: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 64: Heatmaps according to location in All the Onions. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.27 Ulysses, ulysses
Ulysses
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.03% 0.04% 0.17% 00:52:46.60
Llama 1 13B 0.03% 0.07% 0.28% 00:53:08.41
Llama 1 65B 0.13% 0.37% 1.59% 02:55:55.54
Llama 2 70B 0.75% 2.22% 17.20% 02:57:13.25
Llama 3 70B 6.30% 19.77% 62.76% 03:33:22.11
Llama 3.1 70B 10.01% 27.77% 66.53% 03:33:04.45
DeepSeek v1 67B 0.07% 0.17% 0.42% 03:40:03.91
Qwen 2.5 72B 0.12% 0.29% 1.05% 04:11:40.60
Gemma 2 27B 0.06% 0.19% 1.33% 03:19:29.68
Phi 4 0.07% 0.12% 0.29% 01:21:46.13
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 65: Ulysses, ulysses: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 66: Heatmaps according to location in Ulysses. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.28 Sandman Slim, sandmanslim
Sandman Slim
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.13% 0.18% 0.23% 00:35:55.95
Llama 1 13B 0.01% 0.07% 0.15% 00:25:07.35
Llama 1 65B 0.05% 0.12% 0.24% 01:07:04.32
Llama 2 70B 0.06% 0.12% 0.22% 00:56:59.89
Llama 3 70B 0.07% 0.12% 0.33% 01:09:59.60
Llama 3.1 70B 0.09% 0.14% 0.33% 01:10:34.01
DeepSeek v1 67B 0.13% 0.21% 0.34% 01:20:56.42
Qwen 2.5 72B 0.08% 0.20% 0.33% 01:28:06.86
Gemma 2 27B 0.06% 0.12% 0.20% 01:14:49.98
Phi 4 0.02% 0.07% 0.13% 01:09:51.93
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 67: Sandman Slim, sandmanslim: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 68: Heatmaps according to location in Sandman Slim. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.29 Who Is Rich, rich
Who Is Rich
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.64% 0.77% 0.90% 00:33:56.93
Llama 1 13B 0.68% 0.75% 0.88% 00:23:58.84
Llama 1 65B 0.69% 0.76% 0.96% 01:05:50.46
Llama 2 70B 0.68% 0.77% 0.94% 00:56:11.63
Llama 3 70B 0.60% 0.72% 0.85% 01:07:24.48
Llama 3.1 70B 0.60% 0.70% 0.85% 01:06:56.61
DeepSeek v1 67B 0.65% 0.81% 1.02% 01:17:45.64
Qwen 2.5 72B 0.62% 0.76% 0.99% 01:30:40.41
Gemma 2 27B 0.63% 0.72% 0.81% 01:16:21.60
Phi 4 0.59% 0.66% 0.71% 01:07:13.60
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 69: Who Is Rich, rich: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 70: Heatmaps according to location in Who Is Rich. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.30 After I’m Gone, afterimgone
After I’m Gone
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.13% 0.18% 0.23% 00:33:12.10
Llama 1 13B 0.00% 0.04% 0.21% 00:24:18.57
Llama 1 65B 0.01% 0.11% 0.34% 01:05:17.58
Llama 2 70B 0.00% 0.10% 0.36% 00:57:47.55
Llama 3 70B 0.05% 0.15% 0.37% 01:10:27.58
Llama 3.1 70B 0.04% 0.15% 0.37% 01:09:50.46
DeepSeek v1 67B 0.09% 0.19% 0.43% 01:18:42.58
Qwen 2.5 72B 0.08% 0.17% 0.35% 01:28:59.07
Gemma 2 27B 0.02% 0.10% 0.20% 01:07:40.31
Phi 4 0.00% 0.07% 0.14% 01:06:48.44
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 71: After I’m Gone, afterimgone: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 72: Heatmaps according to location in After I’m Gone. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.31 Sunburn, sunburn
Sunburn
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.10% 0.15% 0.20% 00:14:07.97
Llama 1 13B 0.12% 0.13% 0.32% 00:13:09.52
Llama 1 65B 0.12% 0.14% 0.35% 00:47:04.89
Llama 2 70B 0.12% 0.18% 0.35% 00:46:16.38
Llama 3 70B 0.14% 0.19% 0.36% 00:56:19.30
Llama 3.1 70B 0.12% 0.17% 0.36% 00:57:14.94
DeepSeek v1 67B 0.14% 0.17% 0.36% 01:08:03.62
Qwen 2.5 72B 0.14% 0.18% 0.26% 01:14:02.46
Gemma 2 27B 0.12% 0.16% 0.22% 01:01:05.27
Phi 4 0.09% 0.12% 0.17% 00:24:21.27
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 73: Sunburn, sunburn: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 74: Heatmaps according to location in Sunburn. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.32 A Game of Thrones, thrones
A Game of Thrones
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.52% 0.54% 0.70% 00:50:39.51
Llama 1 13B 0.54% 0.57% 0.85% 00:50:41.26
Llama 1 65B 0.73% 1.06% 3.64% 02:52:01.23
Llama 2 70B 0.70% 1.01% 3.65% 02:58:28.25
Llama 3 70B 3.21% 8.50% 49.39% 03:35:35.17
Llama 3.1 70B 3.53% 8.98% 49.00% 03:35:27.98
DeepSeek v1 67B 0.55% 0.59% 1.14% 03:33:39.71
Qwen 2.5 72B 0.55% 0.65% 1.65% 03:57:53.28
Gemma 2 27B 0.54% 0.58% 0.88% 03:15:31.02
Phi 4 0.51% 0.53% 0.63% 01:25:11.53
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 75: A Game of Thrones, thrones: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 76: Heatmaps according to location in A Game of Thrones. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.33 Twilight, twilight
Twilight
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.00% 0.00% 0.00% 00:21:00.06
Llama 1 13B 0.00% 0.00% 0.07% 00:23:01.77
Llama 1 65B 0.13% 0.33% 1.90% 01:13:54.80
Llama 2 70B 0.12% 0.28% 1.76% 01:15:17.73
Llama 3 70B 0.24% 0.81% 9.73% 01:30:38.57
Llama 3.1 70B 0.27% 0.93% 11.14% 01:31:17.49
DeepSeek v1 67B 0.00% 0.00% 0.14% 01:39:12.76
Qwen 2.5 72B 0.00% 0.00% 0.03% 01:54:47.95
Gemma 2 27B 0.00% 0.02% 0.08% 01:34:18.93
Phi 4 0.00% 0.00% 0.01% 00:34:48.12
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 77: Twilight, twilight: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 78: Heatmaps according to location in Twilight. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.34 Beloved, beloved
Beloved
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.00% 0.01% 0.10% 00:15:06.06
Llama 1 13B 0.00% 0.03% 0.35% 00:16:46.39
Llama 1 65B 0.09% 0.37% 1.84% 00:58:18.24
Llama 2 70B 0.27% 0.87% 2.60% 00:57:14.62
Llama 3 70B 1.06% 2.36% 8.26% 01:09:13.78
Llama 3.1 70B 1.52% 3.23% 9.75% 01:10:59.00
DeepSeek v1 67B 0.02% 0.05% 0.49% 01:16:38.75
Qwen 2.5 72B 0.05% 0.26% 1.23% 01:18:30.16
Gemma 2 27B 0.01% 0.09% 0.73% 01:07:15.42
Phi 4 0.05% 0.09% 0.35% 00:28:41.59
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 79: Beloved, beloved: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 80: Heatmaps according to location in Beloved. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.35 The Memory Police, memory
The Memory Police
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.56% 0.63% 0.74% 00:13:17.65
Llama 1 13B 0.62% 0.66% 0.71% 00:13:25.28
Llama 1 65B 0.62% 0.65% 0.79% 00:49:08.03
Llama 2 70B 0.62% 0.66% 0.84% 00:47:45.59
Llama 3 70B 0.59% 0.63% 0.84% 00:59:09.73
Llama 3.1 70B 0.58% 0.63% 0.83% 00:57:40.92
DeepSeek v1 67B 0.63% 0.68% 0.92% 01:06:15.23
Qwen 2.5 72B 0.62% 0.67% 0.87% 01:01:38.03
Gemma 2 27B 0.62% 0.64% 0.70% 01:01:52.56
Phi 4 0.58% 0.59% 0.63% 00:25:00.18
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 81: The Memory Police, memory: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 82: Heatmaps according to location in The Memory Police. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.36 Nineteen Eighty-Four, 1984
Nineteen Eighty-Four
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.22% 0.50% 1.61% 00:21:21.22
Llama 1 13B 1.85% 4.83% 19.66% 00:21:23.13
Llama 1 65B 9.63% 23.91% 53.90% 01:11:45.83
Llama 2 70B 10.24% 25.93% 61.73% 01:12:58.15
Llama 3 70B 14.95% 33.98% 67.01% 01:29:40.53
Llama 3.1 70B 16.53% 36.28% 67.23% 01:31:41.77
DeepSeek v1 67B 1.23% 3.39% 18.57% 01:32:17.67
Qwen 2.5 72B 3.52% 10.16% 32.83% 01:41:22.32
Gemma 2 27B 1.78% 5.49% 17.62% 01:28:44.98
Phi 4 0.57% 1.39% 4.63% 00:35:20.85
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 83: Nineteen Eighty-Four, 1984: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 84: Heatmaps according to location in Nineteen Eighty-Four. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.37 The Subtle Knife, subtleknife
The Subtle Knife
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.00% 0.00% 0.03% 00:17:16.29
Llama 1 13B 0.00% 0.01% 0.10% 00:17:24.09
Llama 1 65B 0.00% 0.03% 0.30% 00:59:12.84
Llama 2 70B 0.01% 0.03% 0.29% 02:29:51.04
Llama 3 70B 0.08% 0.34% 6.03% 01:11:01.69
Llama 3.1 70B 0.09% 0.28% 4.14% 01:11:12.10
DeepSeek v1 67B 0.00% 0.01% 0.14% 01:21:23.03
Qwen 2.5 72B 0.00% 0.00% 0.16% 01:15:51.04
Gemma 2 27B 0.00% 0.00% 0.07% 01:13:01.78
Phi 4 0.00% 0.00% 0.03% 00:28:38.11
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 85: The Subtle Knife, subtleknife: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 86: Heatmaps according to location in The Subtle Knife. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.38 The Making of a Mediterranean Emirate, emirate
The Making of a Mediterranean Emirate
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.29% 0.32% 0.55% 00:18:22.98
Llama 1 13B 0.30% 0.32% 1.73% 00:18:48.93
Llama 1 65B 0.30% 0.33% 2.14% 01:03:15.78
Llama 2 70B 0.30% 0.34% 2.53% 01:05:15.72
Llama 3 70B 0.28% 0.34% 4.44% 01:17:34.58
Llama 3.1 70B 0.28% 0.34% 4.74% 01:18:38.46
DeepSeek v1 67B 0.32% 0.38% 2.37% 01:29:44.07
Qwen 2.5 72B 0.30% 0.36% 2.22% 01:38:18.63
Gemma 2 27B 0.30% 0.33% 1.12% 01:22:59.69
Phi 4 0.28% 0.28% 1.06% 00:32:12.37
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 87: The Making of a Mediterranean Emirate, emirate: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 88: Heatmaps according to location in The Making of a Mediterranean Emirate. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.39 Harry Potter and the Sorcerer’s Stone, hp1
Harry Potter and the Sorcerer’s Stone
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.04% 0.07% 0.39% 00:31:59.52
Llama 1 13B 0.14% 0.41% 2.86% 00:20:10.69
Llama 1 65B 2.95% 8.37% 46.63% 00:56:26.10
Llama 2 70B 3.18% 9.15% 47.38% 00:49:32.78
Llama 3 70B 14.17% 39.67% 85.75% 01:00:39.76
Llama 3.1 70B 17.59% 45.01% 87.31% 01:02:48.59
DeepSeek v1 67B 1.19% 2.85% 25.58% 01:10:07.92
Qwen 2.5 72B 8.49% 22.33% 67.12% 01:17:14.85
Gemma 2 27B 1.21% 4.86% 36.93% 01:05:32.95
Phi 4 0.28% 0.55% 1.59% 00:57:48.55
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 89: Harry Potter and the Sorcerer’s Stone, hp1: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 90: Heatmaps according to location in Harry Potter and the Sorcerer’s Stone. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.40 Harry Potter and the Goblet of Fire, hp4
Harry Potter and the Goblet of Fire
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.01% 0.03% 0.17% 00:32:53.33
Llama 1 13B 0.02% 0.08% 0.55% 00:35:20.49
Llama 1 65B 1.66% 5.09% 43.35% 02:03:02.22
Llama 2 70B 0.77% 2.64% 33.84% 02:01:51.02
Llama 3 70B 7.66% 25.30% 78.01% 02:27:34.90
Llama 3.1 70B 11.19% 32.16% 80.69% 02:33:44.93
DeepSeek v1 67B 0.16% 0.30% 3.37% 02:29:50.51
Qwen 2.5 72B 0.28% 0.85% 4.89% 02:39:47.84
Gemma 2 27B 0.07% 0.24% 2.36% 02:20:58.57
Phi 4 0.00% 0.01% 0.09% 01:02:10.06
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 91: Harry Potter and the Goblet of Fire, hp4: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 92: Heatmaps according to location in Harry Potter and the Goblet of Fire. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.41 The Catcher in the Rye, catcher
The Catcher in the Rye
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.17% 0.33% 0.49% 00:11:43.55
Llama 1 13B 0.32% 0.57% 1.80% 00:12:00.49
Llama 1 65B 1.09% 2.61% 15.39% 00:41:42.65
Llama 2 70B 1.68% 4.12% 23.13% 00:42:18.63
Llama 3 70B 2.70% 7.33% 43.66% 00:50:52.85
Llama 3.1 70B 3.65% 9.83% 46.10% 00:51:32.40
DeepSeek v1 67B 0.36% 0.56% 2.64% 00:55:38.22
Qwen 2.5 72B 0.49% 1.04% 4.42% 00:55:06.67
Gemma 2 27B 0.25% 0.55% 2.19% 00:46:51.43
Phi 4 0.18% 0.27% 0.67% 00:20:24.79
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 93: The Catcher in the Rye, catcher: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 94: Heatmaps according to location in The Catcher in the Rye. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.42 Lean In, leanin
Lean In
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.03% 0.04% 0.56% 00:14:13.58
Llama 1 13B 0.05% 0.17% 2.68% 00:15:00.22
Llama 1 65B 0.31% 0.90% 5.79% 00:49:39.22
Llama 2 70B 0.29% 0.95% 6.21% 00:50:18.34
Llama 3 70B 0.26% 1.17% 9.31% 01:02:16.57
Llama 3.1 70B 0.39% 1.51% 9.73% 01:02:39.76
DeepSeek v1 67B 0.11% 0.43% 4.16% 01:11:49.33
Qwen 2.5 72B 0.19% 0.47% 4.44% 01:19:04.88
Gemma 2 27B 0.05% 0.13% 1.98% 01:05:46.30
Phi 4 0.02% 0.09% 1.22% 00:26:31.67
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 95: Lean In, leanin: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 96: Heatmaps according to location in Lean In. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.43 The Bedwetter, bedwetter
The Bedwetter
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.14% 0.28% 0.39% 00:21:15.89
Llama 1 13B 0.03% 0.12% 0.32% 00:15:56.30
Llama 1 65B 0.09% 0.19% 0.38% 00:41:15.11
Llama 2 70B 0.10% 0.21% 0.55% 00:35:01.42
Llama 3 70B 0.15% 0.24% 0.52% 00:42:34.22
Llama 3.1 70B 0.15% 0.24% 0.54% 00:44:21.50
DeepSeek v1 67B 0.15% 0.34% 0.56% 00:52:18.98
Qwen 2.5 72B 0.12% 0.23% 0.47% 00:54:15.95
Gemma 2 27B 0.07% 0.12% 0.32% 00:50:06.27
Phi 4 0.06% 0.14% 0.27% 00:43:03.81
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 97: The Bedwetter, bedwetter: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 98: Heatmaps according to location in The Bedwetter. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.44 No Visible Bruises, novisible
No Visible Bruises
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.33% 0.36% 0.50% 00:49:33.08
Llama 1 13B 0.32% 0.33% 0.59% 00:33:52.32
Llama 1 65B 0.32% 0.34% 0.85% 01:30:24.99
Llama 2 70B 0.33% 0.36% 0.85% 01:17:15.88
Llama 3 70B 0.33% 0.34% 0.68% 01:33:06.62
Llama 3.1 70B 0.33% 0.35% 0.68% 01:35:40.82
DeepSeek v1 67B 0.35% 0.40% 0.71% 01:41:15.61
Qwen 2.5 72B 0.32% 0.36% 0.68% 01:55:37.64
Gemma 2 27B 0.33% 0.34% 0.54% 01:37:50.88
Phi 4 0.32% 0.34% 0.42% 01:36:16.77
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 99: No Visible Bruises, novisible: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 100: Heatmaps according to location in No Visible Bruises. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.45 Unglued, unglued
Unglued
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.31% 0.82% 2.35% 00:09:36.71
Llama 1 13B 0.38% 1.04% 2.92% 00:09:39.71
Llama 1 65B 0.60% 1.35% 3.35% 00:33:11.48
Llama 2 70B 0.45% 1.32% 3.38% 00:33:34.66
Llama 3 70B 0.64% 1.46% 3.38% 00:40:36.62
Llama 3.1 70B 0.69% 1.50% 3.34% 00:40:17.82
DeepSeek v1 67B 0.70% 1.46% 3.55% 00:50:52.78
Qwen 2.5 72B 0.65% 1.43% 3.18% 00:59:44.79
Gemma 2 27B 0.34% 0.99% 2.53% 00:49:42.29
Phi 4 0.08% 0.25% 1.59% 00:18:33.90
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 101: Unglued, unglued: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 102: Heatmaps according to location in Unglued. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.46 Embraced, embraced
Embraced
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.68% 1.24% 3.07% 00:25:42.20
Llama 1 13B 0.91% 1.63% 3.65% 00:19:19.73
Llama 1 65B 0.92% 1.85% 4.26% 00:52:07.03
Llama 2 70B 0.91% 1.82% 4.33% 00:45:19.79
Llama 3 70B 0.80% 1.49% 4.11% 00:54:43.37
Llama 3.1 70B 0.81% 1.55% 4.17% 00:57:18.49
DeepSeek v1 67B 0.93% 1.78% 4.41% 01:04:30.73
Qwen 2.5 72B 0.86% 1.55% 3.75% 01:11:41.35
Gemma 2 27B 0.83% 1.39% 3.42% 00:59:44.56
Phi 4 0.66% 0.94% 2.53% 00:49:09.94
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 103: Embraced, embraced: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 104: Heatmaps according to location in Embraced. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.47 The Hobbit, hobbit
The Hobbit
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.01% 0.01% 0.21% 00:34:20.87
Llama 1 13B 0.24% 0.60% 2.27% 00:25:50.33
Llama 1 65B 2.18% 5.38% 20.67% 01:05:58.13
Llama 2 70B 1.77% 4.29% 17.59% 00:55:37.34
Llama 3 70B 7.08% 17.75% 58.10% 01:06:14.12
Llama 3.1 70B 9.40% 21.81% 61.18% 01:07:54.81
DeepSeek v1 67B 0.33% 0.62% 2.81% 01:16:25.88
Qwen 2.5 72B 1.27% 2.62% 11.56% 01:20:55.39
Gemma 2 27B 0.28% 0.72% 2.75% 01:14:35.10
Phi 4 0.20% 0.39% 1.09% 01:08:05.60
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 105: The Hobbit, hobbit: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 106: Heatmaps according to location in The Hobbit. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.48 Brown Girl Dreaming, browngirldreaming
Brown Girl Dreaming
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.06% 0.08% 0.27% 00:10:39.32
Llama 1 13B 0.01% 0.07% 0.34% 00:08:01.68
Llama 1 65B 0.07% 0.12% 0.44% 00:22:08.44
Llama 2 70B 0.10% 0.21% 0.64% 00:19:40.27
Llama 3 70B 0.02% 0.14% 0.53% 00:24:06.87
Llama 3.1 70B 0.10% 0.14% 0.54% 00:24:12.94
DeepSeek v1 67B 0.08% 0.16% 0.41% 00:33:58.72
Qwen 2.5 72B 0.05% 0.10% 0.28% 00:39:30.65
Gemma 2 27B 0.07% 0.09% 0.27% 00:31:24.91
Phi 4 0.00% 0.00% 0.07% 00:23:42.67
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 107: Brown Girl Dreaming, browngirldreaming: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 108: Heatmaps according to location in Brown Girl Dreaming. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.49 Another Brooklyn, anotherbrooklyn
Another Brooklyn
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 1.39% 1.71% 1.98% 00:04:24.36
Llama 1 13B 1.09% 1.24% 2.05% 00:04:19.46
Llama 1 65B 1.13% 1.57% 2.71% 00:15:29.82
Llama 2 70B 1.14% 1.62% 2.66% 00:14:57.47
Llama 3 70B 1.25% 1.68% 3.12% 00:18:22.40
Llama 3.1 70B 1.34% 1.69% 3.08% 00:18:25.37
DeepSeek v1 67B 1.63% 1.90% 2.75% 00:29:30.56
Qwen 2.5 72B 1.36% 1.75% 2.65% 00:33:11.13
Gemma 2 27B 1.23% 1.66% 2.05% 00:26:43.76
Phi 4 1.03% 1.23% 1.69% 00:09:55.09
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 109: Another Brooklyn, anotherbrooklyn: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 110: Heatmaps according to location in Another Brooklyn. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.3.50 The Future of the Internet and How to Stop It, zittrain
The Future of the Internet and How to Stop It
Model Greedy $\tau_{\text{min}}\!=\!0.1\%$ Max. Runtime (hh:mm:ss)
Pythia 12B 0.04% 0.09% 0.23% 00:58:48.11
Llama 1 13B 0.05% 0.11% 1.66% 00:42:27.92
Llama 1 65B 0.09% 0.30% 3.04% 01:51:51.40
Llama 2 70B 0.09% 0.25% 3.49% 01:37:27.03
Llama 3 70B 0.10% 0.31% 2.64% 01:58:01.62
Llama 3.1 70B 0.14% 0.44% 2.86% 02:00:02.33
DeepSeek v1 67B 0.06% 0.15% 1.66% 02:02:38.44
Qwen 2.5 72B 0.07% 0.28% 1.96% 02:19:50.42
Gemma 2 27B 0.04% 0.11% 1.60% 02:05:30.57
Phi 4 0.02% 0.04% 0.50% 01:53:41.34
(a) Comparing greedy discoverable extraction rate, probabilistic extraction rate ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$), and (as a baseline) the maximum possible rate of generating training data across different models for the given decoding configuration. The maximum rate should be understood as a very loose upper bound on extraction (i.e., memorization); for sequences with low $p_{\bm{z}}$, it captures generalization.
(b) Distributions over suffix extraction probabilities ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$) for different models.
Figure 111: The Future of the Internet and How to Stop It, zittrain: Measuring probabilistic extraction (T=1T\!=\!1, k=40k\!=\!40). Based on our collective results, it is reasonable to conclude that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Phi 4 was not trained on in-copyright books, and serves as a validity check.
(a) Models known with certainty to have been trained on Books3.
(b) Models that we are confident were trained on (at least parts of) some books that are also contained in Books3.
(c) A model whose developers claim was not trained on whole copyrighted books.
Figure 112: Heatmaps according to location in The Future of the Internet and How to Stop It. At each character, we plot the maximum extraction probability among the overlapping $50$-token suffixes that span that character ($p_{\bm{z}}{\geq}\tau_{\text{min}}{=}0.1\%$). The LLMs in (a) were trained on Books3. Based on our collective results, it is reasonable to conclude that the LLMs in (b) were trained on some (sections of) copyrighted books—whether from Books3 or other sources. Training on (segments of) a given book does not mean that book is memorized by or extractable from all (or even any) models, with respect to our measurement procedure. (c) Phi 4 serves as a validity check; given its training data mix, we do not expect (with exceptions from web-scraped quotes) to extract large, substantive parts of copyrighted books.
D.4 Estimating how much of a book is memorized by a model
We estimate book-specific memorization for a given model $\theta$ and decoding scheme $\phi$ with a metric for extraction coverage. We define this metric and provide measurements for the different models and books we evaluate.
D.4.1 Quantifying extraction coverage
We compute extraction coverage as the fraction of characters in a book that lie within at least one suffix whose extraction probability $p_{\bm{z}}$ exceeds a threshold $\tau$.
Definition 2 (Extraction coverage).
Let $\mathsf{span}({\bm{z}})$ denote the set of character indices spanned by a $50$-token suffix of a given sequence $\bm{z}$}. Also let $\tau_{\text{min}}$ denote the minimum probability we consider to reflect extraction success. For a book $B$, LLM $\theta$, decoding scheme $\phi$, and probability threshold $\tau\in[\tau_{\text{min}},1]$,
$\mathsf{coverage}_{B,\theta,\phi}(\tau)=\frac{1}{L_{B}}\,\Bigl|\bigcup_{{\bm{z}}\in{\mathbb{Z}}(B):\,p_{{\bm{z}},\theta,\phi}\geq\tau}\,\mathsf{span}({\bm{z}})\Bigr|,$ (10)
where $L_{B}$ is the total number of characters in the book.
For details on $\tau_{\text{min}}$, see Equation 9 (Sections 2.1 & 5; Appendices C & E). In the main text, we omit $\theta$ and $\phi$ for brevity, since we always specify the exact model and decoding scheme being considered.
The results of our sliding-window procedure can be used to compute Equation 10. We pick a threshold $\tau$, then we filter the whole (sorted by start position) list of $p_{\bm{z}}$ we obtained with the sliding-window procedure to only contain sequences whose extraction probability is $\geq\!\tau$. Our sliding-window procedure keeps track of character-level metadata so that we can visualize heatmaps; we have information about the start/end character indexes for the prefixes and suffixes. Based on these intervals, we can compute which characters in the whole book are covered with $p_{\bm{z}}\!\geq\!\tau$, and then divide that by the total number of characters in the book.
High extraction coverage does not necessarily mean we can extract the whole book in one go, or that we have identified the minimal set of sequences to try to extract to get maximal coverage of the book. This is simply a procedure for estimating the proportion of total memorized text in a book, in order to do cross-book and cross-model comparisons, as in Section 4.2.
D.4.2 Extended extraction coverage results
We display extraction coverage in two ways, which support different types of comparisons.
  1. 1.
    Model-level histograms. For each model, we plot the distribution of extraction coverage across the $50$ evaluated books. This provides a high-level sense of how coverage varies across books, and enables comparisons of this distribution across models. For example, in the main paper we show two histograms for Llama 1 65B and Llama 3.1 70B at $\tau\!=\!10\%$. These results indicate that most books have little extraction coverage for both models at $\tau=10\%$, but with clear exceptions. They also reveal that some models exhibit substantially higher coverage than others—in this case, Llama 3.1 70B shows higher maximum coverage and more books with significant coverage, i.e., it has memorized much more book text than Llama 1 65B.
  2. 2.
    Book-level percentages. For each book, we compare extraction coverage across different models. This not only indicates how memorized a given book is by a particular model, but also highlights variation in memorization of a given book across models. We provide examples of this comparison in the table in Figure 4 (right), for Harry Potter and the Sorcerer’s Stone and Sandman Slim, evaluated with Llama 3.1 70B, Llama 1 65B, and Pythia 12B.
Here, we provide more extensive extraction coverage results. We include both types of comparisons for the $10$ models for which detailed sliding-window results are given in Appendix D.3. All results are shown for $\tau=\tau_{\text{min}}=0.1\%$, which corresponds to the total (largest) amount of extraction coverage revealed by our experiments. Figure 113 plots histograms of extraction coverage across all $50$ books for each model, and Table LABEL:app:tab:coverage reports per-book extraction coverage results for each model.
Table 4: Extraction coverage for all $50$ books and the $10$ models for which we include full results in Appendix D.3. We use Definition 2 to compute extraction coverage for our results from sliding-window experiments (Appendix D.3). These experiments use $100$-token sequences ($50$-token prefix ++ $50$-token suffix) and top-$k$ decoding (T=1T{=}1, k=40k{=}40). We set $\tau=\tau_{\text{min}}=0.1\%$ to compute the total (highest) amount of extraction coverage that we observe. We highlight instances where coverage with $\tau_{\text{min}}$ is $2.5\%$ in blue, $5\%$ in purple, $10\%$ in red.
Proportion of each book that is extractable with $p_{\bm{z}}\geq 0.1\%$ for different models
Book Pythia 12B Llama 1 13B Llama 1 65B Llama 2 70B Llama 3 70B Llama 3.1 70B DS v1 67B Qwen 2.5 72B Gemma 2 27B Phi 4
A Game of Thrones 0.67% 0.76% 3.34% 3.07% 37.39% 39.39% 0.81% 1.25% 0.87% 0.59%
A Visit from the Goon Squad 0.20% 0.07% 0.29% 0.36% 0.74% 1.06% 0.19% 0.22% 0.17% 0.11%
After I’m Gone 0.36% 0.16% 0.39% 0.31% 0.40% 0.46% 0.41% 0.36% 0.28% 0.24%
Alice’s Adventures in Wonderland 1.65% 18.36% 42.89% 42.72% 47.74% 48.80% 38.86% 38.00% 24.94% 8.33%
All the Onions 2.61% 0.27% 0.88% 1.19% 2.77% 2.77% 2.77% 2.55% 2.11% 0.00%
Another Brooklyn 2.59% 1.97% 3.01% 2.68% 2.84% 3.16% 2.85% 2.62% 2.46% 1.87%
Ararat 0.46% 0.37% 0.43% 0.49% 0.43% 0.45% 0.56% 0.56% 0.42% 0.18%
Beloved 0.09% 0.35% 1.87% 3.58% 9.78% 12.54% 0.34% 1.55% 0.69% 0.32%
Blink 0.48% 0.60% 2.19% 2.87% 6.86% 9.73% 1.51% 2.62% 0.87% 0.65%
Brown Girl Dreaming 0.43% 0.30% 0.81% 0.83% 0.54% 0.55% 0.80% 0.53% 0.58% 0.00%
Catch-22 0.32% 0.60% 2.60% 2.98% 10.01% 15.64% 0.60% 0.93% 0.56% 0.23%
Dante and the Origins of Italian Culture 0.04% 0.00% 0.00% 0.00% 0.36% 0.55% 0.06% 0.08% 0.06% 0.08%
Dead Ringer 0.66% 0.47% 0.68% 0.56% 0.64% 0.65% 0.68% 0.69% 0.50% 0.28%
Down and Out in the Magic Kingdom 0.47% 0.31% 0.46% 0.56% 0.94% 1.11% 0.47% 0.48% 0.37% 0.00%
Drown 0.00% 0.00% 0.00% 0.21% 0.37% 0.37% 0.12% 0.00% 0.00% 0.00%
Dungeons and Dragons and Philosophy 0.86% 0.45% 1.43% 1.31% 1.61% 1.62% 1.29% 0.91% 0.48% 0.31%
Embraced 4.12% 6.38% 6.92% 6.97% 5.68% 6.04% 6.70% 5.48% 4.76% 2.94%
Harry Potter and the Goblet of Fire 0.13% 0.32% 26.15% 14.08% 82.35% 88.66% 1.09% 3.64% 1.12% 0.16%
Harry Potter and the Sorcerer’s Stone 0.41% 1.97% 34.83% 36.41% 93.74% 96.27% 12.00% 58.21% 21.46% 1.73%
Lean In 0.38% 1.05% 4.61% 5.16% 8.76% 10.73% 2.49% 3.03% 1.49% 0.94%
M. Butterfly 1.60% 1.09% 1.46% 1.81% 2.46% 2.91% 1.72% 1.47% 1.20% 0.66%
Nineteen Eighty-Four 1.88% 16.77% 67.27% 71.21% 79.97% 81.19% 12.50% 30.59% 19.16% 5.09%
No Visible Bruises 0.65% 0.44% 0.40% 0.56% 0.56% 0.56% 0.66% 0.59% 0.49% 0.49%
Sandman Slim 0.36% 0.23% 0.32% 0.32% 0.38% 0.43% 0.37% 0.38% 0.34% 0.25%
Sunburn 0.34% 0.27% 0.29% 0.45% 0.44% 0.43% 0.38% 0.33% 0.33% 0.29%
The Beautiful Struggle 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
The Bedwetter 0.57% 0.50% 0.60% 0.61% 0.62% 0.63% 0.69% 0.48% 0.58% 0.39%
The Brief Wondrous Life of Oscar Wao 0.42% 0.42% 0.70% 0.88% 1.52% 1.76% 0.51% 0.52% 0.34% 0.00%
The Catcher in the Rye 0.82% 2.29% 11.57% 18.14% 31.74% 41.30% 2.23% 4.35% 2.73% 1.02%
The Confessions of Max Tivoli 0.11% 0.00% 0.00% 0.00% 0.10% 0.11% 0.13% 0.00% 0.07% 0.00%
The Curious Incident of the Dog in the Night-Time 0.33% 0.46% 1.82% 2.36% 7.47% 9.91% 0.57% 0.87% 0.39% 0.26%
The Da Vinci Code 0.27% 0.40% 1.47% 2.06% 6.27% 10.14% 0.63% 0.84% 0.27% 0.27%
The Fugitive 0.46% 0.36% 0.38% 0.38% 0.42% 0.42% 0.49% 0.52% 0.50% 0.18%
The Future of the Internet and How to Stop It 0.23% 0.46% 1.13% 1.27% 1.88% 2.18% 0.40% 0.93% 0.41% 0.22%
The Great Gatsby 1.45% 14.28% 73.60% 71.82% 85.29% 88.77% 23.69% 26.86% 33.64% 8.71%
The Handmaid’s Tale 0.31% 0.30% 2.82% 3.78% 6.59% 9.57% 0.34% 0.92% 0.28% 0.29%
The Hobbit 0.07% 2.37% 19.78% 15.03% 60.82% 68.69% 2.25% 10.59% 2.55% 1.38%
The Making of a Mediterranean Emirate 0.48% 0.41% 0.58% 0.68% 0.64% 0.63% 0.67% 0.59% 0.45% 0.30%
The Memory Police 0.88% 0.83% 0.83% 0.83% 0.77% 0.77% 0.84% 0.80% 0.76% 0.73%
The Myth of Sisyphus 0.82% 2.12% 5.94% 9.19% 20.22% 24.90% 2.20% 5.90% 3.15% 2.32%
The President’s Vampire 0.59% 0.48% 0.56% 0.56% 0.56% 0.58% 0.60% 0.58% 0.53% 0.11%
The Subtle Knife 0.00% 0.15% 0.38% 0.18% 2.21% 1.75% 0.08% 0.00% 0.07% 0.00%
The Water Dancer 0.71% 0.63% 0.68% 0.74% 0.70% 0.91% 0.76% 0.72% 0.69% 0.63%
The World’s Wife 2.16% 1.39% 3.20% 3.13% 7.41% 7.56% 2.08% 2.78% 2.24% 0.69%
This Is How You Lose Her 0.70% 0.53% 0.86% 1.02% 1.11% 1.15% 0.69% 0.69% 0.41% 0.00%
Twilight 0.00% 0.00% 2.25% 1.65% 4.82% 5.67% 0.05% 0.05% 0.16% 0.00%
Ulysses 0.10% 0.27% 1.55% 10.14% 65.68% 76.11% 0.44% 0.88% 0.66% 0.36%
Unglued 3.07% 4.32% 4.84% 4.58% 4.96% 5.20% 4.85% 4.55% 3.49% 1.74%
We Were Eight Years in Power 0.89% 1.59% 8.40% 7.92% 12.06% 14.26% 1.42% 4.40% 1.43% 0.74%
Who Is Rich 1.04% 0.86% 0.93% 0.99% 0.98% 0.97% 1.04% 1.00% 0.95% 0.84%
X
X
Figure 113: Extraction coverage distributions for $10$ models on the $50$ books we evaluate. We use Definition 2 to compute extraction coverage for our results from sliding-window experiments (Appendix D.3). These experiments use $100$-token sequences ($50$-token prefix ++ $50$-token suffix) and top-$k$ decoding (T=1T{=}1, k=40k{=}40). We set $\tau=\tau_{\text{min}}=0.1\%$ to compute the total (highest) amount of extraction coverage that we observe. As in the rest of this paper, we divide these results into our three model categories: (top) models that we know with certainty were trained on Books3, (middle) models that we can confidently conclude were trained on (at least parts of) some books that were also contained in Books3, and (bottom) a model whose developers claim was not trained on copyrighted books.
D.5 Selection of extracted sequences
This project generated over a million extracted sequences. In an earlier draft of this paper, we included a brief selection of arbitrarily chosen extracted sequences from some of the $50$ books in our sliding-window probabilistic extraction experiments. (See Section 6 and Appendix D). Given the current length of the paper, we omit these sequences for brevity. We provide some one-off examples of extraction in Section 5, Figure 13, and Appendix E. We also provide samples of the diff from our reconstruction experiments, which capture (near-)verbatim extraction in Section 7 and Appendix F.
Appendix E Measurement validity
In this appendix, we provide more detail on the validity experiments introduced in Section 5. These experiments support that our measurements capture true instances of extraction. On this basis, we expand our analysis to models with undisclosed information about training on books—DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B. Because our controls show that our method does not mistake chance generation for extraction, we can be confident that when these models register extraction success, it reflects memorization.
Given the length of this appendix, we provide a summary:
  • Extraction of known training data is evidence of memorization. For models where the training data are known, high-probability generation of a verbatim suffix given its prefix is widely accepted as evidence of memorization [carlini2021extracting, carlini2023extracting, carlini2023quantifying, hayes2025measuringmemorizationlanguagemodels]. Probabilistic extraction provides a more fine-grained version of the same test: rather than returning a single greedy continuation, it assigns a probability to the target suffix and reveals whether it surpasses a minimum acceptable threshold.
  • We adopt a deliberately conservative minimum threshold. Probabilistic extraction produces a continuous measure of extraction, but requires a cutoff. We set $\tau_{\text{min}}$ very high to ensure we never over-count memorization. This makes our results lower bounds: we under-count memorization (Section 2). We capture more instances of memorization than greedy decoding, but we still do not over-count. We defer more fine-grained settings of $\tau_{\text{min}}$ to future work.
  • Validating against non-training data The majority of our experiments—those on Llama and Pythia models—fit the case of extracting known training data from models. It is known that Books3 was included in the training data, and we are running experiments to extract Books3 sequences. But we can also do more to ensure validity: we can test our procedure on known non-training data, and show that suffix probabilities fall below our chosen $\tau_{\text{min}}$. We (and hayes2025measuringmemorizationlanguagemodels) do this for an additional sanity check. In particular, we perform over $200$ negative control experiments (Section 5.2 & Appendix E.4). Work on greedy-decoded discoverable extraction often does not do this; it assumes validity [carlini2023quantifying] or uses shorter prefix lengths to induce less bias on the output [nasr2023scalable] (Appendix E.6). Overall, our results confirm that our choice of $\tau_{\text{min}}$ is very conservative; we could set it lower without risking false positives. However, we are deliberately cautious, and defer this to future work (Section 5.3).
  • Stress-testing extraction with longer prefixes. When validating $\tau_{\text{min}}$ on known training data, we can also ask whether longer prefixes make memorization more discoverable. Because suffix probabilities are conditional on the prefix, extending the prompt can bias the model’s output toward a certain part of its learned distribution (which can also further be affected by the decoding algorithm). As a result, as others have shown [carlini2023quantifying], for experiments on known training data, when a piece of training data is memorized, a longer prefix prompt can raise the probability of the target suffix and expose it as memorized (via greedy decoding). This memorization may not be discoverable with only a $50$-token prompt, but may become discoverable with a longer prompt.
    Our own experiments confirm these patterns, but also show clarify, for a specific work, that longer prompts do not always lead to successful extraction: if a sequence is not memorized, we are unable to find a long enough prefix to make it it extractable (Appendix E.3). This supports the conclusion that probabilistic extraction captures true memorization, and clarifies the limits of analogies that treat models as “monkeys at the typewriter” (Appendix E.5).
  • Bringing the pieces together. Our validity experiments combine baselines, negative controls, and varied-prefix tests. For models known to be trained on Books3, we run baselines with different prefix lengths as points of comparison (Appendix E.3). We then add two sets of negative controls: (i) Phi 4, a model not trained on whole copyrighted books (Appendices D & E.4.1), and (ii) books published after the training cutoff for models trained on Books3 (Appendix E.4.2). Across more than $200$ such experiments, suffix probabilities for non-training data are essentially zero, and the rare nonzero values are orders of magnitude below $\tau_{\text{min}}$. As a result, our procedure does not mistakenly register non-training data as extractable, reinforcing that what we count as extraction reflects true instances of memorization.
  • Applying the procedure to models with unknown training data. Based on these results, we can also apply the procedure to models whose training data are undisclosed. For sequences that register as extractable in these models, we can confidently interpret them as evidence of memorization, regardless of whether the training-data source was Books3 directly or overlapping material from elsewhere (Appendix E.6). Importantly, as we state throughout, for our experiments on these models, our claims about membership in the training data are only about specific extracted sequences.
Appendix outline.
  • Appendix E.1. Motivation for these experiments.
  • Appendix E.2. Setup.
  • Appendix E.3. Additional baseline experiments that for models known to be trained on Books3.
  • Appendix E.4. Additional negative control experiments on non-training data.
  • Appendix E.5. Additional discussion on “the monkey at the typewriter.”
  • Appendix E.6. Discussion of how our work fits in with prior literature on membership inference.
E.1 Motivating our validity experiments
The main point of these experiments is to validate that our measurements capture true instances of extraction, rather than chance generation of verbatim suffixes. They help us make confident claims about both our results for models trained on Books3 and our application of the procedure to models with undisclosed training data. Here, we expand on these motivations (Appendix E.1.1), connect our work to the literature on greedy-decoded discoverable extraction (Appendix E.1.2), and clarify the role of varying prefix length in our baselines and negative controls (Appendix E.1.3).
E.1.1 Setting $\tau_{\text{min}}$ and extending to models with unknown training data
We can be fairly confident that DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B were trained on at least some books: the Pile [gao2020pile] and several other common LLM pre-training datasets either directly include Books3 or other book corpora [lee2023explainers, lee2023talkin]. Indeed, other work on extraction with uncertain training data treats the Pile as part of a proxy for the true training set [nasr2023scalable, nasr2025scalable]. However, without disclosure, we cannot be absolutely certain. And without knowing whether Books3 was in fact included, we cannot—without additional validity experiments—make claims about memorization that are as strong as those we make for models known to be trained on Books3. This is because membership in the training data is a prerequisite for memorization; by definition, a model cannot memorize sequences it has never seen.
This is also why observing $p_{\bm{z}}$ on non-training data is important for setting $\tau_{\text{min}}$. By comparing suffix probabilities from negative controls on non-training data (Section 5.2, Appendix E.4) with those from baselines on known training data (Section 5.1, Appendix E.3), we can draw a conservative line for what $p_{\bm{z}}$ is high enough to be indicative of memorization.
That is, our negative controls clarify how to set the threshold that we use to count extraction success, $\tau_{\text{min}}$—the minimum $p_{\bm{z}}$ at which we interpret our measurements as evidence of memorization. Running our procedure on non-training data, we confirm that the values of $p_{\bm{z}}$ we observe are far below our chosen cutoff. That way, when we count extraction success for a given sequence with $p_{\bm{z}}\geq\tau_{\text{min}}$, we can be confident that it does not reflect chance generation. Our results in Section 5 and the rest of this Appendix confirm this: $\tau_{\text{min}}=0.1\%$ is highly conservative——well above any $p_{\bm{z}}$ observed in negative control experiments.
We therefore use these results to justify our choice of $\tau_{\text{min}}$ in this paper. More extensive experiments would be needed to lower $\tau_{\text{min}}$ or calibrate it model by model. Here, we prefer to be conservative: we would rather under-count memorization than risk over-counting, and defer fine-grained threshold settings to future work.
E.1.2 Relating greedy-decoded and probabilistic discoverable extraction
To our knowledge, we run the largest set of validity experiments in the ML literature on memorization and extraction. Such experiments are not typically performed in work on greedy-decoded discoverable extraction [nasr2023scalable, nasr2025scalable, carlini2021extracting]. hayes2025measuringmemorizationlanguagemodels conducted one small-scale test on non-training data for a small Pythia model, and used this to support validity claims for probabilistic extraction from much larger and higher quality Llama 1 models.
Given these norms, we initially omitted our negative controls from the main paper, deferring discussion of Phi 4 experiments to the Appendix and omitting our varied-prefix experiments entirely from the paper. However, after receiving some feedback about earlier drafts of this work that indicated serious misunderstandings about the nature of memorization and extraction, we realized that including the results of these experiments in the main text could help some readers better understand our results and claims. We now clearly delineate two sets of negative controls, both in this appendix and in the main paper text (Section 5.2): the experiments on Phi 4 (Appendix E.4.1) and the experiments on books published after training-date cutoff (Appendix E.4.2).
Here, we clarify the relationship between prior definitions and our approach, and explain how this motivates our validity experiments.
Greedy-decoded discoverable extraction.
It is widely accepted that successful greedy-decoded discoverable extraction on known training data is evidence of memorization; memorization remains the most parsimonious explanation [hayes2025measuringmemorizationlanguagemodels, carlini2021extracting, nasr2023scalable, carlini2025blog, cooper2024files]. In this procedure, greedy decoding always selects the top-ranked token at each step, making it a deterministic approximation of the most likely sequence. carlini2021extracting describe this choice as a tractable proxy for the $\operatorname*{arg\,max}$ in their definition of model knowledge extraction:
Definition 3 (Model Knowledge Extraction, from carlini2021extracting).
A string $s$ is extractable from an LM $f_{\theta}$ if there exists a prefix $c$ such that:
$s\leftarrow\operatorname*{arg\,max}_{s^{\prime}:|s^{\prime}|=N}f_{\theta}(s^{\prime}|c)$
We abuse notation slightly here to denote by $f_{\theta}(s^{\prime}|c)$ the likelihood of an entire sequence $s^{\prime}$. Since computing the most likely sequence $s$ is intractable for large $N$, the $\operatorname*{arg\,max}$ in [the definition] can be replaced by an appropriate sampling strategy (e.g., greedy sampling) that reflects the way in which the model $f_{\theta}$ generates text in practical applications.
We add emphasis to two areas in this definition, which we discuss below. First, we note that this definition does not mention training data; it is just a definition about a model having knowledge of string $s$. carlini2021extracting connect extraction of training data to memorization through a second definition:
Definition 4 ($k$-Eidetic Memorization, from carlini2021extracting).
A string $s$ is $k$-eidetic memorized (for $k\geq 1$) by an LM $f_{\theta}$ if $s$ is extractable from $f_{\theta}$ and $s$ appears in at most $k$ examples in the training data $X:|\{x\in X:s\subseteq x\}|\leq k$.
Successful extraction, under these definitions, is evidence for memorization. Ideally, $s$ should be computed as the most likely sequence under $f_{\theta}$. However, since this is not tractable to compute in practice, greedy decoding is suggested as a reasonable proxy. It forces top-1 choices regardless of their actual probabilities under the model.
The relationship to probabilistic extraction.
Greedy decoding is equivalent to top-$k$ decoding with T=1T{=}1 and k=1k={1}; one uses the base model distribution (T=1T{=}1) and then only considers the token with top-$1$-ranked logit (k=1k{=}1). Probabilistic extraction relaxes this by considering a wider set of high-probability continuations (e.g., top-$40$). This better reflects the model’s distribution and avoids the problem where greedy decoding yields globally low-probability sequences. hayes2025measuringmemorizationlanguagemodels motivate their metric precisely from this point: the greedy-decoded suffix is locally optimal, but it may lead to a globally relatively low probability sequence under the model.
A hypothetical example illustrates this gap: a greedy-decoded suffix can have vanishing total probability, while a slightly different path within the top-$k$ yields a much higher-probability suffix. For this example, probabilities are computed with respect to the base $T=1$ distribution. Consider a greedy-decoded suffix where the first token has $99.9\%$ probability and is selected, and then every top-$1$ token after it for $49$ tokens has probability $1\%$; the total probability for the $50$-token suffix would be $999\times(0.01)^{49}{\approx}9.99\times 10^{-99}.$ Now imagine we instead were not restricted to always picking the top-$1$ probability token; we relax this requirement to, for example, allow to select from the top-$k$ tokens. Again we consider probabilities computed with respect to the base $T=1$ distribution, so that sequence probabilities are computed off the same logits. For the first token position, the top-$2$ token might have probability $1\%$, and then the remaining top-$1$ $49$ tokens could each have probability $99.9\%$; the total probability for this suffix would be $01\times(0.999)^{49}{\approx}0.001$ (or about $0.1\%$).
The point is, by allowing for a bit more flexibility in the decoding strategy, the model can potentially produce much higher probability sequences; when we observe that such (high-enough-probability, $\geq\tau_{\text{min}}$) sequences match the verbatim suffix from the training data, then we count them as extracted. As long as we choose $\tau_{\text{min}}$ appropriately—i.e., such that it does not capture low-probability sequences that are perhaps reflective of generalization—we should obtain higher fidelity measurements of memorization. We should have fewer false negatives than greedy-decoded discoverable extraction. And notably, as we show in Appendix A.3, this is actually cheaper to compute than how work like carlini2021extracting computes extraction, as we can compute probabilistic extraction directly off of the logits, rather than needing to actually generate completions.
Greedy-decoding is a degenerate case of probabilistic extraction.
Greedy decoding can be seen as a degenerate case of probabilistic extraction, corresponding to $\tau_{\text{min}}=100\%$. Each step assigns all probability mass to a single token under a normalized top-$1$ distribution, even if that token’s true model probability is very low. Probabilistic extraction surfaces this mismatch more clearly and is a better proxy for identifying high-probability memorized sequences. Because it includes the greedy suffix when it matches the training suffix, it never misses those cases, while also capturing others that greedy decoding fails to expose.
E.1.3 The role of prefix length in our experiments
We run our validity experiments with various prefix lengths, not just for $100$-token sequences with $50$-token prefixes and $50$-token suffixes. This is a deliberate choice because additional context biases the model’s output distribution toward particular continuations. For sequences from the training data, this is known to surface memorization that may not appear for shorter prefixes. carlini2023quantifying call this the discoverability phenomenon: “some memorization only becomes apparent …when the model is prompted with a sufficiently long context.” Through the lens of probabilistic extraction, a longer training-data prefix can raise the probability of the target suffix in Equation 1, such that it surpasses $\tau_{\text{min}}$. (It is worth noting that any conditional extraction procedure—i.e., one that uses prompts, rather than unconditionally generating from the LLM—statistically biases the output. This is one of the reasons that extraction under-counts memorization; it only searches a part of the learned distribution.)
By testing various prefix lengths on known training data, we can estimate an upper bound on memorization (with respect to our extraction procedure) for a given work. We expect longer prefixes to expose more memorization, but that increasing prefix length will not typically induce all suffix probabilities to $100\%$; that is, we will observe more extraction only if there is more memorization to uncover. It is not the case that one can simply make the prefix arbitrarily long and then extract any target suffix. We discuss this further in Appendices E.3 and E.5.
This logic also strengthens our negative controls. In this case, we should not be able to measure extraction for any unique sequences of non-training data—even when we intentionally induce more bias with longer prompts. Again, this shows that arbitrarily long suffixes do not all of a sudden mean we will register extraction with our measurements. Here, we should not be able to do so because the sequences were not included in the training data.
E.2 Setup
Books3 books.
Running experiments with longer sequences is significantly more expensive. We therefore limit our results on varied prefixes to a small selection of books from Books3: Beloved [beloved], Harry Potter and the Sorcerer’s Stone [hp1], This Is How You Lose Her [loseher], and We Were Eight Years in Power [wewere]. We picked these $4$ books because each represents (roughly speaking) a different pattern for memorization in our sliding-window results on $50$-token prefixes (Appendix D) for the model that exhibits the most memorization: Llama 3.1 70B. This Is How You Lose Her is barely memorized at all (Appendix D.3.12. Beloved has fragmented memorization “hot-spots;” there are many throughout the book, but they are rather disconnected, separated by lots of regions that do not register any memorization (Appendix D.3.34). We Were Eight Years in Power has several “hot-spots;” some are fragmented, but some indicate long, continuous regions of memorized text (Appendix D.3.7. Harry Potter and the Sorcerer’s Stone is effectively memorized entirely, though not all suffixes have equally high extraction probability (Appendix D.3.39).
Non-training data.
The large majority of sequences in the in-copyright Books3 books above should not be training data for Phi 4; exceptions are highly duplicated text, like boilerplate copyright notices and occasional popular quotes that are included from other sources (e.g., the Bible, newspaper articles, earlier books). We also purchased a set of $4$ books published in 2025—after the release date of the newest model we test, Qwen 2.5 72B (September 19, 2024). (See Table 5.) We run these books through models known to have been trained on Books3 as another point of comparison. Similar to the above note about Phi 4, none of the unique text in these books should ping as extractable. Exceptions are, once again, boilerplate text and popular or otherwise duplicated quotes from earlier sources.
Author Title Publication Date Status
1 Gareth Brown The Society of Unknowable Objects [dante] July 31, 2025 ©
2 Emily Henry Great Big Beautiful Life [greatbig] April 22, 2025 ©
3 Ocean Vuong The Emperor of Gladness [emperorgladness] May 13, 2025 ©
4 Sarah Wynn-Williams Careless People [careless] March 11, 2025 ©
Table 5: The 44 books that post-date training cutoff for all models we test. We use these books for one set of our negative control experiments. See Section 5.2 and Appendix E.4.2.
Overall, we want to see how changing prefix length impacts how much more memorization we can discover. As we increase prefix length, we expect to discover more memorization of Books3 text for models trained on Books3, and to effectively observe no major changes for Phi 4 for unique text from these books nor for unique text from 2025 books.
Models.
Following from above, we run these experiments on $3$ models known to have been trained on Books3. We include Llama 3.1 70B, because this is the model that has memorized the most of all Llama models we test. We also include Llama 3.1 8B, as it is a much smaller model from the same family. It memorizes less than Llama 3.1 70B. (These results omitted for brevity in Appendix D.) While it may be possible that increasing prefix length for Llama 3.1 70B surfaces many more books that are effectively entirely memorized (i.e., like Harry Potter and the Sorcerer’s Stone), we do not expect this to be the case for Llama 3.1 8B, since it is a much smaller capacity model. Together, this will clarify how it is not the case that just making prefixes longer means that one could generate any target suffix they so desire—regardless of whether it was memorized.
We also include Llama 2 13B. This model is a similar size as Phi 4, which we use in our negative controls. They have similar capacity, but were clearly trained very differently. While we do not know all of the differences, we do know that Llama 2 13B was trained on Books3, while Phi 4 was not. For one experiment on non-training data, we also include Qwen 2.5 72B as a point of comparison.
Procedure.
We use the same sliding-window procedure as in Section 4.1, documented in detail in Appendix D.2. We inherit all of those setup details; the only difference is that we run for sequences of different lengths (and associated different prefix lengths) to extract $50$-token suffixes: $75$ tokens ($25$-token prefixes ++ $50$-token suffixes); $100$ tokens ($50$-token prefixes ++ $50$-token suffixes, i.e., the typical setting); $150$ tokens ($100$-token prefixes ++ $50$-token suffixes); $250$ tokens ($200$-token prefixes ++ $50$-token suffixes); $450$ tokens ($400$-token prefixes ++ $50$-token suffixes); and, $850$ tokens ($800$-token prefixes ++ $50$-token suffixes). On our hardware, it takes approximately $13$ hours to run We Were Eight Years in Power through Llama 3.1 70B with $850$-token sequences (compared to $\sim\!1.5$ hours with our typical setup with $100$ tokens, see Appendix D.3.7). This cost is the main reason we do not expand these experiments to a larger set of books and models.
E.3 Results of varying prefix length on known training data
We next provide extended results for baseline experiments that vary prefix lengths on known training data. Our results yield the outcomes we expect: we are able to surface more memorization with increased prefix length—to a point. Except in rare cases like Harry Potter and the Sorcerer’s Stone and Llama 3.1 70B, in which the entire book is memorized, arbitrarily increasing prefix length does not make very training data suffix extractable. This makes sense, as not all training data are memorized; our measurements should therefore not register all training sequences as extractable. These results also make clear that absence of extraction is not itself evidence that a sequence is not a member of the training data. Most sequences from Books3 are not extractable even from models like Llama 3.1 70B, which memorize a lot. So the claim we make with respect to absence of extraction signal is in this one direction. When we make claims about training data membership, these are either based on ground-truth knowledge (e.g., knowing the Books3 was in the training data and that 2025 books were not), or conclusions that we draw based on comparisons to our negative controls (Appendix E.6).
We include results for $3$ different models on $4$ books from Books3, as documented in Appendix E.2. The largest model we test is Llama 3.1 70B (Figure 115). For a point of comparison with a much smaller capacity model, we also include the smaller model from the same family—Llama 3.1 8B (Figure 116) and Llama 2 13B (Figure 117), since this model is of a similar size as Phi 4 (Section 5.2, Appendix E.4.1).
Figure 114: Plotting extraction coverage (for $\tau{=}\tau_{\text{min}}{=}0.1\%$) for the $5$ books we test with varying prefix lengths and Llama 3.1 70B. 44 of these books are in Books3 and thus in Llama’s training data. The remaining book, Great Big Beautiful Life, is used in our negative controls on non-training data; extraction coverage is $0\%$ regardless of prefix length. As we increase prefix length, extraction coverage for each book in the training data begins to level off. While Harry Potter and the Sorcerer’s Stone is entirely memorized by the model, it does not appear to be the case that Beloved, We Were Eight Years in Power, or This Is How You Lose Her would be entirely extractable with respect to $50$-token suffixes and even longer prefixes.
Figure 115: Llama 3.1 70B Baseline: Sliding-window procedure for various prefix lengths. We show results for $4$ books from Books3: Beloved [beloved], Harry Potter and the Sorcerer’s Stone [hp1], This Is How You Lose Her [loseher], and We Were Eight Years in Power [wewere]. Llama models are known to have been trained on Books3. We run probabilistic extraction on Llama 3.1 70B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. Extraction signal generally increases as prefix length increases. However, this only appears to be the case if there is indeed more memorization to surface; increasing prefix length does not always make previously unextractable sequences extractable. There are also diminishing increases as the prefix lengthens.
Figure 116: Llama 3.1 8B Baseline: Sliding-window procedure for various prefix lengths. We show results for $4$ books from Books3: Beloved [beloved], Harry Potter and the Sorcerer’s Stone [hp1], This Is How You Lose Her [loseher], and We Were Eight Years in Power [wewere]. Llama models are known to have been trained on Books3. We run probabilistic extraction on Llama 3.1 8B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. Extraction signal generally increases as prefix length increases. However, this only appears to be the case if there is indeed more memorization to surface; increasing prefix length does not always make previously unextractable sequences extractable. There are also diminishing increases as the prefix lengthens.
These results illustrate the high-level points described above. While for Llama 3.1 70B we are able to extract a lot more memorized sequences, we are able to extract relatively less from Llama 3.1 8B. That model has memorized less overall, which makes sense because it is a significantly smaller model. However, even from Llama 3.1 70B, for some books there is very little overall memorization that we can expose, regardless of prefix length. We can see this with the results in Figure 115 for This Is How You Lose Her. For Llama 3.1 70B, we observe diminishing returns for increasing prompt length. This suggests that there is a maximum number of sequences we can extract from each work, even in the limit of setting the prefix length especially long. We can show this with plots of extraction coverage (Section 4.2, Appendix D.4) as a function of prefix length, as in Figure 114.
Figure 117: Llama 2 13B Baseline: Sliding-window procedure for various prefix lengths. We show results for $4$ books from Books3: Beloved [beloved], Harry Potter and the Sorcerer’s Stone [hp1], This Is How You Lose Her [loseher], and We Were Eight Years in Power [wewere]. Llama models are known to have been trained on Books3. We run probabilistic extraction on Llama 2 13B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. Extraction signal generally increases as prefix length increases. However, this only appears to be the case if there is indeed more memorization to surface; increasing prefix length does not always make previously unextractable sequences extractable. There are also diminishing increases as the prefix lengthens.
E.4 Results of negative control experiments
We provide additional results for our varied-prefix-length negative control experiments (Section 5.2) on both Phi 4 (Appendix E.4.1) and non-training data that post-dates training cutoffs (Appendix E.4.2). While these experiments provide strong evidence for the validity of our procedure, we would want to run much more extensive experiments to responsibly lower $\tau_{\text{min}}$ or to make it model-specific. We defer this to future work, and here opt to pick a very conservative global $\tau_{\text{min}}{=}0.1\%$, which we discuss also in Section 5. The experiments that we run here are of a more substantial scale than the original probabilistic discoverable extraction paper by hayes2025measuringmemorizationlanguagemodels. That work performed one experiment on non-training data on a small Pythia model. The paper uses the results of this experiment as a basis to both validate the experiments on known training data in Pythia models, and to apply the methodology to much larger (and higher quality) Llama 1 and OPT models—i.e., to make claims about extraction of not-known-with-certainty training data, just as we do here with DeepSeek v1 67B, Qwen 2.5 72B, and Gemma 2 27B.
hayes2025measuringmemorizationlanguagemodels finds that “…the number of queries $n$ needed to generate unseen test data is orders of magnitude larger than for generating training data. [They] find that it is generally challenging to generate test data …This supports that, in [their] measurements of [probabilistic] discoverable extraction, matches between training-example targets and generated suffixes are almost surely due to memorization.” We observe the same pattern, though instead of discussing it with respect to the number of queries $n$, as elsewhere, we frame our discussion in terms of $p_{\bm{z}}$. In our results, the $p_{\bm{z}}$ that we observe on non-training data are many orders of magnitude lower than what we deem extraction success in our experiments on known training data. That is, our measurements for extraction are not erroneously capturing chance generation of non-training data. As in hayes2025measuringmemorizationlanguagemodels, this gives us assurance that we can use our approach on models for which we do not know exactly what was included in the training data. We discuss this further in Appendices E.5 and E.6.
E.4.1 Results on Phi 4
We run a large number of negative control experiments using Phi 4 [phi4]. As noted in Appendix D.1, this model was not trained on whole copyrighted books. It was trained on a mix of synthetic data, licensed websites, and public domain books. When running our sliding-window extraction procedure on this model for copyrighted books, we can expect (generally speaking, with important exceptions) that unique text from copyrighted books will not register as extractable with our measurements.
For our main results using the sliding-window procedure with $100$-token sequences (Appendix D.3), we include Phi 4 as one of the $10$ models that we evaluate. Here, in Figure 118, we show additional results for experiments on varied prefixes and $4$ in-copyright books from Books3. We are able to extract non-trivial amounts of public domain books from this model (e.g., Appendix D.3.17). In contrast, we are not able to extract much unique text from in-copyright text. As elsewhere, we extract boilerplate text like copyright notices and publishers’ addresses. And we are also able to extract famous or otherwise popular quotes—either quoted in in-copyright books, or from the books themselves.
We (painstakingly) verify that every sequence that we extract from Phi 4 can be easily found in multiple sources on the Internet. That is, the quotes from books that we extract from Phi 4 are either in the public domain, quoted widely on blogs or platforms like Goodreads, or both. This is indicative (though not definitive) that these quotes could have found their way into Phi 4’s training data from these online sources, rather from the books themselves.
For further support, we can also compare the text at the book locations for which we extract these sequences to heatmaps other models. We note that, almost always, these sequences are extractable (sometimes with enormous probability) from all models we test. Again, while this is not definitive, it is suggestive that this text is highly duplicated in sources that are a part of common pre-training corpora—that they are not necessarily memorized because they are in Books3, but because of their duplicated inclusion in the training data from other sources.
We also observe similar challenges—to a somewhat lesser extent—for our negative control experiments on books published after LLM training-date cutoffs. These books sometimes cite earlier sources—other books, news articles, social media posts, etc.—that pre-date the training-date cutoffs and are widely duplicated online. So, even this seemingly “clean” case also requires significant manual verification to confirm validity, which we discuss in Appendix E.4.2.
Altogether, it is reasonable to conclude that these are true instances of extraction. These results reinforce that our procedure detects true memorization when it exists and does not spuriously flag text where memorization is impossible. Even so, these results show just how challenging it is to run completely clean negative controls for models trained on web-scraped data [lee2023explainers], for which duplicates of the same piece of text across various sources complicate what training-data membership means. A given book may not have been a member of the training data, but this does not mean that all text from that book was not. In these cases, it seems extraordinary likely that these sequences are members of the training data (and memorized). We discuss this further in Appendix E.6.
With the exception of the copyright notice, we are unable to extract any sequences from This Is How You Lose Her [loseher]. The sequences that we extract from Beloved are all posted (multiple times) on platforms like Goodreads. We discuss some of these quotes in Section 6.2 [belovedquote1, belovedquote2, belovedquote3]. Harry Potter and the Sorcerer’s Stone is excerpted and quoted in an extraordinary number of places on the Internet. We were able to find every extracted sequence on $3$ different web pages (after which we stopped counting).
Figure 118: Phi 4 Negative Control: Sliding-window procedure for various prefix lengths. We show results for $4$ books from Books3: Beloved [beloved], Harry Potter and the Sorcerer’s Stone [hp1], This Is How You Lose Her [loseher], and We Were Eight Years in Power [wewere]. Phi 4 was deliberately not trained on whole copyrighted books. We run probabilistic extraction on Llama 3.1 70B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. Extraction signal generally increases as prefix length increases. However, this only appears to be the case if there is indeed more memorization to surface; increasing prefix length does not always make previously unextractable sequences extractable. There are also diminishing increases as the prefix lengthens. We manually verified that every extractable sequence comes from highly duplicated text that can easily be found on the Internet; such sources reasonably were a part of Phi 4’s training data.
We discuss When We Were Eight Years in Power [wewere] in Section 5.2. As noted there, regardless of prefix length, there are only $4$ areas—other than the copyright notice at the beginning of the book—that show extractable sequences. The sequences at the end of the book contain a list of ordered (page) numbers. This kind of list is a very common artifact from converting an epub file to txt—as was done for the collection of Books3 [reisnerbooks3]. Like the copyright notice, such ordered lists of numbers are text that, while high probability, we can discount as being particularly interesting with respect to our negative controls. The other $3$ are all text that is not unique to book, but rather highly duplicated text from the Internet (some of which is in the public domain):
  • $\sim\!176$K: An exact quote from Malcom X’s speech, “Who Taught You To Hate Yourself?” (May 5, 1962) [malcolmx]
  • $\sim\!280$K: An exact quote of Deuteronomy 15:125 from the Bible – King James Version (KJV) [deuteronomy].
  • $\sim\!546$K: A exact quote from Barack Obama’s 2004 DNC Keynote: “hope of a young naval lieutenant bravely patrolling the Mekong Delta; the hope of a mill worker’s son who dares to defy the odds; the hope of a skinny kid with a funny name who believes that America has a place for” [obama2004]
All of these sequences are also highly memorized for other models we test.
Figure 119: Post-Training Cutoff Negative Control: We measure the probability of generating target $50$-token suffixes for $50$-token prefixes (using top-$k$, as elsewhere) for $4$ books published in 2025: Careless People [careless], Great Big Beautiful Life [greatbig], The Emperor of Gladness [emperorgladness], and The Society of Unknowable Objects [unknowableobjects]. All of these books were published after the release dates for all models we evaluate. These experiments serve as a validity check that our measurements for probabilistic extraction are capturing true instances of extraction. By definition, no unique sequences in these books can be memorized because they are not members of the training data; we should not be able to (and do not) obtain high, non-zero probabilities—probabilities that we would consider to be successful instances of extraction for known or suspected training data. Note that, in contrast to our other heatmaps, we set a more permissive (i.e., lower) minimum for $p_{\bm{z}}$ ($p_{\bm{z}}\geq 0.00000001\%$, or $1$ in $10$ billion generations), in order to see if there are sequences that can be generated even with fairly low probability that we would not count as successful extraction.
E.4.2 Results on books from after training cutoffs
We run additional negative control experiments on $4$ books that post-date the training of all models. As a result, these books could not have been included in the training data; we similarly should not expect to extract unique text from them for the different models we evaluate. However, as with our experiments on Phi 4, there are again cases where boilerplate text and popular quotes from other sources come up in our measurements.
Nevertheless, we are not able to extract any text unique to any of these $4$ books for any model we test. In Figure 119, we show results for all $4$ books using $100$-token sequences ($50$-token prefixes ++ $50$-token suffixes) for the $10$ models that we include for our sliding-window experiments on Books3 books (Appendix D). Note that in Figure 119, we expand the range of $p_{\bm{z}}$ that we plot to orders of magnitude below $\tau_{\text{min}}{=}0.1\%$. Even so, the only regions that register with non-zero probability are a select few that surpass $\tau_{\text{min}}$, and they are duplicated generic text—in this case, the copyright notice—and some references drawn from pre-2025 sources. For example, for The Emperor of Gladness [emperorgladness], we extract a verbatim quote from the Bible (Mark 5:13). For Great Big Beautiful Life [greatbig] and The Society of Unknowable Objects [unknowableobjects], we only extract boilerplate text.
For Careless People [careless], we extract the “careless people” quote from The Great Gatsby [gatsby] (Figure 1). It is quoted at the beginning of the book (e.g., $p_{\bm{z}}{=}2.8\%$ for Llama 3.1 70B):
[…]
This is what we register as extracted close to the 0-character position in the heatmaps. (We highlight the portion from The Great Gatsby in blue.) The book also quotes Mark Zuckerberg in several locations—quotes that we confirm are often drawn from Facebook posts that are widely duplicated on social media platforms and in online newspaper articles. These make up the remaining extracted sequences. For example, one quote is from a 2016 post about a SpaceX rocket explosion [spacex] (e.g., $p_{\bm{z}}{=}13.5\%$ for Llama 1 65B):
[…]
    Suffix: hear that SpaceX’s launch failure destroyed our satellite that would have provided connectivity to so many entrepreneurs and everyone else across the continent. Fortunately, we have developed other technologies like Aquila that will connect people as well.
(We highlight the quoted portion from The Great Gatsby in blue; even just a few words from the quote in the prefix are sufficient to generate the suffix with high probability.) Another is from a widely quoted (in news articles) Facebook press release:
[…]
An example of a sequence below $\tau_{\text{min}}$—i.e., that we would not count as extractable, but that we include with our expanded range in Figure 119Llama 1 65 exhibits $p_{\bm{z}}{=}0.02\%$ for a Mark Zuckerberg post about the Charlie Hebdo headquarters attack:
  • need to reject—a group of extremists trying to silence the voices and opinions of everyone else around the world. I won’t let that happen on Facebook. I’m committed to building a service where you can speak freely without fear of violence [hebdo]
Mark Zuckerberg’s manifesto is also quoted [manifesto].
We know the The Great Gatsby quote is in Books3, and widely quoted elsewhere. Similar to our observations for Phi 4, it is very reasonable that these other highly duplicated quotes are similarly in the training data, and that these are true instances of extraction.
Overall, consistent with hayes2025measuringmemorizationlanguagemodels, we do not observe chance generation of non-training data until we significantly lower the probability threshold. In Figure 119, we observe no such instances for $p_{\bm{z}}\geq 10^{-8}\%$, which is $7$ orders of magnitude smaller than our chosen $\tau_{\text{min}}$. (This is an upper bound; this choice was arbitrary for this plot.)
We also include the last book, Great Big Beautiful Life [greatbig], in our varied prefix experiments as another point of comparison. We discuss these results for Llama 3.1 70B in Section 5.2, and provide the full set of results for $5$ models in Figure 120. We extract no sequences (beyond boilerplate text) from any of the $5$ models we test. In Figure 7, we show an extended $p_{\bm{z}}$ range to highlight how low we need to go beneath $\tau_{\text{min}}$ to register chance generation of non-training data. We extend to $10^{-14}\%$$13$ orders of magnitude below $\tau_{\text{min}}{=}0.1\%$—and find exactly $1$ sequence that surpasses this threshold for 200200800800-token prefixes.
For example, for $800$-token prefixes, we register $1$ sequence with $p_{\bm{z}}{=}6.06\times 10^{-14}\%$. We report this as on the order of $10^{-14}\%$ in the main text because that is all we can reliably do for a number this small. The mantissa ($6.06$) is unreliable at this magnitude. This is because, even though we perform all $\log$ probability computations on CPU in full float64 precision, producing the logits during inference on GPU (that we post-process on CPU) involves mixed-precision computations with the model weights/activations in bfloat16 [gdm_bf16_range] and accumulation operations (matrix multiplications, normalization) in float32 [hf_bf16_accum]. Each computation introduces small rounding errors that accumulate and propagate to the final logits. When we exponentiate to to get probabilities, these errors result in only the exponent (not the mantissa) being reliable.
(a) Llama 3.1 70B
(b) Llama 3.1 70B
(c) Llama 2 13B
(d) Phi 4
(e) Qwen 2.5 72B
Figure 120: Post-Training Cutoff Negative Control: Sliding-window procedure for various prefix lengths. We show results for Great Big Beautiful Life [greatbig], which was published in spring 2025—after the release of all of the models we test. (The newest model is from the Qwen 2.5 series, which was released in late September 2024.) This book is not in any of the models’ training data. We run probabilistic extraction on (a) Llama 3.1 70B, (b) Llama 3.1 8B, (c) Llama 2 13B, (d) Phi 4, and (e) Qwen 2.5 72B with top-$k$ (T=1T{=}1, k=40k{=}40) decoding for $50$-token suffixes and prefix lengths in $\{25,50,100,200,400,800\}$. We are unable to extract any unique (i.e., non-copyright notice) sequences using any model or prefix length.
E.5 Limits on “the monkey at the typewriter” metaphor
A common trope is to compare an LLM to a “monkey at the typewriter” [borelmonkey, simpsons]. The basic idea is that, given infinite time, an LLM could presumably output any continuation of a piece of text. Like the monkey, it could by happenstance output a pristine copy of a Shakespeare play. This metaphor seems founded in the idea that LLMs are probabilistic. The monkey is randomly typing letters, and so, too, is an LLM, since there is an element of randomness in the generation process. For a given piece of text (the context), the LLM samples the next token to output based on the next-token conditional probabilities. That sampling is random, with respect to the conditional probability distribution of the possible next tokens.
Already here, this description of what an LLM is doing to generate text should start to sound very different from a “monkey at the typewriter.” In that thought experiment, what the monkey is doing is random in the sense that every next character is equally likely to be produced at every step. But clearly, that is not what is happening with an LLM. The LLM has learned probabilistic relationships from the natural-language and other text (e.g., code) it was trained on. These relationships condition outputs given inputs. So while there is sampling involved—and thus probabilistic outcomes—those probabilities are not “random” in the same sense as the monkey typing. If an LLM could be used to produce a copy of a Shakespeare play (or Harry Potter and the Sorcerer’s Stone, Section 7 & Appendix F), it would not be because it behaves randomly like the monkey. (One could set the temperature to an enormous number, effectively inducing “monkey” behavior via the decoding scheme, but this is not how LLMs are used in practice.)
A “monkey” or a “toothpick”?
These probabilistic relationships can also be easily misunderstood in the opposite direction. One could imagine that a sufficient amount of context used as input would serve to bias the LLM’s output to a particular, deterministic outcome: with a long enough prompt of the user’s choosing, they could get an LLM to produce any particular, desired target output. This is the basic idea in Tyler Cowan’s “toothpick argument”:
  • If you stare at just the exact right part of the toothpick, and measure the length from the tip, expressed in terms of the appropriate unit and converted into binary, and then translated into English, you can find any message you want. You just have to pinpoint your gaze very very exactly (I call this “a prompt”).
    In fact, on your toothpick you can find the lead article from today’s New York Times. With enough squinting, measuring, and translating.
    By producing the toothpick, they put the message there and thus they gave you NYT access, even though you are not a paid subscriber. You simply need to know how to stare (and translate), or in other words how to prompt.
    So let’s sue the toothpick company! [toothpick]
The “toothpick” misunderstanding of an LLM is mutually exclusive with the “monkey” misunderstanding. For the toothpick, prior context has a determinative impact on the output; for the monkey, there is no context long enough that could ever have a determinative impact on the output. Note that in the toothpick metaphor, it should also be possible to get an LLM to generate outputs that are exact copies of non-training data—in this case, “the lead article from today’s [paper]” (emphasis added). (One could, for an open-weight model, manually bias logits. This could force an LLM to produce certain outputs, effectively inducing “toothpick” behavior, but this is also not how LLMs are used in practice.)
Where our results come in.
There is plenty to say about misplaced metaphors about generative AI. We defer to others who have written extensively about this [lee2023talkin, cooper2023report, cooper2024files]. Here, we highlight how our specific results make specific points clear about why these two particular metaphors are both misguided. The key point is that our experiments show that (a) there will be at least some target suffixes that could never be generated by an LLM, even with infinite prompts to the model with a given prefix, and (b) a specific desired target suffix cannot always be produced by an LLM, given a sufficiently long prompt. The evidence for both comes from the same results: for a given LLM $\theta$ and decoding scheme $\phi$, some suffixes are impossible to generate, regardless of (a) how many times one independently prompts the model with a given prefix or (b) how long of a prefix one uses to condition the output.
Our results make this very clear in practice, but we can also make this point conceptually. Intuitively, for a sequence $\bm{z}$}, if the model $\theta$ with decoding scheme $\phi$ assigns the verbatim suffix (given the prefix) zero probability (i.e., $p_{{\bm{z}}}{=}0$), then no amount of independent sampling trials will ever produce that suffix. If instead $p_{{\bm{z}}}>0$, then repeated independent sampling ensures that the probability of seeing $\bm{z}$} at least once tends to $1$ as the number of trials grows. Therefore, for any finite set of sequences $\mathbb{Z}$}, the maximum number of verbatim suffixes that can ever be generated is exactly those with nonzero probability:
$\lim_{n\to\infty}\mathbb{E}\!\left[\#\text{ unique verbatim suffixes produced}\right]\;=\;\big|\{{\bm{z}}\in{\mathbb{Z}}:p_{{\bm{z}}}>0\}\big|.$
Proof.
For a single sequence $\bm{z}$} with per-trial, verbatim-suffix generation probability $p_{{\bm{z}}}$, the probability of generating the suffix at least once after $n$ independent trials is
$\Pr(\text{verbatim suffix produced in }n\text{ independent trials})\;=\;1-(1-p_{{\bm{z}}})^{n}.$
If $p_{{\bm{z}}}=0$, this expression is 0 for all $n$. If $p_{{\bm{z}}}>0$, then $(1-p_{{\bm{z}}})^{n}\to 0$ as $n\to\infty$, so $\Pr(\text{verbatim suffix produced in }n\text{ independent trials})\to 1$. Now let $\mathbb{Z}$} be a finite set of sequences. By linearity of expectation,
$\mathbb{E}[\#\text{verbatim suffixes produced in }n\text{ independent trials}]\;=\;\sum_{{\bm{z}}\in{\mathbb{Z}}}\big[1-(1-p_{{\bm{z}}})^{n}\big].$
Taking $n\to\infty$,
$\lim_{n\to\infty}\mathbb{E}[\#\text{verbatim suffix produced in }n\text{ independent trials}]\;=\;\sum_{{\bm{z}}\in{\mathbb{Z}}}\mathbf{1}[p_{{\bm{z}}}>0].$
Therefore, for this fixed finite set $\mathbb{Z}$}, there is a maximum attainable suffix (given prefix) generation rate: the fraction of $\mathbb{Z}$} whose suffixes can ever be generated. This limit is independent of the target success level $p$ in $(n,p)$-discoverable extraction (Definition 1). As $n$ grows large, the rate reflects not only extractable (memorized) sequences—i.e., those with sufficiently high probability such that we consider them to be instances of extraction (in this work, $\tau_{\text{min}}\!\geq\!0.1\%$)—but also any $\bm{z}$} with $0<p_{{\bm{z}}}<\tau_{\text{min}}$ that we would not count as extraction.
Sequences that cannot be generated in practice.
There are various ways that a given suffix could have $p_{\bm{z}}{=}0$. It might be the case that, for a particular decoding scheme, it is impossible to sample certain tokens, making some suffixes unreachable. This is the case for top-$k$ decoding, which only considers the top-$k$-highest probability tokens at each sampling iteration. Any token not included in the top-$k$ tokens forecloses possible output generation paths (Appendix A.3). For $k=40$ and for Llama 1 models, only $40$ out of 32,00032{,}000 tokens could be generated in a given sampling iteration.
For another, numbers are represented in computers with a finite number of bits. This means that there is a minimum number that can be represented (>0>0) before it gets rounded down to 0. For instance, imagine a token vocabulary of 32,00032{,}000 tokens (as with Llama 1 models). If the model just generated tokens completing randomly (by setting temperature very high, like a “monkey at the typewriter”), then at each iteration every token would have probability $\frac{1}{32000}$. When generating $71$ such tokens in a row (i.e., a $71$-token suffix), the probability of any such suffix is ${\frac{1}{32000}}^{71}\approx 1.3626\times 10^{-320}$. Once we try to generate a 7272nd token, the probability for the whole sequence becomes ${\frac{1}{32000}}^{72}=0$. In terms of how the probability gets represented in the computer, there is underflow (i.e., rounding down to 0). As a result, even with completely-random temperature sampling, not every sequence could possibly be generated.
We often observe 0-probability suffixes regardless of prefix length. This is true for non-training data (Section 5.2, Appendix E.4.2) and for training data that are not memorized (Section 5.1, Appendix E.3). Even if we add more context to statistically bias the output—to try to condition it in a particular direction—there are many cases where we simply are unable to raise $p_{\bm{z}}$ above 0 (let alone above $\tau_{\text{min}}$, the minimum probability we consider evidence for extraction).
E.6 Additional notes on extraction and membership inference
There are several points that are worth addressing with respect to the broader literature on membership inference and extraction. Here, we briefly situate our work—particularly our work with respect to validity—in relation to prior work. This is important to do for clarity reasons, because our work and the types of claims we make are very different from prior work. Prior work tends to focus on average extraction rates over random samples of known or proxy training data. Since we are making claims about specific works from a specific population of text, the ways that we go about evaluating and validating our work are necessarily different. This is not just a detail or aside; it has to do with the entire purpose of this project, and significantly influences our measurement approach in ways that contrast with other work on membership inference.
E.6.1 The relationship between membership inference and extraction
As noted in carlini2021extracting, “Given a set of samples from the model, the problem of training data extraction reduces to one of membership inference: predict whether each sample was present in the training data.” This makes sense intuitively. Extraction is defined with respect to memorization: it is only possible to extract memorized training data, and (tautologically) training data can only be memorized if they are included—i.e., are members—of the training dataset. To demonstrate extraction is therefore to demonstrate memorization, and memorization implies membership.
Not only does extraction imply membership of a piece of data, it actually reproduces that data in model outputs. In this respect, it is a stronger evidentiary claim (and, at least in principle, a harder one to make): membership inference asks a one‐bit question—was this sequence a member of the training data?—while extraction produces the sequence itself. (At the same time, reliable membership inference with even the strongest membership inference attacks is extraordinarily challenging for LLMs [hayes2025strongmia], while extraction is relatively straightforward.) Put differently, one can make membership inference claims without extraction (e.g., with classifiers, see below); these claims can involve evidence where extraction might fail. So, even though successful extraction always implies membership, this does not go in both directions: membership does not imply successful extraction (Appendix E.6.5).
carlini2021extracting goes on to note that “In their most basic form, past membership inference attacks rely on the observation that models tend to assign higher confidence to examples that are present in the training data. Therefore, a potentially high-precision membership inference classifier is to simply choose examples that are assigned the highest likelihood by the model.” At a high level, this is clearly directly related to extraction. Greedy‐decoded discoverable extraction approximates the highest‐likelihood sequence (albeit poorly) with greedy decoding [carlini2023quantifying]. Probabilistic extraction does the same but allows more flexibility than greedy decoding, and can identify sequences with higher likelihood than greedy‐decoded ones [hayes2025measuringmemorizationlanguagemodels].
E.6.2 We know ground-truth membership for most of our experiments
In much work on extraction, researchers do not demonstrate extraction to infer membership. Instead, they run extraction experiments on known training data [hayes2025measuringmemorizationlanguagemodels, carlini2023quantifying, lee2022dedup, hayes2025strongmia], i.e., where ground-truth membership is not in doubt. This makes sense as a setup for drawing valid conclusions about extraction and memorization: membership is a prerequisite for memorization, and memorization is a prerequisite for extraction. Testing extraction on known member sequences therefore enables straightforward conclusions about memorization. Here, memorization is the most plausible explanation for extraction. As Carlini explains, when a sufficiently large and unique piece of training data is extracted, “the only possible explanation is that the model has somewhere internally stored [that piece of training data]” [carlini2025blog]. As Cooper and Grimmelmann similarly note, “in order to be able to extract memorized content from a model at generation time, that memorized content must be encoded in the model’s parameters. There is nowhere else it could be. A model is not a magical portal that pulls fresh information from some parallel universe into our own” [cooper2024files, p. 25]. (See Section 2.)
Most of our experiments fall into this category, as the large majority of our results are for Llama and Pythia models, where it is known that Books3 was included in the training data (Sections 46; Appendices D.1 & E.2). This is precisely the motivation behind our baseline experiments (Section 5.1, Appendix E.3), which we use as a basis of comparison for our negative controls (Section 5.2, Appendix E.4). Not only do we know that these models were trained on the books in Books3, we know that they were trained precisely on the Books3 corpus; we know the exact format of these data, just as they were presented to each of these models during training.
Nevertheless, while work on greedy-decoded discoverable extraction does not do this, we perform validity experiments anyway. We need to set a minimum valid extraction probability $\tau_{\text{min}}$ (Appendix E.1). In this paper, we use a very high global $\tau_{\text{min}}$ to avoid false positives, and we validate this choice with our negative controls. hayes2025measuringmemorizationlanguagemodels do not set an explicit $\tau_{\text{min}}$, and do not make strong claims about when probabilistic extraction is valid; however, they do run a negative control experiment to show more generally that probabilistic extraction is a reasonable metric.
In the original version of this paper, we included the negative controls on Phi 4 (Appendix E.4.1) to support our claims. Combined with the extensive analysis we perform on known training data for Llama, this is how we chose a global $\tau_{\text{min}}$: a conservative threshold that would not over-count extraction (and thus memorization). Even before including our experiments on post-cutoff books, our results provided ample evidence that non-training data exhibit 0 or orders of magnitude smaller probabilities $p_{\bm{z}}$ than training data. The post-cutoff negative controls provide further confirmation.
This was true both for smaller, lower quality models, and higher quality models that exhibit lower loss (on average). While larger models like Llama 3.1 70B fit their training data better than smaller ones, and exhibit lower loss on average, they do not (and cannot) exhibit low loss everywhere, such that our measurements of $\log$ probabilities would be confounded. This is also clear from our experiments on known training data—Books3 books. For many books and large models, we are unable to extract any unique sequences; there are effectively no nonzero sequence probabilities. It is plausible that future work will find lower, model-specific $\tau_{\text{min}}$ thresholds, which may perhaps need to be set higher for larger models than for smaller ones. For now, however, the threshold we set is so conservative that these nuances do not affect our conclusions.
E.6.3 We are inferring membership for models with undisclosed training data
In earlier versions of this paper, we placed our validity discussion in the appendix, but now highlight it in the main text (Section 5). Given our extensive results on known training data—including experiments to vet their validity—we can apply our methodology with confidence to models whose training data are not known with certainty. We do this for three model families: DeepSeek v1, Qwen 2.5, and Gemma 2. For these models, validated extraction serves as evidence of membership. These experiments differ from those on models known to be trained on Books3, but we ensure their validity through extensive negative controls and baselines.
This approach is consistent with prior work. carlini2021extracting performed experiments on GPT-2. nasr2023scalable, nasr2025scalable extracted training data from open-weight models with undisclosed training data and from ChatGPT, a closed system. In both cases, the authors relied on proxy datasets and strategies to reduce bias in their results, such as shorter prompts. More generally, this is an important and reasonable thing to try to achieve. As researchers, we should be developing valid techniques for extraction that are useful in cases for which we do not know ground-truth membership. If membership inference required knowing membership in advance, there would be nothing to infer; at best—when it works—it would confirm what we already know.
We do not need to use shorter prompts in our work because we validate our work with extensive negative controls. Further, we verified that every sequence extracted from Phi 4 and from post-cutoff books was widely duplicated online. Our validity experiments go beyond most prior work, which often does not include negative controls. This is intentional; our work has the potential to impact copyright litigation, not just the scientific literature. We therefore believe it is important to hold our work to a higher bar. Importantly, since training data for these models are undisclosed, we make claims only at the sequence level, not at the level of whole books (Appendix E.6.4).
Finally, the population of data we study differs from most prior work, which evaluates memorization broadly across diverse training data (websites, code, spam, etc.) [carlini2021extracting, carlini2023quantifying, nasr2023scalable, nasr2025scalable, hayes2025measuringmemorizationlanguagemodels]. Such work must account for false positives like “trivial memorization” (e.g., number lists) or “repeated substrings” (e.g., “I love you” loops). LLMs assign high probability to both of these types of sequences, even if they are not memorized. We note these phenomena explicitly where we identify them, but they are generally not relevant our domain of long-form narrative text. The rare cases we identify of trivial memorization—lists of page numbers (an artifact of converting epub to txt), chapter indices—are excluded from our claims about extraction of unique creative expression. We identify one case of repeated substrings. bedwetter exhibits elevated probability in a part of the book that repeats “I will not wet the bed.” four times in a row. But this sequence does not surpass $\tau_{\text{min}}$, anyway, so we do not count it as extracted. (We also exclude boilerplate instances of true memorization from our claims, like copyright notices and author biographies.) Across more than one million sequences, these account for only a few hundred cases. This highlights why the population under study matters for validity: our focus on books yields different considerations than prior work, and we address these systematically. Unlike the work done in carlini2021extracting, these types of edge cases are extremely rare in our population of interest.
E.6.4 Our inferences about membership are about sequences, not whole books
We emphasize that our claims about membership are made at the level of sequences, not entire books. For models such as DeepSeek v1, Qwen 2.5, and Gemma 2, where training data are undisclosed, we cannot establish whole-book membership. Instead, our extraction results are about specific sequences drawn from Books3, a known pretraining dataset, and it is at this level that we make membership claims. In some cases, our results are suggestive that entire books were included in training, given that we can extract high-probability sequences spanning a full work (e.g., Appendix D.3.39), but we do not generalize from sequences to whole-book membership.
We also do not claim that we know the provenance of those extracted sequences. For Llama models we know that the training data included Books3, and thus at least one copy of each sequence from the books we test was included in the training data via Books3. But this is about all we do know with certainty. We do not know very much about these models’ training data details more generally; knowledge of the inclusion of Books3 is a rare exception, not the norm. Books sequences may have been included in the training data via other unknown sources: other book corpora, widely quoted excerpts in web scrapes, etc. The same could be said for these other models, where we do not know if Books3 was included in the training data.
Altogether, these observations highlight that membership questions for LLMs are rarely straightforward. This is already widely known in the literature [e.g., lee2022dedup], with respect to the presence of duplicates in training datasets [carlini2023quantifying, hayes2025measuringmemorizationlanguagemodels]. This discussion illustrates how these challenges manifest specifically for books, and why membership claims should be scoped carefully. While our measurements are at the level of individual sequences, but that is not necessarily the only level at which membership is of interest. For copyright, membership may be meaningful in different ways at different scopes—an entire book, a chapter, a single sentence, etc. [lee2023talkin].
E.6.5 Absence of extraction does not imply anything about membership
We state this with respect to describing the problem of membership inference in Appendix E.6.1, but it bears repeating: extraction implies membership, but membership does not imply extraction. This is clear in our results; many books in Books3 have no extractable unique text for Llama models, even though those models were trained on Books3. So when we see $p_{\bm{z}}{=}0\%$, we should not conclude that a given sequence (or the source that contains it) was not included in the training data. This is trivially true for Llama models, but more generally true for models with undisclosed training data. For example, Qwen 2.5 may have been trained on Books3; just because it exhibits low extraction probabilities on Books3 books is not evidence that it was not trained on this corpus.
Further, $p_{\bm{z}}{=}0\%$ does not necessarily imply anything about how much a particular sequence may have influenced the model or its outputs. Attribution of generations to training data is a different problem, both with respect to technical work in ML and implications for copyright. We do not study these topics in this work. At a very high level, one can see this distinction through our brief results in Appendix F.3. We run our reconstruction procedure on Sandman Slim [sandmanslim] and Llama 3.1 70B. This model does not memorize this book at all, with respect to our extraction procedure. We run this experiment to show that it fails—that reconstruction is not something that trivially works for any book. Nevertheless, even though we fail to reconstruct any text from the book, it is clear that the generation we produce was influenced by the text (in an informal sense). This type of influence is not something that we capture or attempt to capture—either formally or informally—in our measurements in this paper.
E.6.6 Extractability and copyright
Separate from the above technical considerations about membership and extraction, there are other considerations for copyright. A sequence can count as memorized in a technical sense—in this paper, when $p_{\bm{z}}\!\geq\!\tau_{\text{min}}\!=\!0.1\%$—but this does not necessarily mean that all types of memorization are relevant to copyright. As we discuss in Sections 3 and 8, sequences that have very low extraction probabilities may matter less for copyright than those with high ones, since the latter are far easier to extract reliably in model outputs.
Appendix F Reconstructing books with one seed prompt
Given the extent of memorization that we observe for Harry Potter and the Sorcerer’s Stone [hp1] (as well as other books) and Llama 3.1 70B, it seemed possible that we could reconstruct the entire book near-verbatim using only a single seed prompt of ground-truth text drawn from the book.
This result was a natural follow-on to our memorization experiments—based on the results we obtained—but it is not the main focus of our paper. Therefore, this part of our paper reflects only a preliminary investigation into reconstructing large book segments from a model. It is an existence proof that this type of result is possible, not a comprehensive investigation into how to make this type of method work for different texts.
F.1 Reconstruction method
We successfully reconstruct the book near-verbatim with a variety of seed prompts and variations on the same strategy. Overall, we observe that this result is stable—across different hardware, and for different settings. However, the procedure is relatively sensitive to how it is configured; for example, if too long of a stride is used, then it we observe divergence from the true text. We have preliminary results that significantly improve the stability of this approach. We defer this to future work.
Seed prompt.
We obtained high-fidelity reconstructions of the book with several seed prompts, including the first $390$ tokens of the the first chapter (the first few paragraphs) and the first line of the first chapter ($60$ tokens):
  • Mr. and Mrs. Dursley, of number four, Privet Drive, were proud to say that they were perfectly normal, thank you very much. They were the last people you’d expect to be involved in anything strange or mysterious, because they just didn’t hold with such nonsense.
[…]
Overall approach.
We take a sliding-context approach that we discuss in Section 7. We take an initial seed $s$ and top it up to a minimum of $50$ tokens (if it is less than that). We use beam search to generate $50$ tokens at a time, and use a maximum context window of $3000$ tokens for each generation. This means that we append the $50$ tokens we generate to the existing context, and clip the first $50$ tokens for iterations that exceed the maximum context. (This is most of the iterations; only at the start do we have a context that can be less than $3000$ tokens).
We succeeded with $9$ tries in total. (Our first major success occurred in the third run, when we recovered ten chapters—chapters 1-9, and 12—near-verbatim.) The first $3$ runs were debugging and setup (e.g., testing the key idea with greedy decoding and getting a sense of how short a reasonable seed prompt could be). The last $6$ involved changing the size of the number of maximum context tokens (we settled on $3000$) and the number of beam candidates ($8$), as well as minor tweaks to handling EOS tokens.
Handling EOS tokens was the only (minor) complication, and it was simple to resolve. EOS tokens tend to be predicted at the ends of chapters; we therefore remove them and manually replace them with "CHAPTER {n+1}", with {n+1} spelled out (e.g., "TWO", "THREE"). Other attempts involved "Chapter" (instead of all caps) and using the number {n+1} (e.g., "2") instead of spelling out the chapter number (e.g., "TWO"). Other solutions could just down-weight the logits for EOS.
Occasionally, the model does not predict an EOS at the end of a chapter, which leads to some misalignment later on (e.g., inserting "CHAPTER FOUR" for the fifth chapter). This can cause the model to repeat segments of chapters that have already been produced. The fix here was also simple. We keep track of how many tokens have been generated since the last EOS token, and if that number surpasses 10,00010{,}000, we assume that we have missed the end of a chapter and account for this (i.e., move the chapter counter ahead even though we did not see an EOS). In future runs, we were able to remove this logic entirely. We defer discussion to future work.
The exact code that we ran is in Listing 3. Since this code involves beam search, our results should be (in theory) deterministic. Because in practice hardware non-determinism can complicate this, we ran a (non-exhaustive) test by executing the code on two different sets of $4$ GPUs. Both yielded the same exact output. (Though we also spot-checked some of the underlying logits, and there were some slight differences.)
Comparing to the ground-truth text.
The substantive diff between the version of Harry Potter and the Sorcerer’s Stone [hp1] that we have from Books3 and our generated output is minimal. We do some very minor text normalization of the ground-truth book. That is, we trim the beginning and end of the ground-truth book to not contain front matter or anything after the end of the last chapter. We remove _ from both documents (used for italics in the Books3 version) and align ellipses to "..." (which appear as " . . ." in the Books3 version.)
We then compute various similarity metrics betwen the two documents. The cosine similarity of TF-IDF vectors of each document is an astounding $0.9999$. TF-IDF is a limited metric, as it treats documents as bags of words and thus fails to capture word order. So, as two additional points of comparison, we compute similarity using greedy longest common subsequence (LCS) matching at both the word and sentence levels (via difflib’s SequenceMatcher). A score of $1$ indicates a perfect match and 0 indicates no match. We obtain a word-level similarity of $0.992$ and a sentence-level similarity (which is more sensitive to formatting differences) of $0.934$.
The qualitative differences involve minor formatting changes (that we do not normalize) and localization differences. For example, the Books3 version has British spelling (e.g., "Mummy" instead of "Mommy"). These formatting and localization differences are the large majority of the diff between the two texts. Occasionally, the model also misses single lines, e.g., “There was a loud ’Oooooh!”’ (Figure 9).
One of the most interesting outcomes (that we have not investigated in detail) is very occasional apparent hallucinations. For example, in the first chapter, we observe the following diff:
Figure 121: Portion of the diff from the first chapter of Harry Potter and the Sorcerer’s Stone. Crossed-out red text is in the ground-truth book in Books3, but was not generated by our reconstruction procedure; text highlighted in yellow/green is in the reconstructed text but not the ground-truth text.
Selecting the top sequence candidate with beam search always led to this particular outcome in Figure 121, regardless of how we configured our method. Given how specific this particular change was, we investigated it further. It turns out that the generated text reflect the ground-truth from editions of the book printed before 2004. In later editions, this line was changed to the one that we do not generate—the one that is in the ground-truth version we have access to in Books3. This change was done for plot continuity reasons, based on an inconsistency that emerged due to plot development concerning Sirius Black in the later books [sirius]. This line is not actually hallucinated; it is just from an earlier edition of the book than the one we diff with. Of course, this difference still contributes to our quantitative measurements of similarity, even though it is not a meaningful substantive difference.
1from transformers import AutoTokenizer, AutoModelForCausalLM
2import torch, json
3from tqdm import tqdm
4
5# load the first line of HP 1
6with open("seed.txt", "r", encoding="utf-8") as f:
7 seed_text = f.read().strip()
8
9# Load the model and tokenizer
10model_name = "meta-llama/Llama-3.1-70B"
11tokenizer = AutoTokenizer.from_pretrained(model_name)
12model = AutoModelForCausalLM.from_pretrained(
13 model_name,
14 device_map="auto",
15 torch_dtype=torch.bfloat16,
16)
17model.eval()
18
19max_context_tokens = 3000
20max_new_tokens = 50
21# the book is more like 106k;
22# adding buffer for potential formatting differences
23max_story_tokens = 113000
24
25eos_token_id = tokenizer.eos_token_id
26
27input_ids = tokenizer(seed_text, return_tensors="pt").input_ids.to(model.device)
28generated_ids = input_ids
29
30# Set up progress bar
31pbar = tqdm(total=max_story_tokens, desc="Generating story tokens")
32pbar.update(generated_ids.shape[1])
33
34generation_steps = []
35generation_num = 1
36chapter_count = 1
37
38chapter_nums = {1 : "One", 2 : "Two", 3 : "Three", 4 : "Four", 5 : "Five", 6 : "Six", 7 : "Seven", 8 : "Eight", 9 : "Nine", 10 : "Ten", 11 : "Eleven", 12 : "Twelve", 13 : "Thirteen", 14 : "Fourteen", 15 : "Fifteen", 16 : "Sixteen", 17 : "Seventeen"}
39
40tokens_since_last_eos = 0
41
42while True:
43 # Get the tokens for the (up to) 3000-token prompt
44 slice_start = max(0, generated_ids.shape[1] - (max_context_tokens - max_new_tokens))
45 input_ids_window = generated_ids[:, slice_start:]
46 attention_mask = torch.ones_like(input_ids_window)
47
48 # call generation code
49 outputs = model.generate(
50 input_ids_window,
51 attention_mask=attention_mask,
52 max_new_tokens=max_new_tokens,
53 do_sample=False,
54 num_beams=8,
55 early_stopping=False,
56 length_penalty=1.2,
57 temperature=1.0,
58 top_p=1.0,
59 pad_token_id=eos_token_id,
60 )
61
62 # remove the prompt from the output
63 new_tokens = outputs[:, input_ids_window.shape[1]:]
64 new_tokens_list = new_tokens[0].tolist()
65
66 tokens_since_last_eos += len(new_tokens_list)
67
68 # Some fun tricks to deal with EOS tokens
69 if eos_token_id in new_tokens_list:
70 if tokens_since_last_eos >= 10000:
71 print("More than 10000 tokens since last EOS; probably missed a chapter break")
72 print("Incrementing chapter count")
73 chapter_count += 1
74
75 print(f"EOS: replacing with chapter break")
76 new_tokens_list = [t for t in new_tokens_list if t != eos_token_id]
77 chapter_count += 1
78 chapter_text = chapter_nums.get(chapter_count)
79 if chapter_text:
[…]
81 else:
[…]
83 chapter_tokens = tokenizer(chapter_text, add_special_tokens=False, return_tensors="pt").input_ids[0].tolist()
84
85 new_tokens_list.extend(chapter_tokens)
86 new_tokens = torch.tensor([new_tokens_list], device=model.device)
87
88 tokens_since_last_eos = 0
89
90 # Everything below is just saving and printing progress
91 prompt_tokens = input_ids_window[0].tolist()
92 prompt_text = tokenizer.decode(prompt_tokens, skip_special_tokens=True)
93
94 generated_ids = torch.cat([generated_ids, new_tokens], dim=-1)
95 pbar.update(new_tokens.shape[1])
96
97 chunk_text = tokenizer.decode(new_tokens[0], skip_special_tokens=True)
98
[…]
100
101 generation_steps.append({
102 "generation": generation_num,
103 "prompt_text": prompt_text,
104 "generated_text": chunk_text,
105 "total_generated_tokens": generated_ids.shape[1]
106 })
107 generation_num += 1
108
109 with open("generation_log.json", "w", encoding="utf-8") as f:
110 json.dump(generation_steps, f, indent=2)
111
112 with open("generated_ids.json", "w", encoding="utf-8") as f:
113 json.dump(generated_ids[0].tolist(), f)
114
115 if generated_ids.shape[1] >= max_story_tokens:
[…]
117 break
118
119pbar.close()
120
121# save the full story text
122full_text = seed_text + "".join(step["generated_text"] for step in generation_steps)
123
124with open("generated_story.txt", "w", encoding="utf-8") as f:
125 f.write(full_text.strip())
Listing 3: Reconstruction with a single seed prompt.
F.2 Running partial reconstruction for books with partial high memorization signal
We ran the exact same script on the essay “The Case for Reparations,” which is included in Ta Nehisi-Coates’ book, When We Were Eight Years in Power [wewere]. We observed enormous amounts of memorization for this essay in our sliding-window experiments. (See Section 5.1, highly memorized middle section; Appendix E.3). This was one of the books in the Kadrey v. Meta, Inc. [kadreyamendedconsolidated] lawsuit, for which the court in part ruled for fair use because neither party produced evidence that demonstrated sizable amounts of extraction of training data of the books in scope for the suit [kadreyjudgment].
We summarize the first run of this experiment, where we ran the exact same code that reconstructed the first Harry Potter on the first line of the essay ($31$ tokens, including BOS):
  • Clyde Ross was born in 1923, the seventh of thirteen children, near Clarksdale, Mississippi, the home of the blues.
[…]
F.3 Running reconstruction for books with low memorization signal
As a sanity check, we run the same method on Llama 3.1 70B for a book for which we observe effectively no extracted sequences.
For Sandman Slim [sandmanslim] (Appendix D.3.28), we use the first paragraph as the prompt and (including this prompt, shown in blue) produce a $2000$-token document.
Note that the generation gets caught in a loop (we mark where the prompt gets generated in blue later on in the generation.) Interestingly, even though we do not generate anything that we can identify as copied from the book, the generation does contain references to “Sandman Slim,” the main character “James Stark” (who is never referred to directly that way in the book, but as “James Butler Hickok Stark”) as well as another character called “Mason Faim.”
I WAKE UP on a pile of smoldering garbage and leaves in the old Hollywood Forever cemetery behind the Paramount Studio lot on Melrose, though these last details don’t come to me until later. Right now all I know is that I’m back in the world and I’m on fire. My mind hasn’t quite kicked in yet, but my body knows enough to roll off the burning trash and to keep rolling until I can’t feel the heat anymore. I come to a stop against a gravestone and lie there for a moment, trying to remember how I got here and who I am. I don’t always know right away. It’s the amnesia. It’s a hell of a way to travel, but it’s better than being dead. Slowly, like dirty water dropping out of a rusty pipe, it comes back to me. My name is James Stark. Friends call me Sandman Slim. Call me what you like. I’ve been called worse. I got this name because all I do is bring nightmares. True, I spend a lot of time in the City of Angels, but it’s not like they’re putting up a statue to me. Before this, I lived in Hell for eleven years. Even when I tried to stay out of trouble, trouble would come right to my door and make friends. That kind of thing has a way of making a man a little jaded when it comes to trusting his fellow man. Or demons. Even angels. I’m like that. I’ve got a bad attitude. I drink. I’ve been kicked out of every twelve-step program I’ve ever joined. I don’t have many friends, so the ones I have mean a lot to me. They mean enough that if someone hurts them, I’ll tear the universe down to save them. I’m telling you this so you’ll know who I am and what I’m talking about when I say there are things in the dark. Things we don’t always see that move among us. Some of these things are good and some aren’t and sometimes they do things that make you want to rip open your chest and pull out your heart because you’ve seen too much and there’s no going back. I know about these things because I’m one of them. I’m a monster and I’m a killer. I’m good at it and I never let emotion get in the way. When you live where I live, love can get you killed. And when it doesn’t, it’s worse because if you don’t die, you just wish you would and every night you pray for something you lost in Hell. Pray for sweet, sweet death to come make a man out of you. But death doesn’t come because if there’s one thing I learned in Hell, it’s that it will never, ever come for you. No matter how hard you pray. No matter what terrible things you do. But down deep in my heart, I know there’s another truth, one that even Hell can’t touch. And that’s why, no matter what, I can never do the one simple thing that would set me free. The one thing I want more than anything else. I can never, ever kill the man who murdered my friends and started me on this journey to Hell and back. No matter how many times I try, I can’t kill Mason Faim, the man I once called best friend. I wish I wasn’t me sometimes, but to tell you the truth, even when this is all over and they finally let me die, I wouldn’t trade this life for any other. It’s been one hell of a ride and I’d do it all over again if I had to. And I probably will. Over and over and over again. Because that’s all life is. One big ride that you can never get off of. You can just tighten or loosen your grip on the reins and see which way the horses take you. I WAKE UP on a pile of smoldering garbage and leaves in the old Hollywood Forever cemetery behind the Paramount Studio lot on Melrose, though these last details don’t come to me until later. Right now all I know is that I’m back in the world and I’m on fire. My mind hasn’t quite kicked in yet, but my body knows enough to roll off the burning trash and to keep rolling until I can’t feel the heat anymore. I come to a stop against a gravestone and lie there for a moment, trying to remember how I got here and who I am. I don’t always know right away. It’s the amnesia. It’s a hell of a way to travel, but it’s better than being dead. …
We try much longer prompts as well (e.g., the first two pages of the book), and the model does not continue with text from the book.
While these tests are not exhaustive, they are meant to illustrate that this is not behavior that we observe for all books with this approach.
Appendix G Discussion of extended results
We defer detailed discussion to future work. We simply reiterate our main takeaway, which is clear from the extended results we present in the Appendix. Memorization (as measured by a specific technique that focuses on probabilistic extraction, see Appendix A) varies significantly across models for the same book, and across books for the same model (Section 6 & Appendix D). Average extraction rates—the typical metric in ML memorization research—do not reveal this nuance. Instead, it is necessary to examine per-book extraction metrics to get this view.
We need to run more books to make general claims about overall memorization.
We only ran experiments for a small fraction of the entire Books3 dataset—running detailed experiments for $50$ books (Appendix D). We took care to sample books from a variety of sources—in copyright, in the public domain, openly licensed. We included popular books, as well as more obscure ones (Table LABEL:app:tab:sliding-books). Overall, very popular books exhibit the most memorization. It seems likely that these books are duplicated on different parts of the Internet; de-duplication is a challenging problem to implement in practice [lee2022dedup], so it is likely that least some duplicate text persists in training datasets for LLMs.
We do not know exactly what is going on with Llama 3+ models.
Even so, it seems unlikely that (relatively small numbers of) duplicates completely explain the patterns we observe. Llama 3.1 70B exhibits a lot more memorization than any other model. We plan to examine other large models in future work to see if there is a particular role of scale/ training dynamics at this scale with respect to just how much memorization we observe. However, Llama 3.1 70B generally exhibits higher amounts and degrees of extraction compared to Llama 3, Llama 2 70B, and Llama 1 65B (a model of a similar size). It also exhibits more memorization than a model that registers lower loss and superior performance on downstream benchmarks (Qwen 2.5 72B). That is, as our results also show for Llama 3.1 70B on books in the training data that it has not memorized, it is not the case that the model’s loss is “low everywhere,” and that this is responsible for the high suffix probabilities we observe. First, that is not how machine learning works; the model cannot so-strongly fit to all of the massive dataset it was trained on. Second, we do not observe this anyway for our results. And third, we do not observe the same degree of memorization in a model that exhibits lower loss on average than LLama 3.1 70B.
High-level memorization patterns.
Llama 1 65B and Llama 2 70B exhibit less memorization of Books3 text both in general (with respect to average extraction rate, see Appendix C) and on the specific books we test. (We do not test Llama 3 70B in our average extraction rate experiments.) In general, we observe a pattern that later generations of Llama models memorize more than earlier ones, with respect to average extraction rates (Appendix C) and extraction coverage (Appendix D.4.1).
Most books we test exhibited minimal memorization, measured with respect to probabilistic extraction with top-$k$ decoding (T=1T{=}1, k=40k{=}40) and $100$-token sequences ($50$-token prefixes ++ $50$-token suffixes). Our results on longer prefixes (Appendix E), however, indicate that there is a lot more memorization to uncover—even if it is harder to extract, where hardness is loosely equated with requiring a longer prefix.
Nevertheless, much of the memorization that we do observe for books in our main sliding-window experiments frequently falls into one of a few categories: copyright notices, lists of ordered (page) numbers (a common artifact of converting epub files to txt format, see Appendix E.4.1), publisher addresses, chapter number listings, and author biographies. All of these are types of text that are highly duplicated (partially or exactly) in various sources, both the Internet and across books. And so, extraction of a copyright notice from a given book does not necessarily mean it was memorized from that book; it was likely memorized due to the presence of numerous similar pieces of text in the training data. The same is also true for author biographies, which are printed on websites, not just in books.
Another category was text from popular sources that are quoted within books: quotes from the Bible [embraced, unglued, wewere, emperorgladness], philosophers like John Stuart Mill [zittrain], classics like those by Dante Alighieri [dante], text from U.S. government documents [zittrain, wewere], famous political speeches [wewere], public domain books [careless], social media posts referenced in numerous news articles [careless], and famous quotes posted on blogs and platforms like Goodreads [beloved]. In many cases, this was the only such text we were able to extract with non-trivial probability from some books using $50$ token prompts. For our negative controls, we painstakingly verify that each piece of text that we extract within a sequence can be found on multiple sources on the Internet.
With respect to the books that we tested that are within the scope of the Kadrey et al. v. Meta, Inc. class action suit [kadreyamendedconsolidated], we were not able to extract much memorized training data. There were notable exceptions (e.g., wewere). It is of course possible that another extraction technique could reveal additional memorization, but we were unable to meaningfully do so for most books using a $50$-token prefixes. It is also important to note that suit was recently decided in favor of the defendants (Meta), and that memorization of training data was only one issue being argued.
Our memorization measurements do not promise that reconstruction naturally follows.
We emphasize again that, even for highly memorized books, our sliding-window experiments are not extracting enormous amounts of text in one go, i.e., with a single seed prompt to a model. That is not what we focused on in this project. However, we do show that it is possible to reconstruct a highly memorized book by using an LLM autoregressively, starting with a single seed prompt of ground-truth text (Section 7 & Appendix F). This type of experiment differs from the main type of results that we showcase in this paper, which use probabilistic discoverable extraction to quantify memorization (Appendix A). We have made a significant effort throughout to make this point clear. If something remains unclear, please reach out to the corresponding authors.
Final comments on amount of books memorization
We were able to extract relatively enormous amounts of memorized training data in some cases, from both public domain books (e.g.,gatsby, alice, ulysses) and popular in-copyright books of all stripes (e.g., hp1, thrones, leanin, davinci, hobbit, sisyphus). In particular, we intend to investigate this further for Llama 3+ models, which exhibited an order of magnitude more memorization on some books than our worst-case estimates prior to starting this project. It was this enormous extent of memorization that encouraged us to try to reconstruct a book starting from a single seed prompt; we would not have believed such an outcome was possible prior to measuring memorization in Llama 3+ models—certainly not a flagship model from a frontier company, which continues to be downloaded and used by millions of users (Figure 122).
Figure 122: Screenshot of the downloads of Llama 3.1 70B on HuggingFace, May 2025.